### v2

# Soft-QMIX: Integrating Maximum Entropy For Monotonic Value Function Factorization

Abstract
Multi-agent reinforcement learning (MARL) tasks often utilize a centralized training with decentralized execution (CTDE) framework. QMIX is a successful CTDE method that learns a credit assignment function to derive local value functions from a global value function, defining a deterministic local policy. However, QMIX is hindered by its poor exploration strategy. While maximum entropy reinforcement learning (RL) promotes better exploration through stochastic policies, QMIX’s process of credit assignment conflicts with the maximum entropy objective and the decentralized execution requirement, making it unsuitable for maximum entropy RL. In this paper, we propose an enhancement to QMIX by incorporating an additional local Q-value learning method within the maximum entropy RL framework. Our approach constrains the local Q-value estimates to maintain the correct ordering of all actions. Due to the monotonicity of the QMIX value function, these updates ensure that locally optimal actions align with globally optimal actions. We theoretically prove the monotonic improvement and convergence of our method to an optimal solution. Experimentally, we validate our algorithm in matrix games, Multi-Agent Particle Environment and demonstrate state-of-the-art performance in SMAC-v2.  [^1] Our code: https://github.com/WentseChen/Soft-QMIX
Figure 1: Soft-QMIX combines the QMIX value decomposition mechanism with maximum entropy RL under the CTDE framework. The figure explains the algorithm design rationale. The left figure illustrates a naive approach to applying maximum entropy RL in the CTDE context, where blue text represents the desired objectives, and black text indicates corresponding constraints, It reveals that incorporating maximum entropy RL under the CTDE framework implicitly constrains the global value to be the sum of local values, significantly limiting the expressiveness of the critic network. The right figure depicts the Soft-QMIX improvement, where Soft-QMIX first employs QMIX’s value decomposition mechanism to obtain $Q_{i}$. Given the meaningful order of QMIX’s output $Q_{i}$, such as possessing properties like IGM, an order-preserving transformation converts $Q_{i}$ to $f_{i}(Q_{i})$. This transformation is trained by the loss function that minimize the difference between $f_{i}(Q_{i})$ and $Q_{tot}.$
1 Introduction
In the context of solving cooperative tasks using deep multi-agent reinforcement learning (MARL), a standard approach is centralized training with decentralized execution (CTDE) [5, 23, 15]. Decentralized execution implies that during the execution phase, the decision-making of each agent is independent and limited to conditioning on local observations. Centralized training, on the other hand, allows the algorithm to access joint actions and global states while training the value function. Compared to decentralized training, which treats other agents as part of the environment, centralized training helps alleviate the instability often encountered in MARL training [30].
In most cooperative tasks, the multi-agent system receives a single joint reward, and MARL algorithms learn the corresponding joint value function based on this reward. Consequently, there is a need to design a credit assignment mechanism [2] that decomposes the joint value function into several local value functions, guiding the updates of the actor networks through these local value functions. The challenge in credit assignment lies in reconstructing the joint distribution from the combination of multiple marginal distributions, a process that typically involves approximation errors [32]. Therefore, past research has primarily focused on exploring the impact of mechanism design on the model’s capacity [27, 31].
Beyond model capabilities, exploration ability is also a critical factor affecting the performance of MARL algorithms. Maximum entropy RL [14], known for encouraging exploration [7], smoothing the landscape of the objective [1], and enhancing robustness [4], has been widely adopted in single-agent RL. However, integrating MARL with maximum entropy RL is not straightforward. As illustrated in Figure 1, using maximum entropy RL within the framework of CTDE will significantly limit the model’s abilities in a manner akin to VDN [28], which enforces that the sum of local q-values must equal the global q-value.
In this paper, we propose a value-based maximum entropy MARL method compatible with the CTDE framework, named Soft-QMIX. We adopt the QMIX [23] mechanism to train a centralized critic network and utilize its credit assignment mechanism to obtain local Q-functions. We also incorporate maximum entropy RL, utilizing stochastic policies and entropy regularization to improve exploration and robustness. To guarantee monotonically increasing expected returns during training and convergence to the optimal policy, an order-preserving transformation on the original local Q-functions is used for the final policy. Through this design, Soft-QMIX can benefit from both maximum entropy RL and QMIX’s credit assignment mechanism, while still ensuring convergence to the optimal strategy.
Our main contributions can be summarized as follows:
  • We introduce Soft-QMIX, a value-based maximum entropy MARL algorithm that aligns with the CTDE framework. It employs an order-preserving transformation on local q-values, enabling it to derive meaningful order from credit assignment and leverage the exploratory benefits provided by maximum entropy RL.
  • Theoretically, we demonstrate that our approach can monotonically improve the expected q-value and guarantee convergence to the optimal policy.
  • Empirically, we confirm the effectiveness of our algorithm through tests on matrix games, MPE [18] and SMAC-v2 [3], where our method attains state-of-the-art results.
2 Related Works
Multi-Agent RL. In the CTDE framework, MARL primarily involves two common algorithmic approaches: policy gradient methods [17, 34, 15, 5, 13, 37] and value function decomposition methods [28, 23, 31]. Policy gradient methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence. On the other hand, Value function decomposition addresses the credit assignment problem by decomposing the global value function into multiple local value functions. VDN [28] assumes that the global Q-function is the sum of local Q-functions. QMIX [23, 20] permits the mixer function to be any function with non-negative weights. QTRAN [27] optimizes an additional pair of inequality constraints to construct a loss function. QPLEX [31] use the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize. The expressive power of the critic networks in these algorithms increases sequentially, but experimental results show that QMIX has better performance [9]. Therefore, in this work, QMIX is used as the mechanism for credit assignment.
Maximum Entropy RL. Maximum entropy RL [14] is demonstrated to have advantages such as encouraging exploration [7] and increasing robustness [4] in single-agent RL scenarios. In maximum entropy MARL, most works utilize an actor-critic architecture and typically aim to minimize the KL divergence between the local actor and the centralized critic [8, 6, 22, 36]. However, although these methods employ QMIX to train the critic, they are only capable of minimizing the discrepancy in expectation, thus cannot guarantee the IGM condition, which is central to the QMIX algorithm. Our approach is value-based and ensures that the IGM condition is met while employing maximum entropy MARL. Moreover, our method introduces global information into the local policy during training through a hyper-network, which most previous methods either ignored [35] or only applied in centralized training and centralized execution scenarios [16]. In appendix H, we provide a detailed comparison of our work with related works.
3 Preliminary
3.1 Multi-Agent Reinforcement Learning
Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [19]. Formally, a Dec-POMDP is represented as a tuple $(\mathcal{A},S,U,T,r,O,G,\gamma)$, where $\mathcal{A}\equiv\{1,...,n\}$ is the set of $n$ agents. $S$, $U$, $O$, and $\gamma$ are state space, action space, observation space, and discount factor, respectively.
During each discrete time step $t$, each agent $i\in\mathcal{A}$ chooses an action $u^{i}\in U$, resulting in a collective joint action $\mathbf{u}\in\mathbf{U}\equiv U^{n}$. The function $r(s,\mathbf{u})$ defines the immediate reward for all agents when the collective action $\mathbf{u}$ is taken in the state $s$. $\mathcal{P}(s,\mathbf{u},s^{\prime}):S\times\mathbf{U}\times S\rightarrow[0,\infty)$ is the state-transition function, which defines the probability density of the succeeding state $s^{\prime}$ after taking action $\mathbf{u}$ in state $s$. In a Dec-POMDP, each agent receives only partial observations $o\in O$ according to the observation function $G(s,i):S\times\mathcal{A}\rightarrow O$. For simplicity, if a function depends on both $o^{i}$ and $s$, we will disregard $o^{i}$. Each agent uses a policy $\pi_{i}(u^{i}|o^{i})$ to produce its action $u^{i}$ from its local observation $o^{i}$. Note that the policy should be conditioned on the history $o^{i}_{1:t}$. For simplicity, we refer to $o^{i}_{1:t}$ as $o^{i}_{t}$ or $o^{i}$ in the whole paper. The joint policy is denoted as $\pi_{jt}$. $\rho_{\pi}(s_{t},\mathbf{u_{t}})$ is used to represent the state-action marginals of the trajectory distribution induced by a policy $\pi$.
3.2 Individual-Global Condition
In this section, we define two key concepts in credit assignment.
Definition 3.1.
Individual-Global-Max (IGM) For joint q-function $Q_{tot}$, if there exist individual q-functions $\left[Q_{i}\right]_{i=1}^{n}$ such that the following holds:
$\underset{\mathbf{u_{t}}}{\mathrm{argmax}}\ Q_{tot}(\mathbf{u_{t}},s_{t})=\begin{pmatrix}\underset{u_{t}^{1}}{\mathrm{argmax}}\ Q_{1}(u^{1}_{t}|o^{1}_{t})\\ \vdots\\ \underset{u_{t}^{n}}{\mathrm{argmax}}\ Q_{n}(u^{n}_{t}|o^{n}_{t})\end{pmatrix}.$ (1)
Then, we say that $[Q_{i}]_{i=1}^{n}$ satisfy IGM for $Q_{tot}$ under $s_{t}$.
In QMIX, to ensure the validity of IGM, the algorithm constrains the expressive capability of $Q_{tot}$ to the following form: $\frac{\partial Q_{tot}}{\partial Q_{i}}\geq 0,\forall i\in\mathcal{A}.$ Restricting the capacity of $Q_{tot}$ using the above inequality results in $Q_{tot}$ being a monotonic function. The monotonicity constraint serves as a sufficient, but not necessary, condition for IGM. Consequently, the monotonicity constraint is more stringent than those required by IGM.
Definition 3.2.
Individual-Global-Independence (IGI) For joint policy $\pi_{jt}$, if there exist individual policies $\left[\pi_{i}\right]_{i=1}^{n}$ such that the following holds:
$\pi_{jt}(\mathbf{u_{t}}|s_{t})=\prod_{i=1}^{n}\pi_{i}(u^{i}_{t}|o^{i}_{t}).$ (2)
Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGI for $\pi_{jt}$ under $s_{t}$.
Definition 3.1 characterizes the properties we expect a deterministic policy to possess. Definition 3.2, on the other hand, describes the independence of decision-making among agents during decentralized execution. IGI and the IGO [36] condition are very similar; we show in Appendix B that IGI is a sufficient condition for IGO.
3.3 Naive Maximum Entropy MARL Algorithm
In this section, we informally introduce a naive MARL algorithm. We use it to illustrate the issue that naively applying credit assignment to the maximum entropy RL algorithm will lead to limited credit assignment capability.
Assume a value-based framework, where $Q_{i}$ denotes the local Q-function and the function for credit assignment is denoted as $f^{c}$. Hence, the global Q-function, $Q_{tot}$, can be expressed as $Q_{tot}(\mathbf{u},s)=f^{c}\big{(}Q_{1}(u^{i},o^{i}),...,Q_{n}(u^{n},o^{n}),s\big{)}$. Based on maximum entropy RL, the global policy is defined as $\pi_{jt}(\cdot|s)\propto\text{softmax}\big{(}Q_{tot}(\cdot,s)\big{)}$ and the local policy as $\pi_{i}(\cdot|o^{i})\propto\text{softmax}\big{(}f^{t}_{i}(Q_{i}(\cdot,o^{i}),s)\big{)}$, where $f^{t}_{i}$ is an arbitrary function. Due to the IGI condition, we have $f^{c}(Q_{1},...,Q_{n},s)=\sum_{i=1}^{n}f^{t}_{i}(Q_{i},s)$. When $f^{t}_{i}$ is an identity mapping, the aforementioned algorithm is equivalent to VDN [28]. When $f^{t}_{i}$ represents a linear transformation, the algorithm becomes analogous to the DOP, an actor-critic method [33]. In fact, $f^{t}_{i}$ can be any function of $u^{i}$ and $o^{i}$, but cannot be a function of $u^{j}$ or $o^{j}$, where $j\neq i$. This constraint limits the expressiveness of the critic network.
4 Soft-QMIX
In this paragraph, we first briefly introduce the motivation behind the Soft-QMIX algorithm. Then, we detail the process of the Soft-QMIX algorithm and show that the expected returns of the centralized training are monotonically increasing and converge to the optimum. Finally, we propose a practical algorithm.
4.1 Motivation
As shown in Figure 2, we divide the decision-making process into two stages (green and blue blocks). In the first stage, agents rank the Q-values for all actions. We employ the value decomposition mechanism from QMIX to obtain local Q-values but only utilize the order of their outputs, not the specific Q-function values. In the second stage, without changing the order, we assign specific Q-values corresponding to each action. When selecting actions in decentralized execution, only the maximum value is required, necessitating just the first stage for execution. The second stage serves primarily to obtain an accurate estimated q-value. During this phase, we can utilize global information, as it is only used during training.
4.2 Soft Policy Iteration
The objective function of Maximum Entropy RL can be represented in the following form:
$J(\pi_{jt})$ $=\mathbb{E}_{(s_{t},\mathbf{u}_{t})\sim\rho_{\pi_{jt}}}\big{[}r(s_{0},\mathbf{u}_{0})+\sum_{t=1}^{\infty}\gamma^{t}\big{(}r(s_{t},\mathbf{u}_{t})+\alpha H\left(\pi_{jt}(\cdot|s_{t})\big{)}\right)\big{]},$ (3)
where $H(\pi_{jt})$ denotes the entropy of the joint policy $\pi_{jt}$, and the temperature parameter $\alpha$ is used to control the degree of randomness in the policy.
In the policy evaluation step, we aim to approximate the objective function described in Equation 3 using $Q_{tot}:S\times\mathbf{U}\rightarrow\mathbb{R}$. The soft state value function $V_{tot}:S\rightarrow\mathbb{R}$ is defined as follows:
$V_{tot}(s)=\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[Q_{tot}(s,\mathbf{u})-\alpha\log\pi_{jt}(\mathbf{u}|s)].$ (4)
Repeatedly applying the Bellman operation to update $Q_{tot}$ results in the soft $Q_{tot}$ function for policy $\pi_{jt}$. The Bellman operation is defined as follows:
$\mathcal{T}^{\pi_{jt}}Q_{tot}(s,\mathbf{u})=r(s,\mathbf{u})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}}[V_{tot}(s^{\prime})].$ (5)
In reference to the QMIX, a monotonicity constraint is imposed on the capacity of the critic network. Specifically, $Q_{\text{tot}}$ can be formulated as follows:
$Q_{tot}(s,\mathbf{u})=\text{Mixer}\left(Q_{1}(o^{1},u^{1}),...,Q_{n}(o^{n},u^{n}),s\right).$ (6)
The network structure of the Mixer function is the same as QMIX’s, and is elaborated in Appendix G.
Due to the monotonicity constraint, there is moderate error in estimating of $Q_{tot}$. To enhance the precision of value estimation, we introduce the TD($\lambda$[24] method. The estimation error between the maximum entropy objective function and the Bellman equation objective can be expressed as follows:
$\delta_{t}^{V}$ $=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1}),$ (7)
$\epsilon_{t}^{td(\lambda)}$ $=\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u_{t}})-Q_{tot}(s_{t},\mathbf{u_{t}})=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{V}_{t+l}.$
Figure 2: The overall network architecture. Functions $g_{i}$ and $f_{i}$ are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values. The output of the $f_{i}$ function, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled. For decentralized execution, only the green network is needed, and the maximum values of their output are selected.
The derivation and implementation details of Equation 7 are provided in Appendix D. When combining off-policy methods with TD($\lambda$), adopting $\lambda>0$ requires importance sampling corrections [10]. Following previous work [9], we opt for a small $\lambda$ and assume that $\mathcal{T}^{\pi_{jt}}_{\lambda}$ is equivalent to $\mathcal{T}^{\pi_{jt}}$ in expectation. For simpilicity, we will use $\mathcal{T}^{\pi_{jt}}$ instead of $\mathcal{T}^{\pi_{jt}}_{\lambda}$ in theoretical discussions.
The following lemma demonstrates that by iteratively applying the Bellman equation, convergence can be achieved towards the maximum entropy objective $Q_{\text{tot}}$ for $\pi_{jt}$.
Lemma 4.1 (Joint Soft Policy Evaluation).
Let $Q^{0}_{jt}:S\times\mathcal{U}\to\mathbb{R}$ be a bounded mapping where $|\mathcal{U}|<\infty$. Define the iterative process $Q^{k+1}_{tot}=\mathcal{T}^{\pi_{jt}}(Q^{k}_{tot})$. It follows that the sequence $\{Q^{k}_{tot}\}_{k=0}^{\infty}$ converges to the joint soft-Q-function for policy $\pi_{jt}$ as $k\rightarrow\infty$.
Proof. See Appendix C. ∎
In Section 4.1, it was mentioned that decision-making is divided into two stages. During training, the policy is defined as $\pi^{train}_{i}(\cdot|o^{i},s)=\text{softmax}(f^{\text{stage2}}_{i}(f^{\text{stage1}}_{i}(\cdot|o^{i}),s)/\alpha)$, where $f^{\text{stage1}}_{i}$ is a function dependent on local observations, responsible for the first stage of decision-making. $f^{\text{stage2}}_{i}$, on the other hand, is a non-linear order-preserving transformation, responsible for the second stage of decision-making. Although $\pi^{train}_{i}$ depends on the global state, which conflicts with the definition in Dec-POMDP, it remains a valid option because it is utilized solely during training. In the subsequent paragraph, $\pi$ will denote $\pi^{train}_{i}$. $f^{\text{stage2}}_{i}$ can be further decomposed into two order-preserving transformations, $g_{i}$ and $f_{i}$. The output of $g_{i}$ is $Q_{i}$, and $f_{i}$ further transforms $Q_{i}$ into the final policy. The form of order-preserving transformation (OPT) is as follows:
$OPT_{i}(x,s)=W_{{}_{(1\times d_{1})}}^{2}\sigma(W^{1}_{{}_{(d_{1}\times 1)}}x_{{}_{(1\times d_{2})}}+b^{1})+b^{2},$ (8)
where $d_{1}$ is the hidden dimension, $d_{2}$ is the dimension of actions, $b^{1}$ and $b^{2}$ are the bias, $\sigma$ is an activation function, and $W^{1}$ and $W^{2}$ are non-negative weights. The implementation of the order-preserving transformation is similar to QMIX’s and is elaborated in Appendix G, where its weights and biases are provided by a hypernetwork. This hypernetwork takes the global state $s$ as input, making $g_{i}$ and $f_{i}$ functions dependent on the global state. Figure 2 presents the overall framework, where Agents $1$ to Agents $n$ correspond to $f_{i}^{\text{stage1}}$, and $g_{i}$ and $f_{i}$ correspond to $f_{i}^{\text{stage2}}$.
Since $g_{i}$ is an order-preserving transformation, and $[Q_{i}]_{i=1}^{n}$ satisfy the IGM condition for $Q_{tot}$, it follows that $\underset{\mathbf{u}}{\text{argmax}}\,Q_{tot}(\mathbf{u},s)=\left(\underset{u^{1}}{\text{argmax}}\,f_{1}^{\text{stage1}}(u^{1}|o^{1}),\ldots,\underset{u^{n}}{\text{argmax}}\,f_{n}^{\text{stage1}}(u^{n}|o^{n})\right)$. Therefore, during execution, only $f_{i}^{\text{stage1}}$ is needed.
During policy improvement, we update the joint policy towards the exponential of the joint Q-function. We constrain $\pi\in\Pi$, where $\Pi$ is the family of functions that satisfy the IGI condition. In the updating process, the updated policy is projected onto $\Pi$. Specifically, the update formula is as follows:
$\pi_{jt}^{\text{new}}=\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{arg\,min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (9)
[…]
In value-based methods, Equation 9 is intractable. Therefore, in the policy improvement step, we update the policy in this way:
$f_{i}^{\text{new}}=\underset{f^{\prime}_{i}}{\mathrm{argmin}}\frac{1}{2}\left(\sum_{i}f^{\prime}_{i}(Q_{i}(o^{i}_{t},u^{i}_{t},s_{t}))-y_{t}\right)^{2},$ (10)
where $y_{t}$ can be any unbiased estimate of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})\right]$. Here, we assume [^2] This does not necessarily hold in practical implementations, but the difference between these is a term in the loss function of our algorithm. For further discussion, please see the Appendix F. that $\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u_{t}})$ is an unbiased estimate of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}\sum_{i}f_{i}(Q_{i}(s_{t},\mathbf{u_{t}}))\right]$. Under this assumption, we can replace $Q_{\theta}$ and $y_{t}$ in Theorem 4.2 with $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$ respectively, and get that minimizing the L2 norm described in Equation 10 is equivalent to minimizing the KL divergence described in Equation 9 and minimizing the TD loss of value function at the same time. Therefore, in value-based methods, Equation 10 can be used in place of Equation 9.
Theorem 4.2.
If $Q$-function is parameterized as $Q_{\theta}$, and $V_{\theta},\pi_{\theta}$ are defined as below:
$V_{\theta}(s)$ :-$\coloneq\alpha\log\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\exp\left(Q_{\theta}(s,\mathbf{u})/\alpha\right)\right],:$ (11)
$\pi_{jt_{\theta}}(\mathbf{u}|s)$ :-$\coloneq\exp\left(Q_{\theta}(s,\mathbf{u})-V_{\theta}(s)/\alpha\right).:$
and $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right].$ Then, we have
$\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}$ $\left[\|Q_{\theta}(s,\mathbf{u})-y\|^{2}\right]$ (12)
$=\alpha\nabla_{\theta}D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}+\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\|V_{\theta}(s)-y\|^{2}\right]$
Proof. See Appendix E.
The subsequent theorem ensures that updating the policy using Equation 9 results in a monotonic increase of the Q-function under the IGI condition.
Lemma 4.3 (Joint Soft Policy Improvement).
Suppose $\pi_{jt}^{\text{old}}\in\Pi$ and consider $\pi_{jt}^{\text{new}}$ to be the solution to Equation 9. Then $Q^{\pi_{jt}^{\text{new}}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})$ for all $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$ with $|\mathbf{U}|<\infty$.
Proof. See Appendix C.
Theorem 4.4 demonstrates that iteratively applying policy evaluation and policy improvement leads to a globally optimal policy under the IGI condition.
Theorem 4.4 (Joint Soft Policy Iteration).
By continually applying soft policy evaluation and soft policy improvement to any $\pi_{jt}\in\Pi$ converges to a policy $\pi^{*}$ such that $Q^{\pi^{*}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi}(s_{t},\mathbf{u}_{t})$ for all $\pi_{jt}\in\Pi$ and $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$, assuming $|\mathbf{U}|<\infty$.
Proof. See Appendix C.
4.3 The Soft-QMIX Algorithm
Figure 2 illustrates the network architecture of the Soft-QMIX algorithm, which comprises four types of networks. The Mixer network is responsible for handling credit assignment and is parameterized along with its corresponding hypernetwork as $\text{Mixer}_{\theta}$. The $[g_{i}]_{i=1}^{n}$ and $[f_{i}]_{i=1}^{n}$ are $2n$ order-preserving transformations, with each $g_{i},f_{i}$ and their corresponding hypernetwork parameterized as $g_{i,\theta}$ and $f_{i,\theta}$. Agents 1 to $n$ share parameters, also referred to as $f_{i}^{\text{stage1}}$ in the previous section, and are parameterized as $f^{\text{stage1}}_{\theta}$. The updated rules for these parameters are described as follows.
Soft-QMIX aims to minimize two loss functions. $\text{Mixer}_{\theta}$, $g_{i,\theta}$ and $f^{\text{stage1}}_{\theta}$ are updated using the following loss function:
$L_{Q}(\theta)=\frac{1}{2}\left(Q_{tot_{\theta}}(s_{t},\mathbf{u}_{t})-\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2},$ (13)
where $Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big{(}g_{1,\theta}(f^{\text{stage1}}_{\theta}(o^{1},u^{1}),s),\ldots,$ $g_{n_{\theta}}(f^{\text{stage1}}_{\theta}(o^{n},u^{n}),s),s\big{)}$, and $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u}_{t})$ is defined in Equation 7. We utilize the technique of target Q networks to stabilize training, where the target Q network is parameterized as $\bar{\theta}$, with $\bar{\theta}$ being the exponentially weighted moving average of $\theta$. Subsequently, the parameter $\phi$ is updated using the following formula:
$L_{\pi}(\phi)=\frac{1}{2}\left(\sum_{i=1}^{n}f_{i,\phi}(Q_{i}(o^{i}_{t},u^{i}_{t}),s_{t})-Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2}.$ (14)
In this process, the function $f_{i,\phi}$ takes $Q_{i}(o^{i}_{t},u^{i}_{t})$ as input, but during backpropagation, gradients are truncated at $Q_{i}$. This truncation ensures that this loss function updates only $f_{i,\phi}$. After the update, the policy can be expressed in the following form:
$\pi_{i}(u^{i}|o^{i},s)=\text{softmax}(\frac{f_{i}\left(g_{i}\big{(}f^{stage1}(o^{i}),s\big{)},s\right)}{\alpha})$ (15)
It’s important to note that during execution, we only need to call $f^{stage1}$ and apply the argmax operation to it. In this paper, $\alpha$ is fixed during training. Finally, the algorithmic procedure of Soft-QMIX is shown in appendix A.
5 Experiments
In this section, we evaluate our algorithm using matrix games from prior works [27], demonstrating its ability to consistently reach the global optimum. Then, we compare Soft-QMIX with QMIX on MPE [18] and empirically demonstrate Soft-QMIX algorithm achieves state-of-the-art performance on the SMAC-v2 [3] benchmark. We compare our algorithm with four baseline algorithms. MAPPO [34] adapts the on-policy, single-agent PPO [26] algorithm for multi-agent use, utilizing a centralized value function. QMIX [23] is a value-based method employing a monotonic value decomposition to create multiple local deterministic policies from a shared value function. It uses an $\epsilon$-greedy strategy for exploration. FOP [36] implements a maximum entropy policy gradient approach with value factorization similar to QPLEX. HASAC [16] offers an alternative multi-agent soft actor-critic framework, facilitating sequential decision-making by agents in a centralized training with centralized execution context. We elaborate implementation details and hyperparameters in Appendix J.
5.1 Matrix Games
In this section, we demonstrate the effectiveness of our algorithm in the classic one-step matrix games. Table 1 presents a classic non-monotonic payoff matrix, often used in past papers [27, 31] to discuss the representational capability of critic networks. QMIX, due to its monotonicity constraint, often incurs moderate errors in fitting such payoff matrices. However, in practice, during the training process, we are primarily concerned with the value estimates of those joint actions that are frequently visited. For suboptimal joint actions with lower Q values, their errors do not significantly impact the estimated expected returns. As we use TD-$\lambda$ to estimate returns in our experiments, such errors can be further reduced. Conversely, we are more concerned with whether the joint actions with the highest probability of being selected have the globally maximal value function, as we aim for the highest returns when executing with a deterministic policy. For suboptimal joint actions, we also wish to visit them with a certain probability to ensure adequate exploration.
Table 1: Payoff matrix of the one-step matrix game.
A B C
A 8 -12 -12
B -12 0 0
C -12 0 0
In the matrix game shown in Table 2, QMIX exhibits high estimation error due to its monotonicity constraints, and its deterministic policy lacks effective exploration, which leads it to learn only suboptimal joint actions with a reward of 00. FOP, with its network structure similar to QPLEX, can accurately estimate value functions, yet the corresponding policy learned tends to select suboptimal joint actions more frequently. Soft-QMIX has smaller estimation errors for optimal joint actions and larger errors for suboptimal ones. However, the joint action with the maximum Q value in Soft-QMIX matches the true optimal joint action, and the local policy has the highest probability of selecting the optimal joint action. Soft-QMIX possesses this capability because its decision-making is divided into two steps, and in the first step, it ensures the order of probabilities of each action being selected in the policy.
Table 2: (a)-(c) respectively demonstrate the results of Soft-QMIX, FOP, and QMIX in the one-step matrix game. The first row and column display the policies adopted by the row and column players after $10k$ steps of training, respectively. The 3$\times$3 matrix in the bottom right corner represents the $Q_{\text{tot}}$ ultimately learned by the algorithm. The results indicate that Soft-QMIX is the only method among all that consistently chooses the optimal action with the highest probability.
1. 0. 0.
1. 8 -12 -14
0. -8 -28 -30
0. -10 -30 -32
(a) result of Soft-QMIX
.21 .39 .40
.21 8 -12 -12
.40 -12 0 0
.39 -12 0 0
(b) result of FOP
$\epsilon$ 1-$2\epsilon$ $\epsilon$
$\epsilon$8.5 -8.5 -8.5
1-$2\epsilon$8.5 0.2 0.2
$\epsilon$8.5 0.2 0.2
(c) result of QMIX
5.2 Multi-Agent Particle Environment
In this section, we test our algorithm on three custumed Multi-Agent Particle Environment(MPE) [18] scenarios and show that Soft-QMIX do better exploration compare to QMIX and ablate through different choice of $\alpha$. Specifically, we use three different Spread map, where in each map, agents are tasked to navigate to all the landmarks without collide with each other. Harder map means longer initial distance between agents and landmarks. Details are available in Appendix I. Table 3 shows that Soft-QMIX significantly outperforms QMIX on the Hard and Medium map which require more exploration, but shows only a slight advantage on the Simple map. Additionally, we find that higher $\alpha$ values enhance exploration and performance. These findings support our methodological claim that integrating maximum entropy RL with MARL algorithms boosts exploration capabilities. For each scenario, we conduct four experiments across five seeds, reporting the average evaluation episode rewards after 10M environment steps and their standard deviation in our results table.
Table 3: Experiment result in three MPE scenarios, where Soft-QMIX has higher performance gain from QMIX on Hard and Medium map, which need more exploration. Higher $\alpha$ also lead to better exploration and higher performance gain.
eval episode QMIX Soft-QMIX(ours)
rewards $\alpha=0.1$ $\alpha=0.3$ $\alpha=1.0$
Spread(Hard) -171.7 $\pm$ 3.8 -173.7 $\pm$ 3.0 -170.8 $\pm$ 3.1 -168.7 $\pm$ 4.6 (+1.8%)
Spread(Medium) -138.4 $\pm$ 2.9 -139.0 $\pm$ 2.0 -137.4 $\pm$ 3.3 -134.9 $\pm$ 1.8 (+2.5%)
Spread(Simple) -111.9 $\pm$ 1.6 -116.2 $\pm$ 2.7 -114.4 $\pm$ 2.0 -111.4 $\pm$ 0.9 (+0.4%)
5.3 SMAC-v2
We evaluated Soft-QMIX on the SMACv2 [3] benchmark, including six scenarios: protoss_5_vs_5, terran_5_vs_5, zerg_5_vs_5, protoss_10_vs_10, terran_10_vs_10, and zerg_10_vs_10. The results, as shown in Figure 3, were obtained by training each scenario for $10$ million environmental steps, with each curve representing the average of 5 seeds, and HASAC with an average of 3 seeds. Soft-QMIX outperforms the baseline algorithms in all scenarios. MAPPO, being an on-policy algorithm, has lower sampling efficiency, resulting in lower win rates for models trained under 10M environmental steps. Training with MAPPO for 60M environmental steps can achieve win rates comparable to 10M QMIX. QMIX, lacking exploration, tends to converge to local optima, especially in the two Zerg scenarios. HASAC and FOP perform poorly in all the scenarios. Ultimately, Soft-QMIX improves win rates by $5\%$ to $24\%$ in all scenarios compared to baseline algorithms.
Figure 3: Experiment result in six SMAC-v2 scenarios, where Soft-QMIX achieved faster convergence speed and higher performance compared to baseline algorithms.
5.4 Ablation Study
We evaluate the effectiveness of different components in our algorithm through an ablation study. We begin with QMIX, which utilizes epsilon-greedy exploration. We then introduce +entropy, which incorporates softmax sampling (employing $\text{softmax}(Q_{i})$ as the policy) along with a maximum entropy objective function to the QMIX framework. Note that this version of the algorithm does not ensure convergence, leading to a performance decline in two out of three scenarios. The subsequent modification, +function f, integrates the function $f_{i}$ (using $\text{softmax}(f_{i}(Q_{i})$) as the policy) building upon the +entropy setup. Empirical results show that function f yields the most substantial performance improvement across all components. Our study culminates with Soft-QMIX, our finalized algorithm. For each modification, we conduct four experiments across three seeds, reporting the average winning rate and its standard deviation in our results table.
Table 4: The ablation studies reveal that each component enhances performance, with function $f$ contributing most significantly to the performance improvement.
eval win rate QMIX +entropy +function f Soft-QMIX(ours)
protoss 5vs5 0.68 $\pm$ 0.04 0.73 $\pm$ 0.05 0.74 $\pm$ 0.04 0.76 $\pm$ 0.04
terran 5vs5 0.68 $\pm$ 0.05 0.67 $\pm$ 0.04 0.70 $\pm$ 0.05 0.72 $\pm$ 0.04
zerg 5vs5 0.41 $\pm$ 0.05 0.40 $\pm$ 0.04 0.48 $\pm$ 0.05 0.51 $\pm$ 0.04
6 Conclusion
In this paper, we introduced Soft-QMIX, an innovative approach within the CTDE framework, integrating the maximum entropy RL framework with the QMIX-based credit assignment mechanism. Theoretically, we established the monotonic improvement and convergence of Soft-QMIX to a globally optimal solution. Experimentally, our algorithm can solve matrix games and demonstrate state-of-the-art performance on the SMAC-v2 benchmark. Currently, Soft-QMIX is a value-based method, thus one of its limitations is that it can currently only be deployed in tasks with discrete action spaces. In the future, we will try to extend it to tasks with continuous action spaces.
Acknowledgment
This work was supported in part by the U.S. Army Futures Command under Contract No. W519TC-23-C-0030.
Appendix A Soft-QMIX
Algorithm 1 Soft-QMIX
1:  Input: $\theta,\phi$
2:  Initiate: $\bar{\theta}\leftarrow\theta,D\leftarrow\emptyset$
3:  for each iteration do
4:     for each environment step do
5:        get $\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t})$ via Eq. 15
6:        $u^{i}_{t}\sim\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t}),\forall i\in\{1,..,n\}$
7:        $s_{t+1}\sim p(s_{t+1}|s_{t},\mathbf{u}_{t})$
8:        $D\leftarrow D\cup\{(s_{t},\mathbf{u}_{t},r(s_{t},\mathbf{u}_{t}),s_{t+1})\}$
9:     end for
10:     for each gradient step do
11:        calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq. 24
12:        $\theta\leftarrow\theta-\lambda_{Q}\nabla_{\theta}L_{Q}(\theta)$
13:        $\phi\leftarrow\phi-\lambda_{\pi}\nabla_{\phi}L_{\pi}(\phi)$
14:        $\bar{\theta}\leftarrow\tau\theta+(1-\tau)\bar{\theta}$
15:     end for
16:  end for
Appendix B Individual-Global-Independence (IGI)
In this section, we first introduce two concepts, IGI and IGO, where IGI is proposed in our work, and IGO was proposed by A. We then explain that IGI is a sufficient condition for IGO.
Definition B.1.
Individual-Global-Independence (IGI) For joint policy $\pi_{jt}$, if their exist individual policies $\left[\pi_{i}\right]_{i=1}^{n}$ such that the following holds:
$\pi_{jt}(\mathbf{u}|s)=\prod_{i=1}^{n}\pi_{i}(u^{i}|o^{i}).$ (16)
Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGI for $\pi_{jt}$ under $s_{t}$.
Definition B.2.
Individual-Global-Optimal (IGO) For optimal joint policy $\pi^{*}_{jt}$, if their exist individual optimal policies $\left[\pi^{*}_{i}\right]_{i=1}^{n}$ such that the following holds:
$\pi^{*}_{jt}(\mathbf{u}|s)=\prod_{i=1}^{n}\pi^{*}_{i}(u^{i}|o^{i}).$ (17)
Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGO for $\pi_{jt}$ under $s_{t}$.
IGO defines that the optimal joint policy equals the product of optimal local policies. In IGI, it is defined that any joint policy should equal the product of local policies. If finding the optimal policy is viewed as an optimization problem, IGI defines the feasible set of this problem. Hence, any policy not equaling the product of local policies is not a feasible solution and thus cannot be optimal. Therefore, all optimal policies will satisfy the joint policy equals the product of local optimal policies, thereby meeting the IGO condition.
Note that under the assumption of a decentralized execution setting, IGI and IGO are equivalent.
Appendix C Joint Policy Iteration
Lemma C.1 (Joint Soft Policy Evaluation).
Let $Q^{0}_{tot}:S\times\mathbf{U}\to\mathbb{R}$ be a bounded mapping where $|\mathbf{U}|<\infty$. Define the iterative process $Q^{k+1}_{tot}=\mathcal{T}^{\pi_{jt}}_{\lambda}(Q^{k}_{tot})$. It follows that the sequence $\{Q^{k}_{tot}\}_{k=0}^{\infty}$ converges to the joint soft-Q-function for policy $\pi_{jt}$ as $k\rightarrow\infty$.
Proof.
Initiate by defining the entropy-augmented reward function:
$r_{\pi_{jt}}(s_{t},\mathbf{u}_{t})\coloneqq r(s_{t},\mathbf{u}_{t})+\mathbb{E}_{s_{t+1}}\left[H[\pi_{jt}(\cdot|s_{t+1})]\right].$ (18)
Follow previous work [9], we choose a lower $\lambda$ and can assume that $\mathcal{T}^{\pi_{jt}}_{\lambda}$ is an unbiased estimator of $\mathcal{T}^{\pi_{jt}}$. Subsequently, the update mechanism is articulated as:
$Q_{tot}(s_{t},\mathbf{u}_{t})$ $\leftarrow r_{\pi_{jt}}(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1},\mathbf{u}_{t+1}\sim\pi_{jt}}[Q_{tot}(s_{t+1},\mathbf{u}_{t+1})].$ (19)
Conclude by invoking the established convergence theorems pertinent to policy evaluation as explicated by Sutton and Barto. ∎
Lemma C.2 (Joint Soft Policy Improvement).
Suppose $\pi_{jt}^{\text{old}}\in\Pi$ and consider $\pi_{jt}^{\text{new}}$ to be the solution to Equation 9. Then $Q^{\pi_{jt}^{\text{new}}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})$ for all $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$ with $|\mathbf{U}|<\infty$.
Proof.
Let $\pi_{jt}^{\text{old}}\in\Pi$ and let $Q^{\pi_{jt}^{\text{old}}}$ and $V^{\pi_{jt}^{\text{old}}}$ be the corresponding soft state-action value and soft state value. The update rule of $\pi_{\text{new}}$ can be defined as
$\pi_{\text{new}}(\cdot\mid s_{t})$ $=\arg\min_{\pi^{\prime}\in\Pi}\text{D}_{KL}\left(\pi^{\prime}(\cdot)\parallel\exp(Q^{\pi_{jt}^{\text{old}}}(s_{t},\cdot))-\log Z^{\pi_{jt}^{\text{old}}}(s_{t})\right)$ (20)
$=\arg\min_{\pi^{\prime}\in\Pi}J_{\pi_{jt}^{\text{old}}}\left(\pi^{\prime}(\cdot\mid s_{t})\right).$
Since we can always choose $\pi_{\text{new}}=\pi_{\text{old}}\in\Pi$, the following inequality always hold true: $J_{\pi_{jt}^{\text{old}}}(\pi_{\text{new}}(\cdot\mid s_{t}))\leq J_{\pi_{jt}^{\text{old}}}(\pi_{jt}^{\text{old}}(\cdot\mid s_{t}))$. Hence
$\mathbb{E}_{\mathbf{u}_{t}\sim\pi_{jt}^{\text{new}}}\left[\log(\pi_{jt}^{\text{new}}(\mathbf{u}_{t}\mid s_{t}))-Q^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})+\log Z^{\pi_{jt}^{\text{old}}}(s_{t})\right]$ (21)
$\leq\mathbb{E}_{\pi_{jt}^{\text{old}}}\left[\log(\pi_{jt}^{\text{old}}(\mathbf{u}_{t}\mid s_{t}))-Q^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})+\log Z^{\pi_{jt}^{\text{old}}}(s_{t})\right],$
and the inequality reduces to the following form since partition function $Z^{\pi_{jt}^{\text{old}}}$ depends only on the state,
$\mathbb{E}_{\mathbf{u}_{t}\sim\pi_{jt}^{\text{new}}}\left[Q^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t}\mid s_{t})\right]\geq V^{\pi_{jt}^{\text{old}}}(s_{t}).$ (22)
Next, consider the soft Bellman equation:
$Q^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})$ $=r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}\sim p}\left[V^{\pi_{jt}^{\text{old}}}(s_{t+1})\right]$ (23)
$\leq r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}Q^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}$
$\vdots$
$\leq Q^{\pi_{jt}^{\text{new}}}(s_{t},\mathbf{u}_{t}),$
where we can repeatedly expand $Q^{\pi_{jt}^{\text{old}}}$ on the RHS by applying the soft Bellman equation and the bound in Equation 23. Convergence to $Q^{\pi_{jt}^{\text{new}}}$ follows from Lemma B.1.
Theorem C.3 (Joint Soft Policy Iteration).
By continually applying joint soft policy evaluation and joint soft policy improvement to any $\pi_{jt}\in\Pi$ converges to a policy $\pi^{*}_{jt}$ such that $Q^{\pi^{*}_{jt}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi_{jt}}(s_{t},\mathbf{u}_{t})$ for all $\pi_{jt}\in\Pi$ and $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$, assuming $|\mathbf{U}|<\infty$.
Proof.
Let $\pi_{jt}^{i}$ represent the policy after the $i$-th update. By Lemma B.2, the sequence $Q^{\pi_{jt}^{i}}$ is monotonically increasing. Given that $Q^{\pi}$ is bounded above for $\pi_{jt}\in\Pi$ due to bounded rewards and entropy terms, the sequence converges to some $\pi^{*}_{jt}$. It remains to demonstrate that the policy $\pi^{*}_{jt}$ is indeed the optimal policy. At convergence, we infer that $J_{\pi^{*}_{jt}}(s)\geq J_{\pi_{jt}}(s)$ for all $\pi_{jt}\in\Pi$, if $\pi_{jt}\neq\pi^{*}_{jt}$. Using the same iterative argument as in the proof of Lemma B.2, we get $Q^{\pi^{*}_{jt}}(s,\mathbf{u})\geq Q^{\pi}(s,a)$ for all $(s,\mathbf{u})\in S\times\mathbf{U}$, that is, the soft value of any other policy in $\Pi$ is lower than that of the converged policy. Hence $\pi^{*}_{jt}$ is optimal in $\Pi$. ∎
Appendix D TD($\lambda$)
In this paragraph, we will derive the expression of TD($\lambda$) in the context of Maximum Entropy Reinforcement Learning. Specifically, we aim to demonstrate the derivation of the following equation:
$\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u_{t}})-Q_{tot}(s_{t},\mathbf{u_{t}})=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{V}_{t+l}$ (24)
where $\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$.
Proof.
The one-step TD error can be expressed in the following form:
$\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$ (25)
$\delta_{t+1}^{V}=-Q_{tot}(s_{t+1},\mathbf{u}_{t+1})+r_{t+1}+\gamma V_{tot}(s_{t+2})$
Next, the k-step TD error can be represented in the following form:
$\epsilon_{t}^{(1)}:=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})=\delta^{V}_{t}$ (26)
$\epsilon_{t}^{(2)}:=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma r_{t+1}{\color[rgb]{1,0,0}-\log\pi_{jt}(\mathbf{u_{t+1}}|s_{t+1})}+\gamma^{2}V_{tot}(s_{t+1})=\delta^{V}_{t}+\gamma\delta^{V}_{t+1}$
$\epsilon_{t}^{(k)}=\sum_{l=0}^{k-1}\gamma^{l}\delta^{V}_{t+l}$
$\epsilon_{t}^{td(\lambda)}:=(1-\lambda)\left(\epsilon_{t}^{(1)}+\lambda\epsilon_{t}^{(2)}+\lambda^{2}\epsilon_{t}^{(3)}+\cdots\right)$ (27)
$=(1-\lambda)\big{(}\delta_{t}^{V}+\lambda(\delta_{t}^{V}+\gamma\delta_{t+1}^{V})+\lambda^{2}(\delta_{t}^{V}+\gamma\delta_{t+1}^{V}+\gamma^{2}\delta_{t+2}^{V})+\cdots\big{)}$
$=(1-\lambda)\left(\delta_{t}^{V}(1+\lambda+\lambda^{2}+\cdots)+\gamma\delta_{t+1}^{V}(\lambda+\lambda^{2}+\cdots)\right.\left.+\gamma^{2}\delta_{t+2}^{V}(\lambda^{2}+\lambda^{3}+\lambda^{4}+\cdots)+\cdots\right)$
$=(1-\lambda)\left(\delta_{t}^{V}\left(\frac{1}{1-\lambda}\right)+\gamma\delta_{t+1}^{V}\left(\frac{\lambda}{1-\lambda}\right)+\cdots\right)$
$=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{V}$
It is important to note that when computing the TD lambda error in the context of a reverse view and maximum entropy reinforcement learning, it is necessary to subtract the $\log(\pi)$ term from the $t+1$ step’s return as well as subtract $\log(\pi)$ from $Q_{t+1}$ in order to obtain results consistent with Equation 27.
Appendix E Soft-Q Learning and Policy Gradient
Theorem E.1.
If $Q$-function is parameterized as $Q_{\theta}$, and $V_{\theta},\pi_{\theta}$ are defined as below:
$V_{\theta}(s)$ :-$\coloneq\alpha\log\mathbb{E}_{\mathbf{u}\sim\pi_{jt}}\left[\exp\left(Q_{\theta}(s,\mathbf{u})/\alpha\right)\right],:$ (28)
$\pi_{jt_{\theta}}(\mathbf{u}|s)$ :-$\coloneq\exp\left(Q_{\theta}(s,\mathbf{u})-V_{\theta}(s)\right)/\alpha.:$
and $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right].$ Then, we have
$\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}$ $\left[\|Q_{\theta}(s,\mathbf{u})-y\|^{2}\right]$ (29)
$=\alpha\nabla_{\theta}D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$
$+\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\|V_{\theta}(s)-y\|^{2}\right]$
Proof.
The proof is heavily based on [25]. From assumptions described in Eq. 11 we have:
$Q_{\theta}(s,\mathbf{u})=V_{\theta}(s)+\alpha\log\pi_{jt_{\theta}}(\mathbf{u}|s)$ (30)
For simplicity, we further define notation $\Delta_{t}$ as below.
$\Delta_{t}$ $=\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})-Q_{\theta}(s_{t},\mathbf{u_{t}}).$
From assumption, we have $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right]$. Thus, the relationship between $\Delta_{t}$ and $y_{t}$ can be writen as:
$\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[y_{t}]$ $=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right]$ (31)
$=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[Q_{\theta}(s_{t},\mathbf{u_{t}})+\Delta_{t}]$
$=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[V_{\theta}(s_{t})+\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]+\Delta_{t}].$
Here are two equations that will be used in the following proof:
$\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s)\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u}|s)\big{)}\big{]}$ (32)
$=$ $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s)\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u}|s)]\big{)}\big{]}.$
$\mathbb{E}_{a\sim\pi}\big{[}\nabla\log\pi(a)f(s)\big{]}=f(s)\sum_{a}\pi(a)\frac{\nabla\pi(a)}{\pi(a)}$ (33)
$=f(s)\sum_{a}\nabla\pi(a)=f(s)\nabla\sum_{a}\pi(a)=0,$
where $f$ can be any function that does not conditioned on $a$, and $\sum_{a}\pi(a)=1$. Let’s consider the gradient of the $Q$-function’s MSE loss:
$\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\frac{1}{2}\left(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t}\right)^{2}\right]\bigg{|}_{\pi=\pi_{\theta}}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t})\right]\bigg{|}_{\pi=\pi_{\theta}}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\left(V_{\theta}(s_{t})+\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-y_{t}\right)\right]$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\big{(}V_{\theta}(s_{t})+\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-V_{\theta}(s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s_{t})+\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u}|s)\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}$
$\qquad-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}$
$\qquad-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-$
$\qquad\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}$
$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}.$
for simplicity, we ignore the $|_{\pi=\pi_{\theta}}$ after the first two lines of equations. The second equation use Eq. 30. The third equation use Eq. 31. The fourth equation cancled out $V_{\theta}(s_{t})$. The fifth equation use Eq. 30 again. The sixth equation expand out terms. The seventh equation use Eq. 32. The last equation use Eq. 33.
Let’s consider the gradient of the KL divergence.
$\nabla_{\theta}\alpha D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$
$=\nabla_{\theta}\mathbb{E}_{\pi_{jt}}\big{[}\alpha\log\pi_{jt}(\mathbf{u}|s)-y(\mathbf{u},s)+\alpha\log Z(s)\big{]}$
$=\nabla_{\theta}\mathbb{E}_{\pi_{jt}}\big{[}\alpha\log\pi_{jt}(\mathbf{u}|s)-y(\mathbf{u},s)\big{]}$
$=\nabla_{\theta}[\alpha\sum_{\mathbf{u}}\pi_{jt}(\mathbf{u}|s)\log\pi_{jt}(\mathbf{u}|s)]-\nabla_{\theta}[\sum_{\mathbf{u}}\pi_{jt}(\mathbf{u}|s)y(\mathbf{u},s)]$
$=\alpha\sum_{\mathbf{u}}\nabla_{\theta}\pi_{jt}(\mathbf{u}|s)\log\pi_{jt}(\mathbf{u}|s)-\sum_{\mathbf{u}}\nabla_{\theta}\pi_{jt}(\mathbf{u}|s)y(\mathbf{u},s)$
$=\mathbb{E}_{\pi_{jt}}\big{[}\alpha\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\log\pi_{jt}(\mathbf{u}|s)-\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}y(\mathbf{u},s)\big{]}.$
Note that $V_{\theta}(s)+\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u}|s)]$ is not a function of $\mathbf{u}$. With Eq. 33, we have:
$\nabla_{\theta}\alpha D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$ (34)
$=\mathbb{E}_{\pi_{jt}}\big{[}\alpha\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\log\pi_{jt}(\mathbf{u}|s)-\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\Delta_{t}\big{]}$
Finally, we got:
$\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\frac{1}{2}\left(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t}\right)^{2}\right]\bigg{|}_{\pi=\pi_{\theta}}$
$=\nabla_{\theta}\alpha^{2}D_{KL}\big{(}\pi_{jt}(\cdot|s_{t})\|\frac{exp(y(s)/\alpha)}{Z(s)}\big{)}+\frac{1}{2}\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}(V_{\theta}(s_{t})-y_{t})^{2}\big{]},$
where $y(s)$ is the distribution of $y_{t}$ at state $s$.
Appendix F Unbiased Assumption
We assume that $\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u_{t}})$ is an unbiased estimate of $E_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}\sum_{i}f_{i}(Q_{i}(s_{t},\mathbf{u_{t}}))\right]$ in the main paper. Practically, this assumption might fail, leading to non-monotonic increases in the policy improvement step. Here, we initially explain the impact of this error and derive an upper bound for it.
Conclusion: the error’s upper bound is proportional to $|Q_{tot}-\sum_{i}f_{i}(Q_{i})|$ and $\sqrt{D_{KL}\big{(}\pi_{jt}^{\text{old}}|\pi_{jt}^{\text{new}}\big{)}}.$ We use Equation 14 as the loss function, ensuring that the first term decreases throughout training. Additionally, by selecting a sufficiently small learning rate, the second term will also remain minimal.
Derivative: We can replace both $Q_{\theta}$ and $y_{t}$ in Theorem 4.2 with $\sum_{i}f_{i}(Q_{i})$, and get that minimizing the L2 norm described in Equation 10 is equivalent to minimizing the td loss of value function and minimizing the KL divergence described as follow,
$\pi_{jt}^{\text{new}}=\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{arg\,min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(\sum_{i}f_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (35)
[…]
Similar to Equation 22, during policy improvement step, we have:
$\mathbb{E}_{\mathbf{u}_{t}\sim\pi_{jt}^{\text{new}}}\left[\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t},u^{i}_{t}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t}\mid s_{t})\right]\leq V^{\prime\pi_{jt}^{\text{old}}}(s_{t}).$ (36)
During policy evaluation step, we have:
$Q_{tot}^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})$ $=r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}\sim p}\left[V_{tot}^{\pi_{jt}^{\text{old}}}(s_{t+1})\right]$ (37)
$=r(s_{t},\mathbf{u}_{t})+\gamma\big{(}\mathbb{E}_{s_{t+1}\sim p}\left[V^{\prime\pi_{jt}^{\text{old}}}(s_{t+1})\right]+\epsilon_{1}\big{)}$
$\leq r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}+\gamma\epsilon_{1}$
$=r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}+\gamma(\epsilon_{1}+\epsilon_{2})$
$\vdots$
$\leq Q_{tot}^{\pi_{jt}^{\text{new}}}(s_{t},\mathbf{u}_{t})+\epsilon,$
where $\epsilon$ is s the cumulative sum of $\epsilon_{1}$ and $\epsilon_{2}$ over time.$\epsilon_{1}$ and $\epsilon_{2}$ are defined as follows:
$\epsilon_{1}$ $=\mathbb{E}_{s_{t+1}\sim p}\left[V_{tot}^{\pi_{jt}^{\text{old}}}(s_{t+1})-V^{\prime\pi_{jt}^{\text{old}}}(s_{t+1})\right]$ (38)
$\epsilon_{2}$ $=\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
Then, we have:
$\epsilon_{1}+\epsilon_{2}$ $=\mathbb{E}_{s_{t+1}\sim p}\left[V_{tot}^{\pi_{jt}^{\text{old}}}(s_{t+1})-V^{\prime\pi_{jt}^{\text{old}}}(s_{t+1})\right]$ (39)
$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
$=\mathbb{E}_{s_{t+1}\sim p}\left[\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{old}}}\big{[}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{]}\right]$
$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
$=\mathbb{E}_{s_{t+1}\sim p}\big{[}\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{(}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{)}\big{]}$
$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{]},$
where $C=\underset{\mathbf{u}_{t+1}}{\mathrm{arg\,max}}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality [21], we have:
$|\epsilon_{1}+\epsilon_{2}|$ $\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big{]}$ (40)
$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sqrt{2D_{KL}\big{(}\pi_{jt}^{\text{old}}(\cdot|s_{t+1})|\pi_{jt}^{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$
Appendix G Hyper-network Structure
Figure 4: (a) The overall network architecture. Agent 1 to Agent n takes local observations as input and outputs a distribution over local actions $u^{i}$. Functions $g_{i}$ and $f_{i}$ are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values. Both the input and output of the $g_{i}$ and $f_{i}$ networks are k-dimensional, where $k$ is the action dimension. The output of the $f_{i}$ function, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled. The Mixer network takes local $Q_{i}$ as input and outputs the global $Q_{tot}$. Green blocks utilize a recurrent neural network as the backbone, with all agents sharing parameters. For decentralized execution, only the green network is needed, and the maximum values of their output are selected. (b) The blue blocks’ network structure. The red blocks represent the hyper-network, which takes the global state $s_{t}$ as input and outputs the network’s weights and biases. The weights are constrained to be non-negative. The activation function is denoted by $\sigma(\cdot)$. The mixer network takes an n-dimensional Q function as input and outputs a one-dimensional $Q_{tot}$.
Figure 4 shows the overall network structure. Specifically, function $f_{i}$ and function $g_{i}$ share the same network structure as the mixer network, but while the mixer network’s input dimension is (batch_size, agent_num) and its output dimension is (batch_size, 1), both the input and output dimensions of function $f_{i}$ and function $g_{i}$ are (batch_size, 1). Additionally, function $g_{i}$ introduces an additional residual from input to output. The network hyperparameters are as shown in the table below.
layer name hypernet_embed embed_dim num_layer
mixer 64 32 2
function $g_{i}$ 64 32 2
function $f_{i}$ 64 - 1
Table 5: Hyper-parameters used for hyper-networks.
Appendix H Comparison with Baselines
The following paragraphs present the differences between our method and those based on the AC:
1. Using existing maximum entropy MARL algorithms may result in the loss of action order learned through credit assignment. Achieving the IGM condition, which outlines the order of local value functions, is crucial in credit assignment. Our work presents a value-based algorithm that introduces order-preserving transformations, ensuring that the order of the local actor matches the order of the local value functions. However, the commonly used maximum entropy MARL methods, employing the actor-critic framework, may not guarantee alignment between the actor’s action order and that of the local Q-function due to approximation errors.
2. Actor-critic methods utilize the KL divergence as the loss function, while our approach uses the MSE loss. In the CTDE framework, all methods maintain a more accurate centralized Q-function alongside multiple imprecise local policies. The core idea is to distill knowledge from the centralized Q-function to the local policies. Actor-critic methods typically minimize the KL divergence between the actor and the softmax critic, whereas our method minimizes the MSE loss between their logits. [12] pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
3. AC methods can use global information when training the critic, resulting in higher model capacity. Our method introduces global information through hyper-networks and the mixer network, which results in lower model capacity.
4. In fact, it’s challenging to theoretically prove that value-based methods are inherently superior to AC methods. However, through experiments, we have demonstrated that our method surpasses both FOP [36] and HASAC [16].
Below are comparisons with individual papers.
mSAC [22] introduces a method similar to MASAC. It employs the AC architecture, thus encountering issues 1 and 2 mentioned above, but does not fully utilize global information to train the critic, hence lacking the advantage noted in point 3.
FOP [36] shares a similar issue with [22], in that the AC framework it utilizes can actually train the critic with global state information. Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX [31]. [9] experimentally shown that QMIX outperforms QPLEX. Additionally, FOP’s proof concept is that under the IGO conditions, if we locally move local policies towards local optimal policies, the distance between the joint policy and the optimal joint policy also decreases. In contrast, our proof approach first establishes that our method is equivalent to the single-agent SAC algorithm and then completes the proof using existing conclusions.
HASAC [16] is a CTCE version of MASAC, with its core contribution being the introduction of sequential decision-making. Therefore, we do not compare it here.
PAC [38] utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
Appendix I MPE setup
In the MPE [18] Spread scenario, three agents and three landmarks are present. Agents work together to cover all landmarks while avoiding collisions. We created three different maps: Spread(Simple), which matches the original benchmark’s simple_spread; Spread(Medium) and Spread(Hard), which increase the initial distances between agents and landmarks by 1.5X and 2X, respectively.
Appendix J Hyper-parameters
For all baseline algorithms, we implemented them using the corresponding open-source frameworks and chose the hyperparameters provided by these frameworks for replication. We used MAPPO implemented by  Huang et al., QMIX by  Hu et al., FOP by  Zhang et al., and HARL by  Zhong et al.. Our method, Soft-QMIX, shares the same code base as QMIX. The experiments were conducted on a computer equipped with 92 GB of RAM, a 40-core CPU, and a GeForce RTX 2080 Ti GPU. The following are the hyperparameters used in the experiments.
Hyperparameter Value
Learning Rate Actor 5e-4
Learning Rate Critic 1e-3
Data Chunk Length 8
Env Num 8
Episode Length 400
PPO Epoch 5
Actor Train Interval Step 1
Use Recurrent Policy True
Use Adv Normalize True
Use Value Active Masks False
Use Linear LR Decay True
Table 6: MAPPO hyper-parameters used for SMACv2. We utilize the hyperparameters used in SMACv2 [3].
Parameter Name Value
n_rollout_threads 8
num_env_steps 10000000
warmup_steps 10000
train_interval 50
update_per_train 1
use_valuenorm False
use_linear_lr_decay False
use_proper_time_limits True
hidden_sizes [256, 256]
activation_func relu
use_feature_normalization True
final_activation_func tanh
initialization_method orthogonal
gain 0.01
lr 0.0003
critic_lr 0.0005
[…]
$\gamma$ 0.99
buffer_size 1000000
batch_size 1000
polyak 0.005
n_step 20
use_huber_loss False
use_policy_active_masks True
share_param False
fixed_order False
Table 7: HASAC hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper [16].
Parameter Name Value
Action Selector epsilon greedy
$\epsilon$-Start 1.0
$\epsilon$-Finish 0.05
$\epsilon$-Anneal Time 1e5 (5e5 for MPE)
Runner parallel
Batch Size Run 4
Buffer Size 5000
Batch Size 128
Optimizer Adam
$t_{\max}$ 1005000
Target Update Interval 200
Mac n_mac
Agent n_rnn
Agent Output Type q
Learner nq_learner
Mixer qmix
Mixing Embed Dimension 32
Hypernet Embed Dimension 64
Learning Rate 0.001
$\lambda$ 0.4
Table 8: QMIX hyper-parameters used for SMACv2 and MPE. We utilize the hyperparameters used in SMACv2 [3].
Parameter Name Value
$\alpha$($\_5$ vs _ 5) 0.03
$\alpha$($\_10$ vs _ 10) 0.005
Table 9: Soft-QMIX hyper-parameters used for SMACv2. Note that besides the hyperparameters mentioned in the table, all other values for Soft-QMIX are the same as those for QMIX.
Parameter Name Value
Action Selector Multinomial
$\epsilon$-Start 1.0
$\epsilon$-Finish 0.05
$\epsilon$-Anneal Time 50000
Buffer Size 5000
$t_{\max}$ 1005000
Target Update Interval 200
Agent Output Type pi_logits
Learner fop_learner
Head Number 4
Mixing Embed Dimension 32
Learning Rate 0.0005
Burn In Period 100
$\lambda$ 0.4
Table 10: FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper [36].
Appendix K Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@
-Soft-QMIX: Integrating Maximum Entropy For Monotonic Value Function Factorization
+ME-IGM: Individual-Global-Max in Maximum Entropy Multi-Agent Reinforcement Learning
 Abstract
-Multi-agent reinforcement learning (MARL) tasks often utilize a centralized training with decentralized execution (CTDE) framework.
-QMIX is a successful CTDE method that learns a credit assignment function to derive local value functions from a global value function, defining a deterministic local policy.
-However, QMIX is hindered by its poor exploration strategy.
-While maximum entropy reinforcement learning (RL) promotes better exploration through stochastic policies, QMIX’s process of credit assignment conflicts with the maximum entropy objective and the decentralized execution requirement, making it unsuitable for maximum entropy RL.
-In this paper, we propose an enhancement to QMIX by incorporating an additional local Q-value learning method within the maximum entropy RL framework.
-Our approach constrains the local Q-value estimates to maintain the correct ordering of all actions.
-Due to the monotonicity of the QMIX value function, these updates ensure that locally optimal actions align with globally optimal actions.
-We theoretically prove the monotonic improvement and convergence of our method to an optimal solution.
-Experimentally, we validate our algorithm in matrix games, Multi-Agent Particle Environment and demonstrate state-of-the-art performance in SMAC-v2.
-[^1] Our code: https://github.com/WentseChen/Soft-QMIX
-Figure 1: Soft-QMIX combines the QMIX value decomposition mechanism with maximum entropy RL under the CTDE framework.
-The figure explains the algorithm design rationale.
-The left figure illustrates a naive approach to applying maximum entropy RL in the CTDE context, where blue text represents the desired objectives, and black text indicates corresponding constraints, It reveals that incorporating maximum entropy RL under the CTDE framework implicitly constrains the global value to be the sum of local values, significantly limiting the expressiveness of the critic network.
-The right figure depicts the Soft-QMIX improvement, where Soft-QMIX first employs QMIX’s value decomposition mechanism to obtain $Q_{i}$.
-Given the meaningful order of QMIX’s output $Q_{i}$, such as possessing properties like IGM, an order-preserving transformation converts $Q_{i}$ to $f_{i}(Q_{i})$.
-This transformation is trained by the loss function that minimize the difference between $f_{i}(Q_{i})$ and $Q_{tot}.$
+Multi-agent credit assignment is a fundamental challenge for cooperative multi-agent reinforcement learning (MARL), where a team of agents learn from shared reward signals.
+The Individual-Global-Max (IGM) condition is a widely used principle for multi-agent credit assignment, requiring that the joint action determined by individual Q-functions maximizes the global Q-value.
+Meanwhile, the principle of maximum entropy has been leveraged to enhance exploration in MARL.
+However, we identify a critical limitation in existing maximum entropy MARL methods: a misalignment arises between local policies and the joint policy that maximizes the global Q-value, leading to violations of the IGM condition.
+To address this misalignment, we propose an order-preserving transformation.
+Building on it, we introduce ME-IGM, a novel maximum entropy MARL algorithm compatible with any credit assignment mechanism that satisfies the IGM condition while enjoying the benefits of maximum entropy exploration.
+We empirically evaluate two variants of ME-IGM: ME-QMIX and ME-QPLEX, in non-monotonic matrix games, and demonstrate their state-of-the-art performance across 17 scenarios in SMAC-v2 and Overcooked.
 1 Introduction
-In the context of solving cooperative tasks using deep multi-agent reinforcement learning (MARL), a standard approach is centralized training with decentralized execution (CTDE) [5, 23, 15].
-Decentralized execution implies that during the execution phase, the decision-making of each agent is independent and limited to conditioning on local observations.
-Centralized training, on the other hand, allows the algorithm to access joint actions and global states while training the value function.
-Compared to decentralized training, which treats other agents as part of the environment, centralized training helps alleviate the instability often encountered in MARL training [30].
-In most cooperative tasks, the multi-agent system receives a single joint reward, and MARL algorithms learn the corresponding joint value function based on this reward.
-Consequently, there is a need to design a credit assignment mechanism [2] that decomposes the joint value function into several local value functions, guiding the updates of the actor networks through these local value functions.
-The challenge in credit assignment lies in reconstructing the joint distribution from the combination of multiple marginal distributions, a process that typically involves approximation errors [32].
-Therefore, past research has primarily focused on exploring the impact of mechanism design on the model’s capacity [27, 31].
-Beyond model capabilities, exploration ability is also a critical factor affecting the performance of MARL algorithms.
-Maximum entropy RL [14], known for encouraging exploration [7], smoothing the landscape of the objective [1], and enhancing robustness [4], has been widely adopted in single-agent RL.
-However, integrating MARL with maximum entropy RL is not straightforward.
-As illustrated in Figure 1, using maximum entropy RL within the framework of CTDE will significantly limit the model’s abilities in a manner akin to VDN [28], which enforces that the sum of local q-values must equal the global q-value.
-In this paper, we propose a value-based maximum entropy MARL method compatible with the CTDE framework, named Soft-QMIX.
-We adopt the QMIX [23] mechanism to train a centralized critic network and utilize its credit assignment mechanism to obtain local Q-functions.
-We also incorporate maximum entropy RL, utilizing stochastic policies and entropy regularization to improve exploration and robustness.
-To guarantee monotonically increasing expected returns during training and convergence to the optimal policy, an order-preserving transformation on the original local Q-functions is used for the final policy.
-Through this design, Soft-QMIX can benefit from both maximum entropy RL and QMIX’s credit assignment mechanism, while still ensuring convergence to the optimal strategy.
-Our main contributions can be summarized as follows:
-  • We introduce Soft-QMIX, a value-based maximum entropy MARL algorithm that aligns with the CTDE framework.
-It employs an order-preserving transformation on local q-values, enabling it to derive meaningful order from credit assignment and leverage the exploratory benefits provided by maximum entropy RL.
-  • Theoretically, we demonstrate that our approach can monotonically improve the expected q-value and guarantee convergence to the optimal policy.
-  • Empirically, we confirm the effectiveness of our algorithm through tests on matrix games, MPE [18] and SMAC-v2 [3], where our method attains state-of-the-art results.
-2 Related Works
-Multi-Agent RL.
-In the CTDE framework, MARL primarily involves two common algorithmic approaches: policy gradient methods [17, 34, 15, 5, 13, 37] and value function decomposition methods [28, 23, 31].
-Policy gradient methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence.
-On the other hand, Value function decomposition addresses the credit assignment problem by decomposing the global value function into multiple local value functions.
-VDN [28] assumes that the global Q-function is the sum of local Q-functions.
-QMIX [23, 20] permits the mixer function to be any function with non-negative weights.
-QTRAN [27] optimizes an additional pair of inequality constraints to construct a loss function.
-QPLEX [31] use the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize.
-The expressive power of the critic networks in these algorithms increases sequentially, but experimental results show that QMIX has better performance [9].
-Therefore, in this work, QMIX is used as the mechanism for credit assignment.
-Maximum Entropy RL.
-Maximum entropy RL [14] is demonstrated to have advantages such as encouraging exploration [7] and increasing robustness [4] in single-agent RL scenarios.
-In maximum entropy MARL, most works utilize an actor-critic architecture and typically aim to minimize the KL divergence between the local actor and the centralized critic [8, 6, 22, 36].
-However, although these methods employ QMIX to train the critic, they are only capable of minimizing the discrepancy in expectation, thus cannot guarantee the IGM condition, which is central to the QMIX algorithm.
-Our approach is value-based and ensures that the IGM condition is met while employing maximum entropy MARL.
-Moreover, our method introduces global information into the local policy during training through a hyper-network, which most previous methods either ignored [35] or only applied in centralized training and centralized execution scenarios [16].
-In appendix H, we provide a detailed comparison of our work with related works.
-3 Preliminary
+Figure 1: The figure illustrates the improvement of our approach compared to existing maximum entropy MARL methods.
+The left figure shows a straightforward approach to applying maximum entropy MARL in the CTDE context, where blue texts represent the desired objectives, and black texts indicate corresponding constraints.
+It reveals that existing maximum entropy MARL methods under the CTDE framework implicitly constrain the global Q-value to be the sum of local Q-values (as in VDN), significantly limiting the expressiveness of the critic network.
+The right figure depicts the improvements in ME-IGM.
+ME-IGM first applies any credit assignment mechanism that satisfies the IGM condition, such as QMIX and QPLEX, to obtain $Q_{i}$ for each agent $i$.
+Given the meaningful order of local Q-values (for different actions), an order-preserving transformation $f_{i}$ converts $Q_{i}$ to $f_{i}(Q_{i})$ as the policy logits.
+This transformation is optimized using a loss function that minimizes the expected difference between $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$, which guarantees monotonic policy improvement in maximum entropy MARL.
+Collaborative multi-agent tasks, where a team of agents works together to complete a task and receives joint rewards, are inherently challenging for reinforcement learning (RL) agents.
+This difficulty arises due to the absence of individual reward signals for each agent and the significantly higher exploration requirements compared to single-agent scenarios.
+The challenge is further exacerbated by the exponential growth of the joint action space as the number of agents increases.
+In this work, we propose leveraging maximum entropy RL to promote exploration and adopting the individual-global-max (IGM) credit assignment mechanism (Rashid et al.
+2020) to effectively distribute joint rewards among the agents.
+Maximum entropy RL (Levine 2018), recognized for promoting exploration (Haarnoja et al.
+2018), smoothing the objective landscape (Ahmed et al.
+2019), and improving robustness (Eysenbach and Levine 2021), has been extensively applied in single-agent RL.
+However, as illustrated in Figure 1, naively extending maximum entropy RL to multi-agent settings under the centralized training and decentralized execution (CTDE) paradigm leads to uniform credit assignment, which limits the ability of multi-agent reinforcement learning (MARL) agents to learn from joint rewards.
+On the other hand, numerous studies have explored the development of effective credit assignment mechanisms.
+Among them, the IGM condition has demonstrated strong empirical performance by aligning the maximum local Q-values with the maximum global Q-value.
+Previous works have primarily applied the IGM condition in deterministic policy settings, where policies are derived using the argmax of local Q-values.
+However, in maximum entropy MARL methods, stochastic local policies are learned, and misalignment can occur between local policies and the maximum global Q-values.
+That is, selecting joint actions based on the highest logits of local policies does not necessarily lead to maximizing the global Q-value.
+This limitation highlights a critical challenge in achieving effective credit assignment within the maximum entropy MARL framework.
+This challenge, as we will demonstrate, is not unique to IGM-based methods but also represents a broader obstacle for actor-critic-based maximum entropy MARL algorithms, such as FOP (Zhang et al.
+2021), hindering their ability to achieve optimal coordination.
+Our main contributions are summarized as follows:
+  • We are the first to identify and empirically demonstrate a critical issue of existing maximum entropy MARL methods: the misalignment between stochastic local policies and the maximum global Q-values.
+  • We introduce ME-IGM, the first CTDE maximum entropy MARL method that fully complies with the IGM condition.
+The core of ME-IGM is an order-preserving transformation (OPT), which maps local Q-values to policy logits while preserving their relative order.
+This ensures that selecting actions with the highest logits collectively leads to a joint action that maximizes the global Q-value.
+OPT can be seamlessly integrated into mainstream value decomposition MARL methods for a maximum entropy extension to enhance exploration.
+Additionally, we propose a theoretically grounded and straightforward objective for training these transformation operators.
+  • We confirm the effectiveness of our algorithm through extensive evaluation on matrix games, SMAC-v2 (Ellis et al.
+2022), and Overcooked (Carroll et al.
+2019), where our method attains state-of-the-art results.
+2 Background
 2.1 Multi-Agent Reinforcement Learning
-Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [19].
+Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek and Amato 2016).
 Formally, a Dec-POMDP is represented as a tuple $(\mathcal{A},S,U,T,r,O,G,\gamma)$, where $\mathcal{A}\equiv\{1,...,n\}$ is the set of $n$ agents.
@@
 For simplicity, if a function depends on both $o^{i}$ and $s$, we will disregard $o^{i}$.
-Each agent uses a policy $\pi_{i}(u^{i}|o^{i})$ to produce its action $u^{i}$ from its local observation $o^{i}$.
-Note that the policy should be conditioned on the history $o^{i}_{1:t}$.
+Each agent uses a policy $\pi_{i}(u^{i}|o^{i})$ to produce its action $u^{i}$.
+Note that the policy should be conditioned on the observation history $o^{i}_{1:t}$.
 For simplicity, we refer to $o^{i}_{1:t}$ as $o^{i}_{t}$ or $o^{i}$ in the whole paper.
@@
 $\rho_{\pi}(s_{t},\mathbf{u_{t}})$ is used to represent the state-action marginals of the trajectory distribution induced by a policy $\pi$.
-3.2 Individual-Global Condition
-In this section, we define two key concepts in credit assignment.
+2.2 Individual-Global-Max Condition
+Individual-Global-Max (IGM) (Rashid et al. 2020) is a commonly used credit assignment method in value-decomposition-based cooperative MARL and is defined as follows:
 Definition 2.1.
@@
 Then, we say that $[Q_{i}]_{i=1}^{n}$ satisfy IGM for $Q_{tot}$ under $s_{t}$.
-In QMIX, to ensure the validity of IGM, the algorithm constrains the expressive capability of $Q_{tot}$ to the following form: $\frac{\partial Q_{tot}}{\partial Q_{i}}\geq 0,\forall i\in\mathcal{A}.$ Restricting the capacity of $Q_{tot}$ using the above inequality results in $Q_{tot}$ being a monotonic function.
-The monotonicity constraint serves as a sufficient, but not necessary, condition for IGM.
-Consequently, the monotonicity constraint is more stringent than those required by IGM.
-Definition 3.2.
-Individual-Global-Independence (IGI) For joint policy $\pi_{jt}$, if there exist individual policies $\left[\pi_{i}\right]_{i=1}^{n}$ such that the following holds:
-$\pi_{jt}(\mathbf{u_{t}}|s_{t})=\prod_{i=1}^{n}\pi_{i}(u^{i}_{t}|o^{i}_{t}).$ (2)
-Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGI for $\pi_{jt}$ under $s_{t}$.
-Definition 3.1 characterizes the properties we expect a deterministic policy to possess.
-Definition 3.2, on the other hand, describes the independence of decision-making among agents during decentralized execution.
-IGI and the IGO [36] condition are very similar; we show in Appendix B that IGI is a sufficient condition for IGO.
-3.3 Naive Maximum Entropy MARL Algorithm
-In this section, we informally introduce a naive MARL algorithm.
-We use it to illustrate the issue that naively applying credit assignment to the maximum entropy RL algorithm will lead to limited credit assignment capability.
-Assume a value-based framework, where $Q_{i}$ denotes the local Q-function and the function for credit assignment is denoted as $f^{c}$.
-Hence, the global Q-function, $Q_{tot}$, can be expressed as $Q_{tot}(\mathbf{u},s)=f^{c}\big{(}Q_{1}(u^{i},o^{i}),...,Q_{n}(u^{n},o^{n}),s\big{)}$.
-Based on maximum entropy RL, the global policy is defined as $\pi_{jt}(\cdot|s)\propto\text{softmax}\big{(}Q_{tot}(\cdot,s)\big{)}$ and the local policy as $\pi_{i}(\cdot|o^{i})\propto\text{softmax}\big{(}f^{t}_{i}(Q_{i}(\cdot,o^{i}),s)\big{)}$, where $f^{t}_{i}$ is an arbitrary function.
-Due to the IGI condition, we have $f^{c}(Q_{1},...,Q_{n},s)=\sum_{i=1}^{n}f^{t}_{i}(Q_{i},s)$.
-When $f^{t}_{i}$ is an identity mapping, the aforementioned algorithm is equivalent to VDN [28].
-When $f^{t}_{i}$ represents a linear transformation, the algorithm becomes analogous to the DOP, an actor-critic method [33].
-In fact, $f^{t}_{i}$ can be any function of $u^{i}$ and $o^{i}$, but cannot be a function of $u^{j}$ or $o^{j}$, where $j\neq i$.
-This constraint limits the expressiveness of the critic network.
-4 Soft-QMIX
-In this paragraph, we first briefly introduce the motivation behind the Soft-QMIX algorithm.
-Then, we detail the process of the Soft-QMIX algorithm and show that the expected returns of the centralized training are monotonically increasing and converge to the optimum.
-Finally, we propose a practical algorithm.
-4.1 Motivation
-As shown in Figure 2, we divide the decision-making process into two stages (green and blue blocks).
-In the first stage, agents rank the Q-values for all actions.
-We employ the value decomposition mechanism from QMIX to obtain local Q-values but only utilize the order of their outputs, not the specific Q-function values.
-In the second stage, without changing the order, we assign specific Q-values corresponding to each action.
-When selecting actions in decentralized execution, only the maximum value is required, necessitating just the first stage for execution.
-The second stage serves primarily to obtain an accurate estimated q-value.
-During this phase, we can utilize global information, as it is only used during training.
-4.2 Soft Policy Iteration
-The objective function of Maximum Entropy RL can be represented in the following form:
-$J(\pi_{jt})$ $=\mathbb{E}_{(s_{t},\mathbf{u}_{t})\sim\rho_{\pi_{jt}}}\big{[}r(s_{0},\mathbf{u}_{0})+\sum_{t=1}^{\infty}\gamma^{t}\big{(}r(s_{t},\mathbf{u}_{t})+\alpha H\left(\pi_{jt}(\cdot|s_{t})\big{)}\right)\big{]},$ (3)
-where $H(\pi_{jt})$ denotes the entropy of the joint policy $\pi_{jt}$, and the temperature parameter $\alpha$ is used to control the degree of randomness in the policy.
-In the policy evaluation step, we aim to approximate the objective function described in Equation 3 using $Q_{tot}:S\times\mathbf{U}\rightarrow\mathbb{R}$.
-The soft state value function $V_{tot}:S\rightarrow\mathbb{R}$ is defined as follows:
-$V_{tot}(s)=\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[Q_{tot}(s,\mathbf{u})-\alpha\log\pi_{jt}(\mathbf{u}|s)].$ (4)
-Repeatedly applying the Bellman operation to update $Q_{tot}$ results in the soft $Q_{tot}$ function for policy $\pi_{jt}$. The Bellman operation is defined as follows:
-$\mathcal{T}^{\pi_{jt}}Q_{tot}(s,\mathbf{u})=r(s,\mathbf{u})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}}[V_{tot}(s^{\prime})].$ (5)
-In reference to the QMIX, a monotonicity constraint is imposed on the capacity of the critic network. Specifically, $Q_{\text{tot}}$ can be formulated as follows:
-$Q_{tot}(s,\mathbf{u})=\text{Mixer}\left(Q_{1}(o^{1},u^{1}),...,Q_{n}(o^{n},u^{n}),s\right).$ (6)
-The network structure of the Mixer function is the same as QMIX’s, and is elaborated in Appendix G.
-Due to the monotonicity constraint, there is moderate error in estimating of $Q_{tot}$.
-To enhance the precision of value estimation, we introduce the TD($\lambda$[24] method.
-The estimation error between the maximum entropy objective function and the Bellman equation objective can be expressed as follows:
-$\delta_{t}^{V}$ $=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1}),$ (7)
-$\epsilon_{t}^{td(\lambda)}$ $=\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u_{t}})-Q_{tot}(s_{t},\mathbf{u_{t}})=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{V}_{t+l}.$
-Figure 2: The overall network architecture.
-Functions $g_{i}$ and $f_{i}$ are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values.
-The output of the $f_{i}$ function, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled.
-For decentralized execution, only the green network is needed, and the maximum values of their output are selected.
-The derivation and implementation details of Equation 7 are provided in Appendix D.
-When combining off-policy methods with TD($\lambda$), adopting $\lambda>0$ requires importance sampling corrections [10].
-Following previous work [9], we opt for a small $\lambda$ and assume that $\mathcal{T}^{\pi_{jt}}_{\lambda}$ is equivalent to $\mathcal{T}^{\pi_{jt}}$ in expectation.
-For simpilicity, we will use $\mathcal{T}^{\pi_{jt}}$ instead of $\mathcal{T}^{\pi_{jt}}_{\lambda}$ in theoretical discussions.
-The following lemma demonstrates that by iteratively applying the Bellman equation, convergence can be achieved towards the maximum entropy objective $Q_{\text{tot}}$ for $\pi_{jt}$.
-Lemma 4.1 (Joint Soft Policy Evaluation).
-Let $Q^{0}_{jt}:S\times\mathcal{U}\to\mathbb{R}$ be a bounded mapping where $|\mathcal{U}|<\infty$.
-Define the iterative process $Q^{k+1}_{tot}=\mathcal{T}^{\pi_{jt}}(Q^{k}_{tot})$.
-It follows that the sequence $\{Q^{k}_{tot}\}_{k=0}^{\infty}$ converges to the joint soft-Q-function for policy $\pi_{jt}$ as $k\rightarrow\infty$.
-Proof. See Appendix C. ∎
-In Section 4.1, it was mentioned that decision-making is divided into two stages.
-During training, the policy is defined as $\pi^{train}_{i}(\cdot|o^{i},s)=\text{softmax}(f^{\text{stage2}}_{i}(f^{\text{stage1}}_{i}(\cdot|o^{i}),s)/\alpha)$, where $f^{\text{stage1}}_{i}$ is a function dependent on local observations, responsible for the first stage of decision-making.
-$f^{\text{stage2}}_{i}$, on the other hand, is a non-linear order-preserving transformation, responsible for the second stage of decision-making.
-Although $\pi^{train}_{i}$ depends on the global state, which conflicts with the definition in Dec-POMDP, it remains a valid option because it is utilized solely during training.
-In the subsequent paragraph, $\pi$ will denote $\pi^{train}_{i}$.
-$f^{\text{stage2}}_{i}$ can be further decomposed into two order-preserving transformations, $g_{i}$ and $f_{i}$.
-The output of $g_{i}$ is $Q_{i}$, and $f_{i}$ further transforms $Q_{i}$ into the final policy.
-The form of order-preserving transformation (OPT) is as follows:
-$OPT_{i}(x,s)=W_{{}_{(1\times d_{1})}}^{2}\sigma(W^{1}_{{}_{(d_{1}\times 1)}}x_{{}_{(1\times d_{2})}}+b^{1})+b^{2},$ (8)
-where $d_{1}$ is the hidden dimension, $d_{2}$ is the dimension of actions, $b^{1}$ and $b^{2}$ are the bias, $\sigma$ is an activation function, and $W^{1}$ and $W^{2}$ are non-negative weights.
-The implementation of the order-preserving transformation is similar to QMIX’s and is elaborated in Appendix G, where its weights and biases are provided by a hypernetwork.
-This hypernetwork takes the global state $s$ as input, making $g_{i}$ and $f_{i}$ functions dependent on the global state.
-Figure 2 presents the overall framework, where Agents $1$ to Agents $n$ correspond to $f_{i}^{\text{stage1}}$, and $g_{i}$ and $f_{i}$ correspond to $f_{i}^{\text{stage2}}$.
-Since $g_{i}$ is an order-preserving transformation, and $[Q_{i}]_{i=1}^{n}$ satisfy the IGM condition for $Q_{tot}$, it follows that $\underset{\mathbf{u}}{\text{argmax}}\,Q_{tot}(\mathbf{u},s)=\left(\underset{u^{1}}{\text{argmax}}\,f_{1}^{\text{stage1}}(u^{1}|o^{1}),\ldots,\underset{u^{n}}{\text{argmax}}\,f_{n}^{\text{stage1}}(u^{n}|o^{n})\right)$.
-Therefore, during execution, only $f_{i}^{\text{stage1}}$ is needed.
-During policy improvement, we update the joint policy towards the exponential of the joint Q-function.
-We constrain $\pi\in\Pi$, where $\Pi$ is the family of functions that satisfy the IGI condition.
-In the updating process, the updated policy is projected onto $\Pi$.
-Specifically, the update formula is as follows:
-$\pi_{jt}^{\text{new}}=\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{arg\,min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (9)
-[…]
-In value-based methods, Equation 9 is intractable. Therefore, in the policy improvement step, we update the policy in this way:
-$f_{i}^{\text{new}}=\underset{f^{\prime}_{i}}{\mathrm{argmin}}\frac{1}{2}\left(\sum_{i}f^{\prime}_{i}(Q_{i}(o^{i}_{t},u^{i}_{t},s_{t}))-y_{t}\right)^{2},$ (10)
-where $y_{t}$ can be any unbiased estimate of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})\right]$.
-Here, we assume [^2] This does not necessarily hold in practical implementations, but the difference between these is a term in the loss function of our algorithm.
-For further discussion, please see the Appendix F.
-that $\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u_{t}})$ is an unbiased estimate of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}\sum_{i}f_{i}(Q_{i}(s_{t},\mathbf{u_{t}}))\right]$.
-Under this assumption, we can replace $Q_{\theta}$ and $y_{t}$ in Theorem 4.2 with $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$ respectively, and get that minimizing the L2 norm described in Equation 10 is equivalent to minimizing the KL divergence described in Equation 9 and minimizing the TD loss of value function at the same time.
-Therefore, in value-based methods, Equation 10 can be used in place of Equation 9.
+2.3 Related Works
+Multi-Agent Credit Assignment: CTDE MARL primarily consists of two algorithmic branches: actor-critic methods (Lowe et al.
+2017; Yu et al.
+2022; Lin et al.
+2023; Foerster et al.
+2018; Kuba et al.
+2021; Zhong et al.
+2023) and value decomposition methods (Sunehag et al.
+2017; Rashid et al.
+2020; Wang et al.
+2020).
+Actor-critic methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence between local policies and the softmax of the global Q-value.
+On the other hand, value decomposition methods address the credit assignment problem by decomposing the global value function into multiple local value functions while satisfying the IGM condition.
+In particular, VDN (Sunehag et al.
+2017) assumes that the global Q-function is the sum of local Q-functions.
+QMIX (Rashid et al.
+2020) allows the mixing function, which aggregates local Q-values into the global Q-value, to be any function with non-negative partial derivatives, i.e., $\frac{\partial Q_{tot}}{\partial Q_{i}}\geq 0,\forall i\in\mathcal{A}.$ QTRAN (Son et al.
+2019) introduces an additional set of inequality constraints to construct its loss function.
+QPLEX (Wang et al.
+2020) uses the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize.
+Our method, ME-IGM, enhances the exploration capability of these MARL algorithms by leveraging the maximum entropy RL framework and is compatible with all the aforementioned value decomposition methods.
+Maximum Entropy MARL: Maximum entropy RL is demonstrated to encourage exploration in single-agent RL (Haarnoja et al.
+2018).
+Most maximum entropy MARL works utilize an actor-critic architecture, minimizing the KL divergence between the local actors and the softmax of the centralized critic (He et al.
+2022; Guo and Wu 2022; Zhang et al.
+2021).
+However, as demonstrated later, these methods cannot guarantee that the joint action (at each state) determined by local policies maximizes the global Q-value, thus violating the IGM condition.
+In this case, we propose a novel order preserving transformation to ensure that the IGM condition is met while employing maximum entropy MARL.
+3 Misalignment Between Local Policies and the Maximum Global Q-Value
+The Misalignment Problem of Maximum Entropy MARL: While the IGM condition defines the relationship between local Q-values and the global Q-value, what is often more important in practice is the alignment between local policies and the global Q-value.
+In other words, joint actions that maximize the global Q-value can be obtained by querying local policies.
+Previous work in maximum entropy MARL (Zhang et al.
+2021; Guo and Wu 2022) often decompose the global Q-value into individual local Q-values to satisfy the IGM condition.
+These local Q-values are then used to guide the training of local stochastic policies.
+Compared with value decomposition methods such as QMIX, they train stochastic policy networks in addition to the Q-functions, which can potentially improve multi-agent exploration in complex environments.
+However, while the value decomposition ensures alignment between global and local Q-values via the IGM condition, there is typically no explicit constraints imposed between the local policies and the local Q-values during the policy training phase.
+This lack of coupling often leads to misalignment during testing, where the local policies fail to consistently select joint actions that maximize the global Q-value, defeating the purpose of enforcing the IGM condition.
+Consequently, the learned policies may struggle to achieve optimal global coordination during execution.
+Figure 2: Illustration of misalignment between local policies and the maximum global Q-value.
+When naively combining the IGM condition with maximum entropy MARL, local policies often select suboptimal joint actions, leading to a lower Q-value, depicted by the blue curve.
+In contrast, the optimal joint action should achieve the global-max Q-value, shown as the orange curve.
+Empirical Demonstration: To analyze the performance gap caused by the misalignment between local policies and the maximum global Q-value, we evaluate the difference between the maximum global Q-value and the global Q-value selected by local policies in the SMAC-v2 zerg 5vs5 scenario, using a state-of-the-art Maximum Entropy MARL algorithm (Zhang et al.
+2021).
+Figure 2 shows the results, averaged over five random seeds with standard deviations included.
+We can observe that there is a significant Q-value gap, emphasizing the potential for improvement by developing algorithms that enable the selection of joint actions with the highest global Q-value through local stochastic policies.
+4 Order-Preserving Transformation
+Here, we propose a novel order-preserving transformation (OPT) to solve the aforementioned misalignment problem between local policies and the maximum global Q-value.
+4.1 Formulation of OPT
+OPT is used to map local Q-values to the logits of local policies while preserving their relative order. Formally, the OPT is defined as:
+$\text{OPT}(x,s)=W^{2}\sigma(W^{1}x+b^{1})+b^{2}$ (2)
+Here, $W^{1}_{{}_{(d_{1}\times 1)}}$, $W_{{}_{(1\times d_{1})}}^{2}$, $b^{1}_{{}_{(d_{1}\times 1)}}$, and $b^{2}_{{}_{(1\times 1)}}$ are generated by a hyper-network [^3] For details of the hyper-network, please refer to the Appendix.
+which takes the global state $s$ as input.
+Consider $x$ as a local Q-value of dimension $(1\times d_{2})$, then $W^{1}x+b^{1}$ is a set of hidden variables of size $(d_{1}\times d_{2})$.
+Notably, when all entries in $W^{1}$ are non-negative, the relative order of elements in $x$ is preserved in each of the $d_{1}$ hidden variables.
+Next, a nonlinear activation function $\sigma$, such as ELU (Clevert 2015), is applied, followed by another order-preserving linear transformation using $W^{2}$ and $b^{2}$.
+Consequently, the final output, $\text{OPT}(x,s)$, maintains both the dimension and the relative order of the entries in the input $x$.
+$\text{OPT}(x,s)$ is used as the logits of the local policy corresponding to the local Q-value $x$.
+Through aforementioned order-preserving design, each agent can select its action corresponding to the highest local policy logit, which is also the maximum point of its local Q-function.
+The local Q functions can be learned with any value decomposition methods satisfying the IGM condition, to ensure that the selected local actions collectively compose the optimal joint action that maximize the global Q-function.
+Thus, the proposed OPT addresses the aforementioned misalignment issue commonly found in maximum entropy MARL.
+4.2 Training Scheme of OPT
+Figure 3: The overall pipeline of ME-IGM.
+Here, we first formulate the ideal objective of Maximum Entropy MARL and its practical counterpart that admits an analytical-form solution.
+We then analyze the gap between these two objective functions, providing a theoretically sound objective design for Maximum Entropy MARL based on OPTs.
+Maximum Entropy MARL trains the individual and global Q-functions as in value decomposition MARL methods, while training a policy network in a Maximum Entropy RL framework based on the global Q-values.
+Formally, the joint policy is trained as follows:
+$\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (3)
+where $Z^{\pi_{\text{old}}}(s_{t})=\sum_{\mathbf{u}}\exp(Q_{tot}^{\pi_{\text{old}}}(s_{t},\mathbf{u})/\alpha)$ and $\alpha$ is the temperature in the softmax distribution.
+To simplify and decentralize the policy improvement process, we propose replacing $Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)$ with $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})$, which leads to:
+$\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(\sum_{i}\frac{\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})}{\alpha})}{Z^{\prime\pi_{\text{old}}}(s_{t})}\right),$ (4)
+where $Z^{\prime\pi_{\text{old}}}(s_{t})$ is the normalization factor.
+The joint policy $\pi^{\prime}_{jt}$ is implemented as a set of independent local policies for decentralized execution, and we have the analytical solution for the local policy of agent $i$ as:
+$\pi_{i}^{\text{new}}(a|o^{i}_{t})\propto\exp(\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},a),s_{t})/\alpha).$ (5)
+We can see that the policy is defined based on the local Q-value and utilizes the OPT to ensure that the relative order of policy logits aligns with the local Q-values.
+There is a gap between $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})$ and $Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)$.
+Also, we find that the policy improvement using Equation (4) no longer guarantees a monotonic increase in the global Q-value, unlike the approach based on Equation (3).
+Specifically, we have the following theorem:
 Theorem 4.1.
-If $Q$-function is parameterized as $Q_{\theta}$, and $V_{\theta},\pi_{\theta}$ are defined as below:
-$V_{\theta}(s)$ :-$\coloneq\alpha\log\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\exp\left(Q_{\theta}(s,\mathbf{u})/\alpha\right)\right],:$ (11)
-$\pi_{jt_{\theta}}(\mathbf{u}|s)$ :-$\coloneq\exp\left(Q_{\theta}(s,\mathbf{u})-V_{\theta}(s)/\alpha\right).:$
-and $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right].$ Then, we have
-$\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}$ $\left[\|Q_{\theta}(s,\mathbf{u})-y\|^{2}\right]$ (12)
-$=\alpha\nabla_{\theta}D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}+\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\|V_{\theta}(s)-y\|^{2}\right]$
-Proof. See Appendix E.
-The subsequent theorem ensures that updating the policy using Equation 9 results in a monotonic increase of the Q-function under the IGI condition.
-Lemma 4.3 (Joint Soft Policy Improvement).
-Suppose $\pi_{jt}^{\text{old}}\in\Pi$ and consider $\pi_{jt}^{\text{new}}$ to be the solution to Equation 9.
-Then $Q^{\pi_{jt}^{\text{new}}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})$ for all $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$ with $|\mathbf{U}|<\infty$.
-Proof. See Appendix C.
-Theorem 4.4 demonstrates that iteratively applying policy evaluation and policy improvement leads to a globally optimal policy under the IGI condition.
-Theorem 4.4 (Joint Soft Policy Iteration).
-By continually applying soft policy evaluation and soft policy improvement to any $\pi_{jt}\in\Pi$ converges to a policy $\pi^{*}$ such that $Q^{\pi^{*}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi}(s_{t},\mathbf{u}_{t})$ for all $\pi_{jt}\in\Pi$ and $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$, assuming $|\mathbf{U}|<\infty$.
-Proof. See Appendix C.
-4.3 The Soft-QMIX Algorithm
-Figure 2 illustrates the network architecture of the Soft-QMIX algorithm, which comprises four types of networks.
-The Mixer network is responsible for handling credit assignment and is parameterized along with its corresponding hypernetwork as $\text{Mixer}_{\theta}$.
-The $[g_{i}]_{i=1}^{n}$ and $[f_{i}]_{i=1}^{n}$ are $2n$ order-preserving transformations, with each $g_{i},f_{i}$ and their corresponding hypernetwork parameterized as $g_{i,\theta}$ and $f_{i,\theta}$.
-Agents 1 to $n$ share parameters, also referred to as $f_{i}^{\text{stage1}}$ in the previous section, and are parameterized as $f^{\text{stage1}}_{\theta}$.
-The updated rules for these parameters are described as follows.
-Soft-QMIX aims to minimize two loss functions. $\text{Mixer}_{\theta}$, $g_{i,\theta}$ and $f^{\text{stage1}}_{\theta}$ are updated using the following loss function:
-$L_{Q}(\theta)=\frac{1}{2}\left(Q_{tot_{\theta}}(s_{t},\mathbf{u}_{t})-\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2},$ (13)
-where $Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big{(}g_{1,\theta}(f^{\text{stage1}}_{\theta}(o^{1},u^{1}),s),\ldots,$ $g_{n_{\theta}}(f^{\text{stage1}}_{\theta}(o^{n},u^{n}),s),s\big{)}$, and $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u}_{t})$ is defined in Equation 7.
-We utilize the technique of target Q networks to stabilize training, where the target Q network is parameterized as $\bar{\theta}$, with $\bar{\theta}$ being the exponentially weighted moving average of $\theta$.
-Subsequently, the parameter $\phi$ is updated using the following formula:
-$L_{\pi}(\phi)=\frac{1}{2}\left(\sum_{i=1}^{n}f_{i,\phi}(Q_{i}(o^{i}_{t},u^{i}_{t}),s_{t})-Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2}.$ (14)
-In this process, the function $f_{i,\phi}$ takes $Q_{i}(o^{i}_{t},u^{i}_{t})$ as input, but during backpropagation, gradients are truncated at $Q_{i}$.
-This truncation ensures that this loss function updates only $f_{i,\phi}$.
-After the update, the policy can be expressed in the following form:
-$\pi_{i}(u^{i}|o^{i},s)=\text{softmax}(\frac{f_{i}\left(g_{i}\big{(}f^{stage1}(o^{i}),s\big{)},s\right)}{\alpha})$ (15)
-It’s important to note that during execution, we only need to call $f^{stage1}$ and apply the argmax operation to it.
-In this paper, $\alpha$ is fixed during training.
-Finally, the algorithmic procedure of Soft-QMIX is shown in appendix A.
-5 Experiments
-In this section, we evaluate our algorithm using matrix games from prior works [27], demonstrating its ability to consistently reach the global optimum.
-Then, we compare Soft-QMIX with QMIX on MPE [18] and empirically demonstrate Soft-QMIX algorithm achieves state-of-the-art performance on the SMAC-v2 [3] benchmark.
-We compare our algorithm with four baseline algorithms.
-MAPPO [34] adapts the on-policy, single-agent PPO [26] algorithm for multi-agent use, utilizing a centralized value function.
-QMIX [23] is a value-based method employing a monotonic value decomposition to create multiple local deterministic policies from a shared value function.
-It uses an $\epsilon$-greedy strategy for exploration.
-FOP [36] implements a maximum entropy policy gradient approach with value factorization similar to QPLEX.
-HASAC [16] offers an alternative multi-agent soft actor-critic framework, facilitating sequential decision-making by agents in a centralized training with centralized execution context.
-We elaborate implementation details and hyperparameters in Appendix J.
-5.1 Matrix Games
-In this section, we demonstrate the effectiveness of our algorithm in the classic one-step matrix games.
-Table 1 presents a classic non-monotonic payoff matrix, often used in past papers [27, 31] to discuss the representational capability of critic networks.
-QMIX, due to its monotonicity constraint, often incurs moderate errors in fitting such payoff matrices.
-However, in practice, during the training process, we are primarily concerned with the value estimates of those joint actions that are frequently visited.
-For suboptimal joint actions with lower Q values, their errors do not significantly impact the estimated expected returns.
-As we use TD-$\lambda$ to estimate returns in our experiments, such errors can be further reduced.
-Conversely, we are more concerned with whether the joint actions with the highest probability of being selected have the globally maximal value function, as we aim for the highest returns when executing with a deterministic policy.
-For suboptimal joint actions, we also wish to visit them with a certain probability to ensure adequate exploration.
-Table 1: Payoff matrix of the one-step matrix game.
+Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$.
+Updating the policy with Equation (3) ensures monotonic improvement in the global Q-value.
+That is, $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}.$
+Algorithm 1 ME-IGM
+1:Input: $\theta,\phi$
+2:Initiate: ${\theta^{-}}\leftarrow\theta,D\leftarrow\emptyset$
+3:for each iteration do
+4:  for each environment step do
+5:   get $\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t})$ via Eq. 10
+6:   $u^{i}_{t}\sim\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t}),\forall i\in\{1,..,n\}$
+7:   $s_{t+1}\sim\mathcal{P}(s_{t},\mathbf{u}_{t},\cdot)$
+8:   $D\leftarrow D\cup\{(s_{t},\mathbf{u}_{t},r(s_{t},\mathbf{u}_{t}),s_{t+1})\}$
+9:  end for
+10:  for each gradient step do
+11:   calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq. 12
+12:   $\theta\leftarrow\theta-\lambda_{\theta}\nabla_{\theta}L(\theta)$
+13:   $\phi\leftarrow\phi-\lambda_{\phi}\nabla_{\phi}L(\phi)$
+14:   $\omega\leftarrow\omega-\lambda_{\omega}\nabla_{\omega}L(\omega)$
+15:   ${\theta^{-}}\leftarrow\tau\theta+(1-\tau){\theta^{-}}$
+16:  end for
+17:end for
+In contrast, updating the policy with Equation (4) only ensures that $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})-\epsilon\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}$.
+The gap $\epsilon$ satisfies:
+$\epsilon\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim\mathcal{P}}\big{[}C\sqrt{2D_{KL}\big{(}\pi_{\text{old}}(\cdot|s_{t+1})|\pi_{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$ (6)
+where $\gamma$ is the discount factor, $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function and $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})\triangleq|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}),s_{t+1})|$.
+Proof is provided in the Appendix.
+This theorem implies that we must enforce $\epsilon\leq 0$ to guarantee monotonic policy improvement when using Equation (4) as the objective.
+This motivates us to minimize the factor $C$ and adopt the following objective to train the OPTs.
+$\min_{{\text{OPT}_{1:n}}}\mathbb{E}_{s,\mathbf{u}\sim\rho_{\pi_{jt}^{new}}}[\Delta_{Q}(s,\mathbf{u})^{2}],$ (7)
+$\pi_{jt}^{new}=(\pi_{1}^{new},\cdots,\pi_{n}^{new})$ and $\pi_{i}^{new}$ is defined in Eq. (5).
+(a) Payoff Matrix
 A B C
@@
 C -12 0 0
-In the matrix game shown in Table 2, QMIX exhibits high estimation error due to its monotonicity constraints, and its deterministic policy lacks effective exploration, which leads it to learn only suboptimal joint actions with a reward of 00.
-FOP, with its network structure similar to QPLEX, can accurately estimate value functions, yet the corresponding policy learned tends to select suboptimal joint actions more frequently.
-Soft-QMIX has smaller estimation errors for optimal joint actions and larger errors for suboptimal ones.
-However, the joint action with the maximum Q value in Soft-QMIX matches the true optimal joint action, and the local policy has the highest probability of selecting the optimal joint action.
-Soft-QMIX possesses this capability because its decision-making is divided into two steps, and in the first step, it ensures the order of probabilities of each action being selected in the policy.
-Table 2: (a)-(c) respectively demonstrate the results of Soft-QMIX, FOP, and QMIX in the one-step matrix game.
-The first row and column display the policies adopted by the row and column players after $10k$ steps of training, respectively.
-The 3$\times$3 matrix in the bottom right corner represents the $Q_{\text{tot}}$ ultimately learned by the algorithm.
-The results indicate that Soft-QMIX is the only method among all that consistently chooses the optimal action with the highest probability.
+(b) Result of ME-QMIX (Ours)
 1. 0. 0.
@@
 0. -10 -30 -32
-(a) result of Soft-QMIX
+(c) Result of FOP
 .21 .39 .40
@@
 .39 -12 0 0
-(b) result of FOP
-$\epsilon$ 1-$2\epsilon$ $\epsilon$
+(d) Result of QMIX
+[…]
 $\epsilon$8.5 -8.5 -8.5
-1-$2\epsilon$8.5 0.2 0.2
+[…]
 $\epsilon$8.5 0.2 0.2
-(c) result of QMIX
-5.2 Multi-Agent Particle Environment
-In this section, we test our algorithm on three custumed Multi-Agent Particle Environment(MPE) [18] scenarios and show that Soft-QMIX do better exploration compare to QMIX and ablate through different choice of $\alpha$.
-Specifically, we use three different Spread map, where in each map, agents are tasked to navigate to all the landmarks without collide with each other.
-Harder map means longer initial distance between agents and landmarks.
-Details are available in Appendix I.
-Table 3 shows that Soft-QMIX significantly outperforms QMIX on the Hard and Medium map which require more exploration, but shows only a slight advantage on the Simple map.
-Additionally, we find that higher $\alpha$ values enhance exploration and performance.
-These findings support our methodological claim that integrating maximum entropy RL with MARL algorithms boosts exploration capabilities.
-For each scenario, we conduct four experiments across five seeds, reporting the average evaluation episode rewards after 10M environment steps and their standard deviation in our results table.
-Table 3: Experiment result in three MPE scenarios, where Soft-QMIX has higher performance gain from QMIX on Hard and Medium map, which need more exploration.
-Higher $\alpha$ also lead to better exploration and higher performance gain.
-eval episode QMIX Soft-QMIX(ours)
-rewards $\alpha=0.1$ $\alpha=0.3$ $\alpha=1.0$
-Spread(Hard) -171.7 $\pm$ 3.8 -173.7 $\pm$ 3.0 -170.8 $\pm$ 3.1 -168.7 $\pm$ 4.6 (+1.8%)
-Spread(Medium) -138.4 $\pm$ 2.9 -139.0 $\pm$ 2.0 -137.4 $\pm$ 3.3 -134.9 $\pm$ 1.8 (+2.5%)
-Spread(Simple) -111.9 $\pm$ 1.6 -116.2 $\pm$ 2.7 -114.4 $\pm$ 2.0 -111.4 $\pm$ 0.9 (+0.4%)
-5.3 SMAC-v2
-We evaluated Soft-QMIX on the SMACv2 [3] benchmark, including six scenarios: protoss_5_vs_5, terran_5_vs_5, zerg_5_vs_5, protoss_10_vs_10, terran_10_vs_10, and zerg_10_vs_10.
-The results, as shown in Figure 3, were obtained by training each scenario for $10$ million environmental steps, with each curve representing the average of 5 seeds, and HASAC with an average of 3 seeds.
-Soft-QMIX outperforms the baseline algorithms in all scenarios.
-MAPPO, being an on-policy algorithm, has lower sampling efficiency, resulting in lower win rates for models trained under 10M environmental steps.
-Training with MAPPO for 60M environmental steps can achieve win rates comparable to 10M QMIX.
-QMIX, lacking exploration, tends to converge to local optima, especially in the two Zerg scenarios.
-HASAC and FOP perform poorly in all the scenarios.
-Ultimately, Soft-QMIX improves win rates by $5\%$ to $24\%$ in all scenarios compared to baseline algorithms.
-Figure 3: Experiment result in six SMAC-v2 scenarios, where Soft-QMIX achieved faster convergence speed and higher performance compared to baseline algorithms.
+[…]
+5 The Overall Framework: ME-IGM
+Figure 3 illustrates the framework of ME-IGM, which consists of three types of networks.
+Agents 1 to $n$ share a network $f_{\theta}$ for predicting local Q-values.
+The Mixer network, which aggregates local Q-values and manages the credit assignment among agents, is parameterized through its associated hyper-network as $\text{Mixer}_{\theta}$ [^4] Our algorithm is compatible with any Mixer network design that satisfies the IGM condition.
+For instance, during evaluation, we utilize the Mixer networks of QMIX and QPLEX, leading to two variants of our algorithm: ME-QMIX and ME-QPLEX..
+$[\text{OPT}_{i,\phi}]_{i=1}^{n}$ are $n$ order-preserving transformations, and $\text{OPT}_{i,\phi}(f_{\theta}(o_{t}^{i}),s_{t})$ defines a local stochastic policy for agent $i$.
+The update rules for these networks are described as follows.
+ME-IGM optimizes two loss functions to update $\theta$ and $\phi$ separately. $\text{Mixer}_{\theta}$ and $f_{\theta}$ are trained by minimizing $L(\theta)$:
+$\mathbb{E}_{\tau_{t}}\left[\frac{1}{2}\left(Q_{tot_{\theta}}(s_{t},\mathbf{u}_{t})-\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2}\right]$ (8)
+where $\tau_{t}=\{o_{t:t+1}^{1:n},u_{t:t+1}^{1:n},s_{t:t+1},r_{t}\}$, $Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big{(}f_{\theta}(o^{1},u^{1}),\ldots,f_{\theta}(o^{n},u^{n}),s\big{)}$, and the estimated Q-target $\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})=r_{t}+\gamma Q_{tot,\theta^{-}}(s_{t+1},\mathbf{u}_{t+1})$.
+As a common practice, we utilize a target Q network $Q_{tot,\theta^{-}}$ to stabilize training, with $\theta^{-}$ being the exponentially weighted moving average of $\theta$.
+Notably, local actions $u_{t:t+1}^{1:n}$ are sampled from local stochastic policies $\pi_{i}$ (as defined in Equation (10)) during roll-out, rather than selecting the argmax of local Q-functions.
+This offers two advantages: (1) the softmax policy function enhances exploration, and (2) incorporating state $s$ in the sampling process allows for better utilization of global information during centralized training.
+Further, to train $[\text{OPT}_{i,\phi}]_{i=1}^{n}$, we adopt a sample-based version of Equation (7), denoted as $L(\phi)$:
+$\mathbb{E}_{\tau_{t}}\left[\left(\sum_{i=1}^{n}\text{OPT}_{i,\phi}(f_{\theta}(o^{i}_{t},u^{i}_{t}),s_{t})-Q_{tot,\theta}(s_{t},\mathbf{u}_{t})\right)^{2}\right]$ (9)
+OPTs define the local policies for each agent as follows:
+$\pi_{i}(u^{i}|o^{i},s)=\text{softmax}\left(\text{OPT}_{i,\phi}\left(f_{\theta}(o^{i}),s\right)/\alpha_{\omega}\right)$ (10)
+$\alpha_{\omega}$ is the temperature and, similar to Haarnoja et al. (2018), can be updated by minimizing the following loss function:
+$L(\omega)=\mathbb{E}_{\tau_{t}}\left[-\alpha_{\omega}\log\pi_{jt}(\mathbf{u}_{t}|s_{t})-\alpha_{\omega}\bar{\mathcal{H}}\right],$ (11)
+where $\bar{\mathcal{H}}$ is a predefined target entropy.
+Note that the local policies (i.e., $\pi_{i}$) are only used to collect training roll-outs $\tau_{t}$.
+During execution, we would select the argmax of $f_{\theta}$, which does not require any centralized information (e.g., $s$), since OPTs preserve the order of the input vectors.
+The pseudo code of ME-IGM algorithm is presented in Algorithm 1.
+6 Experiments
+In this section, we first evaluate our method on a classic matrix game, demonstrating its capability to learn a globally optimal policy despite the non-monotonic nature of the overall payoff matrix.
+Next, we compare our algorithm against a set of baselines on the SMAC-v2 and Overcooked benchmark, highlighting its state-of-the-art performance in cooperative MARL.
+Moreover, we conduct ablation studies to evaluate the benefits of entropy-driven exploration, assess the necessity of incorporating the order-preserving transformation, and analyze the impact of key hyperparameters.
+Figure 4: The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked.
+It shows that ME-QMIX and ME-QPLEX achieve higher returns and exhibit faster convergence, compared to QMIX and QPLEX which do not adopt maximum entropy.
+MARL ME-IGM (Ours) IL+MARL
+Task Scenario MAPPO IPPO QMIX QPLEX HASAC FOP ME-QMIX ME-QPLEX IMAX-PPO InQ
+Protoss 5_vs_5 58.0 54.6 70.2 53.3 22.9 37.8 75.9$\pm$3.0 69.8$\pm$1.8 68.1 78.7
+10_vs_10 58.3 58.0 69.0 53.7 12.1 8.6 78.5$\pm$2.7 78.7$\pm$3.4 59.6 79.8
+10_vs_11 18.2 20.3 42.5 22.8 5.7 0.4 50.4$\pm$2.9 44.9$\pm$3.2 21.3 48.7
+20_vs_20 38.1 44.5 69.7 27.2 - 2.0 73.9$\pm$3.0 38.3$\pm$4.0 76.3 80.6
+20_vs_23 5.1 4.1 16.5 4.8 - 0.3 23.4$\pm$1.0 18.7$\pm$1.3 11.8 24.2
+Terran 5_vs_5 52.0 56.2 58.4 70.0 37.9 57.1 70.2$\pm$2.6 70.0$\pm$4.8 53.3 69.9
+10_vs_10 58.1 57.3 65.8 66.1 20.4 15.8 72.6$\pm$1.8 69.4$\pm$0.2 58.4 72.2
+10_vs_11 28.6 31.0 39.4 41.4 9.1 7.5 49.6$\pm$1.6 51.0$\pm$0.4 28.4 53.9
+20_vs_20 52.8 49.6 57.6 23.9 - 0.2 59.3$\pm$1.3 45.1$\pm$7.2 35.9 65.4
+20_vs_23 11.2 10.0 10.0 7.0 - 0.0 18.7$\pm$0.7 17.4$\pm$2.9 4.7 17.7
+Zerg 5_vs_5 41.0 37.2 37.2 47.8 29.1 35.2 50.8$\pm$4.6 52$\pm$3.9 48.6 55.0
+10_vs_10 39.1 49.4 40.8 41.6 17.9 8.0 49.1$\pm$0.6 42.9$\pm$1.8 50.6 57.6
+10_vs_11 31.2 26.0 28.0 31.1 14.0 1.5 32.6$\pm$1.7 45.3$\pm$0.6 34.8 41.5
+20_vs_20 31.9 31.2 30.4 15.8 - 0.2 34.3$\pm$5.2 42.1$\pm$5.1 26.7 43.3
+20_vs_23 15.8 8.3 10.1 6.7 - 0.2 14.8$\pm$6.4 17.5$\pm$3.2 8.2 21.3
+Table 2: We record the average win rates of different algorithms on SMAC-v2.
+The results demonstrate that our method, ME-IGM, outperforms all previous baseline algorithms and achieves performance comparable to IL+MARL methods, despite not having access to expert opponent demonstrations or opponent modeling.
+6.1 Evaluation on Matrix Games
+We first demonstrate the effectiveness of our algorithm in a classic one-step matrix game, shown as Table 1.
+It is a non-monotonic payoff matrix commonly used in previous studies (Son et al.
+2019; Wang et al.
+2020).
+We evaluate ME-QMIX, a variant of ME-IGM built upon QMIX, by comparing its performance against two state-of-the-art MARL methods: QMIX (Rashid et al.
+2020), a value decomposition MARL method, and FOP (Zhang et al.
+2021), a Maximum Entropy MARL method.
+We are concerned with whether the joint action with the highest selection probability corresponds to the globally maximal Q value, ensuring optimal returns when executing a deterministic policy.
+As shown in Table 1, QMIX suffers from high estimation error in payoff values of this non-monotonic game, due to its monotonicity constraints (as introduced in Section 2.3).
+Its deterministic policy converges to a suboptimal joint action (B, B) with a reward of 0.
+FOP, which employs a network architecture similar to QPLEX, is capable of accurately estimating value functions (i.e., the payoff).
+However, its policy tends to select suboptimal joint actions more frequently due to the misalignment between local policies and the maximum global Q-values, as discussed in Section 3.
+ME-QMIX exhibits smaller estimation errors for the payoff of optimal joint actions (which is our primary concern) while having larger errors for suboptimal ones.
+Crucially, the joint action with the maximum Q-value in ME-QMIX aligns with the true optimal joint action, and the local policy assigns the highest probability to selecting this optimal joint action, avoiding the misalignment issue through the use of OPTs.
+6.2 Evaluation on Overcooked
+To demonstrate the advantage of employing the principle of maximum entropy, we compare ME-IGM with two leading value decomposition MARL methods (which do not involve maximum entropy): QMIX and QPLEX, on the Overcooked benchmark (Carroll et al.
+2019) – a widely used coordination-focused multi-agent environment.
+In particular, we evaluate them in two representative layouts: cramped room and asymmetric advantages.
+These scenarios pose distinct challenges, including close collaboration under spatial constraints and asymmetric advantages in task division, and require coordinated exploration.
+For each scenario, we conduct repeated experiments with different random seeds to get the training curves of 10 million environment steps.
+As shown in Figure 4, ME-QMIX and ME-QPLEX exhibit significantly faster convergence and achieve higher performance levels in the early stages of training compared to their counterparts, QMIX and QPLEX.
+6.3 Evaluation on SMAC-v2
+We further conduct a comprehensive evaluation of ME-IGM on the SMACv2 (Ellis et al.
+2022) benchmark across 15 scenarios, comparing its performance against 8 baselines.
+The results, as shown in Table 2, are obtained by training each algorithm on each scenario for 101010 millions of environment steps.
+Each value in the table represents the average performance over three runs with different random seeds.
+ME-IGM outperforms the baseline algorithms across all scenarios by a significant margin.
+MAPPO and IPPO, as on-policy algorithms, have lower sampling efficiency, resulting in lower win rates within  [^10] M environment steps.
+QMIX and QPLEX, lacking exploration, tend to converge to local optimum.
+HASAC (Liu et al.
+2023) and FOP perform poorly across all scenarios.
+Surprisingly, ME-IGM achieves performance on par with (or even surpassing) IL+MARL methods, despite not relying on expert opponent demonstrations or opponent modeling.
+The results of MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, InQ are taken from (Bui, Mai, and Nguyen 2024).
+A detailed comparison with actor-critic baselines is presented in the Appendix.
 6.4 Ablation Study
-We evaluate the effectiveness of different components in our algorithm through an ablation study.
-We begin with QMIX, which utilizes epsilon-greedy exploration.
-We then introduce +entropy, which incorporates softmax sampling (employing $\text{softmax}(Q_{i})$ as the policy) along with a maximum entropy objective function to the QMIX framework.
-Note that this version of the algorithm does not ensure convergence, leading to a performance decline in two out of three scenarios.
-The subsequent modification, +function f, integrates the function $f_{i}$ (using $\text{softmax}(f_{i}(Q_{i})$) as the policy) building upon the +entropy setup.
-Empirical results show that function f yields the most substantial performance improvement across all components.
-Our study culminates with Soft-QMIX, our finalized algorithm.
-For each modification, we conduct four experiments across three seeds, reporting the average winning rate and its standard deviation in our results table.
-Table 4: The ablation studies reveal that each component enhances performance, with function $f$ contributing most significantly to the performance improvement.
-eval win rate QMIX +entropy +function f Soft-QMIX(ours)
-protoss 5vs5 0.68 $\pm$ 0.04 0.73 $\pm$ 0.05 0.74 $\pm$ 0.04 0.76 $\pm$ 0.04
-terran 5vs5 0.68 $\pm$ 0.05 0.67 $\pm$ 0.04 0.70 $\pm$ 0.05 0.72 $\pm$ 0.04
-zerg 5vs5 0.41 $\pm$ 0.05 0.40 $\pm$ 0.04 0.48 $\pm$ 0.05 0.51 $\pm$ 0.04
+Figure 5: Comparison of exploration strategies in CTDE MARL: ME-QMIX with maximum entropy vs. QMIX with epsilon-greedy and extended epsilon annealing period.
+Benefits of entropy-driven exploration: To further demonstrate the benefits of using the maximum entropy framework for exploration, we compare ME-QMIX with a modified version of QMIX, which incorporates significantly higher exploration by applying an extended epsilon annealing period in the epsilon-greedy method, on Protoss 5v5.
+Figure 5 illustrates that simply increasing the epsilon annealing time for exploration does not improve QMIX’s performance.
+ME-QMIX achieves superior performance, with the systematic exploration enabled by the maximum entropy framework.
+QMIX +Entropy +OPT
+Protoss 5vs5 0.68 $\pm$ 0.04 0.73 $\pm$ 0.05 0.74 $\pm$ 0.04
+Terran 5vs5 0.68 $\pm$ 0.05 0.67 $\pm$ 0.04 0.70 $\pm$ 0.05
+Zerg 5vs5 0.41 $\pm$ 0.05 0.40 $\pm$ 0.04 0.48 $\pm$ 0.05
+Table 3: Our algorithm (+OPT) vs. its two ablated versions: +Entropy, obtained by removing OPTs, and QMIX, obtained by further removing softmax local policies.
+The necessity of using OPTs: In Table 3, we compare QMIX with its two variants: +Entropy and +OPT.
+Both variants learn local stochastic policies based on Q-functions trained with the QMIX objective.
+Specifically, +OPT corresponds to our proposed ME-IGM, while +Entropy removes the OPTs from ME-IGM (i.e., no OPTs in Equations (4), (9), and (10)).
+We test each algorithm on three SMAC-v2 scenarios across four random seeds, reporting the average win rate along with its standard deviation in the table.
+Without OPTs, +Entropy suffers from the misalignment problem discussed in Section 3 and fails to ensure monotonic policy improvement, as analyzed in Section 4.2, leading to a performance decline in two out of three scenarios.
+On the other hand, +OPT achieves a substantial performance improvement across all evaluated scenarios.
+(a) Impact of the learning rate for $\alpha_{\omega}$
+(b) Impact of the target entropy $\bar{\mathcal{H}}$
+Figure 6: The impact of different hyperparameters for training $\alpha_{\omega}$ on the performance of ME-QMIX.
+(a) A lower learning rate for $\alpha_{\omega}$ corresponds to increased exploration.
+(b) ME-QMIX’s performance is not sensitive to the choice of the target entropy $\bar{\mathcal{H}}$ in Equation (10).
+ME-QMIX (MLP) ME-QMIX (OPT)
+Protoss 5v5 0.58±0.030.58$\pm$ 0.030.58 ± 0.03 0.74±0.040.74$\pm$ 0.040.74 ± 0.04
+Terran 5v5 0.61±0.040.61$\pm$ 0.040.61 ± 0.04 0.70±0.050.70$\pm$ 0.050.70 ± 0.05
+Zerg 5v5 0.41±0.040.41$\pm$ 0.040.41 ± 0.04 0.48±0.050.48$\pm$ 0.050.48 ± 0.05
+Table 4: ME-QMIX (OPT) vs. its ablated version: ME-QMIX (MLP), where OPT is replaced with a 3-layer MLP.
+To further validate the design of the order-preserving transformation, we replace the OPT layers in ME-QMIX with a 3-layer MLP, which does not enforce non-negative weights, referring to this variant as ME-QMIX (MLP).
+As shown in Table 4, our method, ME-QMIX (OPT), consistently outperforms ME-QMIX (MLP) across three different SMAC-v2 scenarios.
+This underscores the importance of the order-preserving transformation (i.e., Eq.
+(2)) in resolving the misalignment between local policies and the maximum global Q-value, thereby enhancing overall performance.
+The impact of hyperparameters: Compared to QMIX, ME-QMIX introduces two additional hyperparameters: the learning rate and the target entropy $\bar{\mathcal{H}}$ for updating $\alpha_{\omega}$ in Eq.
+(10).
+We set the other hyperparameters of ME-QMIX following the design choices of QMIX.
+To analyze the impact of these additional parameters, we conduct a parameter scan on the learning rate of $\alpha_{\omega}$ and the target entropy $\bar{\mathcal{H}}$.
+Figure 6(a) illustrates that the learning rate of $\alpha_{\omega}$ influences the exploration level.
+A lower learning rate results in higher policy entropy at the beginning of training, promoting exploration.
+While this initially leads to weaker performance (e.g., the blue curve in Figure 6(a)), it eventually achieves higher returns.
+Furthermore, the results in Figure 6(b) suggest that the algorithm’s performance is not sensitive to the choice of the target entropy.
+Notably, in this study, we did not fine-tune configurations for each individual testing scenario.
+Please check the Appendix for the tables of hyperparameters.
 7 Conclusion
-In this paper, we introduced Soft-QMIX, an innovative approach within the CTDE framework, integrating the maximum entropy RL framework with the QMIX-based credit assignment mechanism.
-Theoretically, we established the monotonic improvement and convergence of Soft-QMIX to a globally optimal solution.
-Experimentally, our algorithm can solve matrix games and demonstrate state-of-the-art performance on the SMAC-v2 benchmark.
-Currently, Soft-QMIX is a value-based method, thus one of its limitations is that it can currently only be deployed in tasks with discrete action spaces.
-In the future, we will try to extend it to tasks with continuous action spaces.
+In this paper, we introduce ME-IGM, a novel maximum entropy MARL approach that applies the CTDE framework and satisfies the IGM condition.
+We identify and address the misalignment problem between local policies and the maximum global Q-value (in maximum entropy MARL) by introducing an order-preserving transformation (OPT).
+We further propose a theoretically sound and straightforward objective for updating these transformation operators.
+Empirically, we demonstrate the effectiveness of ME-IGM on matrix games and its state-of-the-art performance on the challenging Overcooked and SMAC-v2 benchmarks.
+However, as a value-based method, ME-IGM is currently limited to tasks with discrete action spaces.
+In future work, we plan to extend its use to tasks involving continuous action spaces.
 Acknowledgment
 This work was supported in part by the U.S. Army Futures Command under Contract No. W519TC-23-C-0030.
-Appendix A Soft-QMIX
-Algorithm 1 Soft-QMIX
-1:  Input: $\theta,\phi$
-2:  Initiate: $\bar{\theta}\leftarrow\theta,D\leftarrow\emptyset$
-3:  for each iteration do
-4:     for each environment step do
-5:        get $\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t})$ via Eq. 15
-6:        $u^{i}_{t}\sim\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t}),\forall i\in\{1,..,n\}$
-7:        $s_{t+1}\sim p(s_{t+1}|s_{t},\mathbf{u}_{t})$
-8:        $D\leftarrow D\cup\{(s_{t},\mathbf{u}_{t},r(s_{t},\mathbf{u}_{t}),s_{t+1})\}$
-9:     end for
-10:     for each gradient step do
-11:        calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq. 24
-12:        $\theta\leftarrow\theta-\lambda_{Q}\nabla_{\theta}L_{Q}(\theta)$
-13:        $\phi\leftarrow\phi-\lambda_{\pi}\nabla_{\phi}L_{\pi}(\phi)$
-14:        $\bar{\theta}\leftarrow\tau\theta+(1-\tau)\bar{\theta}$
-15:     end for
-16:  end for
-Appendix B Individual-Global-Independence (IGI)
-In this section, we first introduce two concepts, IGI and IGO, where IGI is proposed in our work, and IGO was proposed by A. We then explain that IGI is a sufficient condition for IGO.
-Definition B.1.
-Individual-Global-Independence (IGI) For joint policy $\pi_{jt}$, if their exist individual policies $\left[\pi_{i}\right]_{i=1}^{n}$ such that the following holds:
-$\pi_{jt}(\mathbf{u}|s)=\prod_{i=1}^{n}\pi_{i}(u^{i}|o^{i}).$ (16)
-Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGI for $\pi_{jt}$ under $s_{t}$.
-Definition B.2.
-Individual-Global-Optimal (IGO) For optimal joint policy $\pi^{*}_{jt}$, if their exist individual optimal policies $\left[\pi^{*}_{i}\right]_{i=1}^{n}$ such that the following holds:
-$\pi^{*}_{jt}(\mathbf{u}|s)=\prod_{i=1}^{n}\pi^{*}_{i}(u^{i}|o^{i}).$ (17)
-Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGO for $\pi_{jt}$ under $s_{t}$.
-IGO defines that the optimal joint policy equals the product of optimal local policies.
-In IGI, it is defined that any joint policy should equal the product of local policies.
-If finding the optimal policy is viewed as an optimization problem, IGI defines the feasible set of this problem.
-Hence, any policy not equaling the product of local policies is not a feasible solution and thus cannot be optimal.
-Therefore, all optimal policies will satisfy the joint policy equals the product of local optimal policies, thereby meeting the IGO condition.
-Note that under the assumption of a decentralized execution setting, IGI and IGO are equivalent.
-Appendix C Joint Policy Iteration
-Lemma C.1 (Joint Soft Policy Evaluation).
-Let $Q^{0}_{tot}:S\times\mathbf{U}\to\mathbb{R}$ be a bounded mapping where $|\mathbf{U}|<\infty$.
-Define the iterative process $Q^{k+1}_{tot}=\mathcal{T}^{\pi_{jt}}_{\lambda}(Q^{k}_{tot})$.
-It follows that the sequence $\{Q^{k}_{tot}\}_{k=0}^{\infty}$ converges to the joint soft-Q-function for policy $\pi_{jt}$ as $k\rightarrow\infty$.
+[…]
+[…]
+$\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u_{t}})-Q_{tot}(s_{t},\mathbf{u_{t}})=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{V}_{t+l}$ (12)
+where $\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$.
 Proof.
-Initiate by defining the entropy-augmented reward function:
-$r_{\pi_{jt}}(s_{t},\mathbf{u}_{t})\coloneqq r(s_{t},\mathbf{u}_{t})+\mathbb{E}_{s_{t+1}}\left[H[\pi_{jt}(\cdot|s_{t+1})]\right].$ (18)
-Follow previous work [9], we choose a lower $\lambda$ and can assume that $\mathcal{T}^{\pi_{jt}}_{\lambda}$ is an unbiased estimator of $\mathcal{T}^{\pi_{jt}}$.
-Subsequently, the update mechanism is articulated as:
-$Q_{tot}(s_{t},\mathbf{u}_{t})$ $\leftarrow r_{\pi_{jt}}(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1},\mathbf{u}_{t+1}\sim\pi_{jt}}[Q_{tot}(s_{t+1},\mathbf{u}_{t+1})].$ (19)
-Conclude by invoking the established convergence theorems pertinent to policy evaluation as explicated by Sutton and Barto. ∎
-Lemma C.2 (Joint Soft Policy Improvement).
+The one-step TD error can be expressed in the following form:
+$\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$ (13)
+$\delta_{t+1}^{V}=-Q_{tot}(s_{t+1},\mathbf{u}_{t+1})+r_{t+1}+\gamma V_{tot}(s_{t+2})$
+Next, the k-step TD error can be represented in the following form:
+$\epsilon_{t}^{(1)}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})=\delta^{V}_{t}$ (14)
+$\epsilon_{t}^{(2)}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma r_{t+1}{-\log\pi_{jt}(\mathbf{u_{t+1}}|s_{t+1})}+\gamma^{2}V_{tot}(s_{t+1})=\delta^{V}_{t}+\gamma\delta^{V}_{t+1}$
+$\epsilon_{t}^{(k)}=\sum_{l=0}^{k-1}\gamma^{l}\delta^{V}_{t+l}$
+$\epsilon_{t}^{td(\lambda)}=(1-\lambda)\left(\epsilon_{t}^{(1)}+\lambda\epsilon_{t}^{(2)}+\lambda^{2}\epsilon_{t}^{(3)}+\cdots\right)$ (15)
+$=(1-\lambda)\big{(}\delta_{t}^{V}+\lambda(\delta_{t}^{V}+\gamma\delta_{t+1}^{V})+\lambda^{2}(\delta_{t}^{V}+\gamma\delta_{t+1}^{V}+\gamma^{2}\delta_{t+2}^{V})+\cdots\big{)}$
+$=(1-\lambda)\left(\delta_{t}^{V}(1+\lambda+\lambda^{2}+\cdots)+\gamma\delta_{t+1}^{V}(\lambda+\lambda^{2}+\cdots)\right.\left.+\gamma^{2}\delta_{t+2}^{V}(\lambda^{2}+\lambda^{3}+\lambda^{4}+\cdots)+\cdots\right)$
+$=(1-\lambda)\left(\delta_{t}^{V}\left(\frac{1}{1-\lambda}\right)+\gamma\delta_{t+1}^{V}\left(\frac{\lambda}{1-\lambda}\right)+\cdots\right)$
+$=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{V}$
+It is important to note that when computing the TD lambda error in the context of a reverse view and maximum entropy reinforcement learning, it is necessary to subtract the $\log(\pi)$ term from the $t+1$ step’s return as well as subtract $\log(\pi)$ from $Q_{t+1}$ in order to obtain results consistent with Equation 15.
+Appendix B B. Proof of Theorem 4.1
+Lemma B.1 (Joint Soft Policy Improvement).
 Suppose $\pi_{jt}^{\text{old}}\in\Pi$ and consider $\pi_{jt}^{\text{new}}$ to be the solution to Equation 3.
@@
 Let $\pi_{jt}^{\text{old}}\in\Pi$ and let $Q^{\pi_{jt}^{\text{old}}}$ and $V^{\pi_{jt}^{\text{old}}}$ be the corresponding soft state-action value and soft state value.
-The update rule of $\pi_{\text{new}}$ can be defined as
+The update rule of $\pi_{\text{new}}$ can be defined as:
 $\pi_{\text{new}}(\cdot\mid s_{t})$ $=\arg\min_{\pi^{\prime}\in\Pi}\text{D}_{KL}\left(\pi^{\prime}(\cdot)\parallel\exp(Q^{\pi_{jt}^{\text{old}}}(s_{t},\cdot))-\log Z^{\pi_{jt}^{\text{old}}}(s_{t})\right)$ (16)
-$=\arg\min_{\pi^{\prime}\in\Pi}J_{\pi_{jt}^{\text{old}}}\left(\pi^{\prime}(\cdot\mid s_{t})\right).$
+$=\arg\min_{\pi^{\prime}\in\Pi}J_{\pi_{jt}^{\text{old}}}\left(\pi^{\prime}(\cdot\mid s_{t})\right),$
+where $Z^{\pi_{jt}^{\text{old}}}(s_{t})$ is the normalization term.
 Since we can always choose $\pi_{\text{new}}=\pi_{\text{old}}\in\Pi$, the following inequality always hold true: $J_{\pi_{jt}^{\text{old}}}(\pi_{\text{new}}(\cdot\mid s_{t}))\leq J_{\pi_{jt}^{\text{old}}}(\pi_{jt}^{\text{old}}(\cdot\mid s_{t}))$.
@@
 $\leq Q^{\pi_{jt}^{\text{new}}}(s_{t},\mathbf{u}_{t}),$
-where we can repeatedly expand $Q^{\pi_{jt}^{\text{old}}}$ on the RHS by applying the soft Bellman equation and the bound in Equation 23.
-Convergence to $Q^{\pi_{jt}^{\text{new}}}$ follows from Lemma B.1.
-Theorem C.3 (Joint Soft Policy Iteration).
-By continually applying joint soft policy evaluation and joint soft policy improvement to any $\pi_{jt}\in\Pi$ converges to a policy $\pi^{*}_{jt}$ such that $Q^{\pi^{*}_{jt}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi_{jt}}(s_{t},\mathbf{u}_{t})$ for all $\pi_{jt}\in\Pi$ and $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$, assuming $|\mathbf{U}|<\infty$.
+where we can repeatedly expand $Q^{\pi_{jt}^{\text{old}}}$ on the RHS by applying the soft Bellman equation, progressing from the second line to the final one.
+Theorem B.2.
+Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$.
+Updating the policy with Equation (4) ensures that $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})-\epsilon\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}$.
+The gap $\epsilon$ satisfies:
+$\epsilon\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim\mathcal{P}}\big{[}C\sqrt{2D_{KL}\big{(}\pi_{\text{old}}(\cdot|s_{t+1})|\pi_{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$ (20)
+where $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function and $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ |Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|$.
 Proof.
-Let $\pi_{jt}^{i}$ represent the policy after the $i$-th update.
-By Lemma B.2, the sequence $Q^{\pi_{jt}^{i}}$ is monotonically increasing.
-Given that $Q^{\pi}$ is bounded above for $\pi_{jt}\in\Pi$ due to bounded rewards and entropy terms, the sequence converges to some $\pi^{*}_{jt}$.
-It remains to demonstrate that the policy $\pi^{*}_{jt}$ is indeed the optimal policy.
-At convergence, we infer that $J_{\pi^{*}_{jt}}(s)\geq J_{\pi_{jt}}(s)$ for all $\pi_{jt}\in\Pi$, if $\pi_{jt}\neq\pi^{*}_{jt}$.
-Using the same iterative argument as in the proof of Lemma B.2, we get $Q^{\pi^{*}_{jt}}(s,\mathbf{u})\geq Q^{\pi}(s,a)$ for all $(s,\mathbf{u})\in S\times\mathbf{U}$, that is, the soft value of any other policy in $\Pi$ is lower than that of the converged policy.
-Hence $\pi^{*}_{jt}$ is optimal in $\Pi$.
-∎
-Appendix D TD($\lambda$)
-In this paragraph, we will derive the expression of TD($\lambda$) in the context of Maximum Entropy Reinforcement Learning.
-Specifically, we aim to demonstrate the derivation of the following equation:
-$\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u_{t}})-Q_{tot}(s_{t},\mathbf{u_{t}})=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{V}_{t+l}$ (24)
-where $\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$.
-Proof.
-The one-step TD error can be expressed in the following form:
-$\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$ (25)
-$\delta_{t+1}^{V}=-Q_{tot}(s_{t+1},\mathbf{u}_{t+1})+r_{t+1}+\gamma V_{tot}(s_{t+2})$
-Next, the k-step TD error can be represented in the following form:
-$\epsilon_{t}^{(1)}:=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})=\delta^{V}_{t}$ (26)
-$\epsilon_{t}^{(2)}:=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma r_{t+1}{\color[rgb]{1,0,0}-\log\pi_{jt}(\mathbf{u_{t+1}}|s_{t+1})}+\gamma^{2}V_{tot}(s_{t+1})=\delta^{V}_{t}+\gamma\delta^{V}_{t+1}$
-$\epsilon_{t}^{(k)}=\sum_{l=0}^{k-1}\gamma^{l}\delta^{V}_{t+l}$
-$\epsilon_{t}^{td(\lambda)}:=(1-\lambda)\left(\epsilon_{t}^{(1)}+\lambda\epsilon_{t}^{(2)}+\lambda^{2}\epsilon_{t}^{(3)}+\cdots\right)$ (27)
-$=(1-\lambda)\big{(}\delta_{t}^{V}+\lambda(\delta_{t}^{V}+\gamma\delta_{t+1}^{V})+\lambda^{2}(\delta_{t}^{V}+\gamma\delta_{t+1}^{V}+\gamma^{2}\delta_{t+2}^{V})+\cdots\big{)}$
-$=(1-\lambda)\left(\delta_{t}^{V}(1+\lambda+\lambda^{2}+\cdots)+\gamma\delta_{t+1}^{V}(\lambda+\lambda^{2}+\cdots)\right.\left.+\gamma^{2}\delta_{t+2}^{V}(\lambda^{2}+\lambda^{3}+\lambda^{4}+\cdots)+\cdots\right)$
-$=(1-\lambda)\left(\delta_{t}^{V}\left(\frac{1}{1-\lambda}\right)+\gamma\delta_{t+1}^{V}\left(\frac{\lambda}{1-\lambda}\right)+\cdots\right)$
-$=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{V}$
-It is important to note that when computing the TD lambda error in the context of a reverse view and maximum entropy reinforcement learning, it is necessary to subtract the $\log(\pi)$ term from the $t+1$ step’s return as well as subtract $\log(\pi)$ from $Q_{t+1}$ in order to obtain results consistent with Equation 27.
-Appendix E Soft-Q Learning and Policy Gradient
-Theorem E.1.
-If $Q$-function is parameterized as $Q_{\theta}$, and $V_{\theta},\pi_{\theta}$ are defined as below:
-$V_{\theta}(s)$ :-$\coloneq\alpha\log\mathbb{E}_{\mathbf{u}\sim\pi_{jt}}\left[\exp\left(Q_{\theta}(s,\mathbf{u})/\alpha\right)\right],:$ (28)
-$\pi_{jt_{\theta}}(\mathbf{u}|s)$ :-$\coloneq\exp\left(Q_{\theta}(s,\mathbf{u})-V_{\theta}(s)\right)/\alpha.:$
-and $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right].$ Then, we have
-$\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}$ $\left[\|Q_{\theta}(s,\mathbf{u})-y\|^{2}\right]$ (29)
-$=\alpha\nabla_{\theta}D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$
-$+\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\|V_{\theta}(s)-y\|^{2}\right]$
-Proof.
-The proof is heavily based on [25]. From assumptions described in Eq. 11 we have:
-$Q_{\theta}(s,\mathbf{u})=V_{\theta}(s)+\alpha\log\pi_{jt_{\theta}}(\mathbf{u}|s)$ (30)
-For simplicity, we further define notation $\Delta_{t}$ as below.
-$\Delta_{t}$ $=\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})-Q_{\theta}(s_{t},\mathbf{u_{t}}).$
-From assumption, we have $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right]$.
-Thus, the relationship between $\Delta_{t}$ and $y_{t}$ can be writen as:
-$\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[y_{t}]$ $=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right]$ (31)
-$=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[Q_{\theta}(s_{t},\mathbf{u_{t}})+\Delta_{t}]$
-$=\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}[V_{\theta}(s_{t})+\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]+\Delta_{t}].$
-Here are two equations that will be used in the following proof:
-$\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s)\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u}|s)\big{)}\big{]}$ (32)
-$=$ $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s)\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u}|s)]\big{)}\big{]}.$
-$\mathbb{E}_{a\sim\pi}\big{[}\nabla\log\pi(a)f(s)\big{]}=f(s)\sum_{a}\pi(a)\frac{\nabla\pi(a)}{\pi(a)}$ (33)
-$=f(s)\sum_{a}\nabla\pi(a)=f(s)\nabla\sum_{a}\pi(a)=0,$
-where $f$ can be any function that does not conditioned on $a$, and $\sum_{a}\pi(a)=1$. Let’s consider the gradient of the $Q$-function’s MSE loss:
-$\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\frac{1}{2}\left(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t}\right)^{2}\right]\bigg{|}_{\pi=\pi_{\theta}}$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t})\right]\bigg{|}_{\pi=\pi_{\theta}}$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\left(V_{\theta}(s_{t})+\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-y_{t}\right)\right]$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\big{(}V_{\theta}(s_{t})+\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-V_{\theta}(s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\nabla_{\theta}Q_{\theta}(s_{t},\mathbf{u_{t}})\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s_{t})+\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u}|s)\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})-\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]-\Delta_{t}\big{)}\big{]}$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}$
-$\qquad-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}$
-$\qquad-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-$
-$\qquad\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})]\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}$
-$=\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}-\big{(}\nabla_{\theta}V_{\theta}(s_{t})\big{)}\Delta_{t}+\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\big{(}\alpha\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}-\big{(}\alpha\nabla_{\theta}\log\pi_{jt_{\theta}}(\mathbf{u_{t}}|s_{t})\big{)}\Delta_{t}\big{]}.$
-for simplicity, we ignore the $|_{\pi=\pi_{\theta}}$ after the first two lines of equations.
-The second equation use Eq.
-30.
-The third equation use Eq.
-31.
-The fourth equation cancled out $V_{\theta}(s_{t})$.
-The fifth equation use Eq.
-30 again.
-The sixth equation expand out terms.
-The seventh equation use Eq.
-32.
-The last equation use Eq.
-33.
-Let’s consider the gradient of the KL divergence.
-$\nabla_{\theta}\alpha D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$
-$=\nabla_{\theta}\mathbb{E}_{\pi_{jt}}\big{[}\alpha\log\pi_{jt}(\mathbf{u}|s)-y(\mathbf{u},s)+\alpha\log Z(s)\big{]}$
-$=\nabla_{\theta}\mathbb{E}_{\pi_{jt}}\big{[}\alpha\log\pi_{jt}(\mathbf{u}|s)-y(\mathbf{u},s)\big{]}$
-$=\nabla_{\theta}[\alpha\sum_{\mathbf{u}}\pi_{jt}(\mathbf{u}|s)\log\pi_{jt}(\mathbf{u}|s)]-\nabla_{\theta}[\sum_{\mathbf{u}}\pi_{jt}(\mathbf{u}|s)y(\mathbf{u},s)]$
-$=\alpha\sum_{\mathbf{u}}\nabla_{\theta}\pi_{jt}(\mathbf{u}|s)\log\pi_{jt}(\mathbf{u}|s)-\sum_{\mathbf{u}}\nabla_{\theta}\pi_{jt}(\mathbf{u}|s)y(\mathbf{u},s)$
-$=\mathbb{E}_{\pi_{jt}}\big{[}\alpha\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\log\pi_{jt}(\mathbf{u}|s)-\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}y(\mathbf{u},s)\big{]}.$
-Note that $V_{\theta}(s)+\alpha\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[\log\pi_{jt_{\theta}}(\mathbf{u}|s)]$ is not a function of $\mathbf{u}$. With Eq. 33, we have:
-$\nabla_{\theta}\alpha D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}$ (34)
-$=\mathbb{E}_{\pi_{jt}}\big{[}\alpha\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\log\pi_{jt}(\mathbf{u}|s)-\big{(}\nabla_{\theta}\log\pi_{jt}(\mathbf{u}|s)\big{)}\Delta_{t}\big{]}$
-Finally, we got:
-$\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\left[\frac{1}{2}\left(Q_{\theta}(s_{t},\mathbf{u_{t}})-y_{t}\right)^{2}\right]\bigg{|}_{\pi=\pi_{\theta}}$
-$=\nabla_{\theta}\alpha^{2}D_{KL}\big{(}\pi_{jt}(\cdot|s_{t})\|\frac{exp(y(s)/\alpha)}{Z(s)}\big{)}+\frac{1}{2}\nabla_{\theta}\mathbb{E}_{s_{t},\mathbf{u_{t}}\sim\pi_{jt_{\theta}}}\big{[}(V_{\theta}(s_{t})-y_{t})^{2}\big{]},$
-where $y(s)$ is the distribution of $y_{t}$ at state $s$.
-Appendix F Unbiased Assumption
-We assume that $\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u_{t}})$ is an unbiased estimate of $E_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}\sum_{i}f_{i}(Q_{i}(s_{t},\mathbf{u_{t}}))\right]$ in the main paper.
-Practically, this assumption might fail, leading to non-monotonic increases in the policy improvement step.
-Here, we initially explain the impact of this error and derive an upper bound for it.
-Conclusion: the error’s upper bound is proportional to $|Q_{tot}-\sum_{i}f_{i}(Q_{i})|$ and $\sqrt{D_{KL}\big{(}\pi_{jt}^{\text{old}}|\pi_{jt}^{\text{new}}\big{)}}.$ We use Equation 14 as the loss function, ensuring that the first term decreases throughout training.
-Additionally, by selecting a sufficiently small learning rate, the second term will also remain minimal.
-Derivative: We can replace both $Q_{\theta}$ and $y_{t}$ in Theorem 4.2 with $\sum_{i}f_{i}(Q_{i})$, and get that minimizing the L2 norm described in Equation 10 is equivalent to minimizing the td loss of value function and minimizing the KL divergence described as follow,
-$\pi_{jt}^{\text{new}}=\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{arg\,min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(\sum_{i}f_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (35)
-[…]
-Similar to Equation 22, during policy improvement step, we have:
-$\mathbb{E}_{\mathbf{u}_{t}\sim\pi_{jt}^{\text{new}}}\left[\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t},u^{i}_{t}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t}\mid s_{t})\right]\leq V^{\prime\pi_{jt}^{\text{old}}}(s_{t}).$ (36)
-During policy evaluation step, we have:
+We define the value function as $V^{\prime\pi_{jt}^{\text{new}}}(s)=\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{new}}}(o^{i}_{t+1},u^{i}_{t+1}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}$.
+Then, similar to Equation 18, during policy improvement step, we have:
+$\mathbb{E}_{\mathbf{u}_{t}\sim\pi_{jt}^{\text{new}}}\left[\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t},u^{i}_{t}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t}\mid s_{t})\right]\geq V^{\prime\pi_{jt}^{\text{old}}}(s_{t}).$ (21)
+By expanding the soft Bellman equation, we have:
 $Q_{tot}^{\pi_{jt}^{\text{old}}}(s_{t},\mathbf{u}_{t})$ $=r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}\sim p}\left[V_{tot}^{\pi_{jt}^{\text{old}}}(s_{t+1})\right]$ (22)
 $=r(s_{t},\mathbf{u}_{t})+\gamma\big{(}\mathbb{E}_{s_{t+1}\sim p}\left[V^{\prime\pi_{jt}^{\text{old}}}(s_{t+1})\right]+\epsilon_{1}\big{)}$
-$\leq r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}+\gamma\epsilon_{1}$
+$\leq r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}+\gamma\epsilon_{1}$
 $=r(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1}}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\log\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}\mid s_{t+1})\big{]}\big{]}+\gamma(\epsilon_{1}+\epsilon_{2})$
@@
 $\leq Q_{tot}^{\pi_{jt}^{\text{new}}}(s_{t},\mathbf{u}_{t})+\epsilon,$
-where $\epsilon$ is s the cumulative sum of $\epsilon_{1}$ and $\epsilon_{2}$ over time.$\epsilon_{1}$ and $\epsilon_{2}$ are defined as follows:
+where $p$ is the transition function $\mathcal{P}(s_{t},a_{t},\cdot)$, and $\epsilon$ is the cumulative sum of $\epsilon_{1}$ and $\epsilon_{2}$ over time.
+$\epsilon_{1}$ and $\epsilon_{2}$ are defined as follows:
 $\epsilon_{1}$ $=\mathbb{E}_{s_{t+1}\sim p}\left[V_{tot}^{\pi_{jt}^{\text{old}}}(s_{t+1})-V^{\prime\pi_{jt}^{\text{old}}}(s_{t+1})\right]$ (23)
-$\epsilon_{2}$ $=\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
+$\epsilon_{2}$ $=\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
 Then, we have:
 $\epsilon_{1}+\epsilon_{2}$ $=\mathbb{E}_{s_{t+1}\sim p}\left[V_{tot}^{\pi_{jt}^{\text{old}}}(s_{t+1})-V^{\prime\pi_{jt}^{\text{old}}}(s_{t+1})\right]$ (24)
-$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
-$=\mathbb{E}_{s_{t+1}\sim p}\left[\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{old}}}\big{[}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{]}\right]$
-$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
-$=\mathbb{E}_{s_{t+1}\sim p}\big{[}\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{(}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{)}\big{]}$
-$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{]},$
-where $C=\underset{\mathbf{u}_{t+1}}{\mathrm{arg\,max}}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}f_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality [21], we have:
-$|\epsilon_{1}+\epsilon_{2}|$ $\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big{]}$ (40)
-$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sqrt{2D_{KL}\big{(}\pi_{jt}^{\text{old}}(\cdot|s_{t+1})|\pi_{jt}^{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$
-Appendix G Hyper-network Structure
+$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
+$=\mathbb{E}_{s_{t+1}\sim p}\left[\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{old}}}\big{[}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{]}\right]$
+$\quad+\mathbb{E}_{s_{t+1}\sim p}\big{[}\mathbb{E}_{\mathbf{u}_{t+1}\sim\pi_{jt}^{\text{new}}}\big{[}\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))-Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})\big{]}\big{]}$
+$=\mathbb{E}_{s_{t+1}\sim p}\big{[}\sum_{\mathbf{u}_{t+1}}\big{(}\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})\big{)}\big{(}Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))\big{)}\big{]}$
+$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})||Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|\big{]}$
+$\leq\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big{]},$
+where $C=\underset{\mathbf{u}_{t+1}}{\max}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality (Pinsker 1964), we have:
+$\epsilon\leq\frac{\gamma}{1-\gamma}|\epsilon_{1}+\epsilon_{2}|$ $\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big{]}$ (25)
+$\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim p}\big{[}C\cdot\sqrt{2D_{KL}\big{(}\pi_{jt}^{\text{old}}(\cdot|s_{t+1})|\pi_{jt}^{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$
+Appendix C C. Hyper-network Structure
 Figure 7: (a) The overall network architecture.
-Agent 1 to Agent n takes local observations as input and outputs a distribution over local actions $u^{i}$.
-Functions $g_{i}$ and $f_{i}$ are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values.
-Both the input and output of the $g_{i}$ and $f_{i}$ networks are k-dimensional, where $k$ is the action dimension.
-The output of the $f_{i}$ function, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled.
+Single Q-Net takes local observations as input and outputs a distribution over local actions $u^{i}$, also denoted as local $Q_{i}$.
+Functions OPTs are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values.
+Both the input and output of the OPTs are k-dimensional, where $k$ is the action dimension.
+The output of the OPT, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled.
 The Mixer network takes local $Q_{i}$ as input and outputs the global $Q_{tot}$.
@@
 Figure 7 shows the overall network structure.
-Specifically, function $f_{i}$ and function $g_{i}$ share the same network structure as the mixer network, but while the mixer network’s input dimension is (batch_size, agent_num) and its output dimension is (batch_size, 1), both the input and output dimensions of function $f_{i}$ and function $g_{i}$ are (batch_size, 1).
-Additionally, function $g_{i}$ introduces an additional residual from input to output.
+Specifically, OPTs share the same network structure as the mixer network, but while the mixer network’s input dimension is (batch_size, agent_num) and its output dimension is (batch_size, 1), both the input and output dimensions of OPTs are (batch_size, action_num).
 The network hyperparameters are as shown in the table below.
-layer name hypernet_embed embed_dim num_layer
+In the SMAC-v2 experiment, we select a single layer for the OPT model to ensure fairness in comparison.
+Using two layers of OPT would significantly increase the number of parameters, making it less comparable to the baseline algorithms.
+While this may slightly reduce our method’s performance, we adopt this choice for a fair comparison.
+layer name hyper-net_embed embed_dim num_layer
 mixer 64 32 2
-function $g_{i}$ 64 32 2
-function $f_{i}$ 64 - 1
-Table 5: Hyper-parameters used for hyper-networks.
-Appendix H Comparison with Baselines
-The following paragraphs present the differences between our method and those based on the AC:
+OPT 64 - 1
+Table 5: Hyperparameters used for hyper-networks.
+Appendix D D. Comparison with Baselines
+The following paragraphs present the differences between our method and those based on the Actor-Critic framework:
 1.
@@
 Actor-critic methods typically minimize the KL divergence between the actor and the softmax critic, whereas our method minimizes the MSE loss between their logits.
-[12] pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
+(Kim et al.
+2021) pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
 3.
@@
 In fact, it’s challenging to theoretically prove that value-based methods are inherently superior to AC methods.
-However, through experiments, we have demonstrated that our method surpasses both FOP [36] and HASAC [16].
+However, through experiments, we have demonstrated that our method surpasses both FOP (Zhang et al.
+2021) and HASAC (Liu et al.
+2023).
 Below are comparisons with individual papers.
-mSAC [22] introduces a method similar to MASAC.
+mSAC (Pu et al.
+2021) introduces a method similar to MASAC.
 It employs the AC architecture, thus encountering issues 1 and 2 mentioned above, but does not fully utilize global information to train the critic, hence lacking the advantage noted in point 3.
-FOP [36] shares a similar issue with [22], in that the AC framework it utilizes can actually train the critic with global state information.
-Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX [31].
-[9] experimentally shown that QMIX outperforms QPLEX.
+FOP (Zhang et al.
+2021) shares a similar issue with (Pu et al.
+2021), in that the AC framework it utilizes can actually train the critic with global state information.
+Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX (Wang et al.
+2020).
+(Hu et al.
+2021) experimentally shown that QMIX outperforms QPLEX.
 Additionally, FOP’s proof concept is that under the IGO conditions, if we locally move local policies towards local optimal policies, the distance between the joint policy and the optimal joint policy also decreases.
 In contrast, our proof approach first establishes that our method is equivalent to the single-agent SAC algorithm and then completes the proof using existing conclusions.
-HASAC [16] is a CTCE version of MASAC, with its core contribution being the introduction of sequential decision-making. Therefore, we do not compare it here.
-PAC [38] utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
-Appendix I MPE setup
-In the MPE [18] Spread scenario, three agents and three landmarks are present.
-Agents work together to cover all landmarks while avoiding collisions.
-We created three different maps: Spread(Simple), which matches the original benchmark’s simple_spread; Spread(Medium) and Spread(Hard), which increase the initial distances between agents and landmarks by 1.5X and 2X, respectively.
-Appendix J Hyper-parameters
+HASAC (Liu et al.
+2023), in contrast to FOP (Zhang et al.
+2021), avoids the restrictive Individual-Global-Optimal (IGO) assumption used for joint policy factorization by employing a theoretically grounded sequential optimization of individual policies via joint soft policy decomposition.
+However, our primary contribution addresses the misalignment problem, which is orthogonal to HASAC’s focus.
+Moreover, since HASAC does not rely on either the IGM or IGO assumptions, establishing a direct theoretical comparison is nontrivial.
+Nevertheless, we empirically demonstrate that our method consistently outperforms HASAC across benchmark tasks presented in the paper.
+PAC (Zhou, Lan, and Aggarwal 2022) utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
+Appendix E E. Hyper-parameters
 For all baseline algorithms, we implemented them using the corresponding open-source frameworks and chose the hyperparameters provided by these frameworks for replication.
-We used MAPPO implemented by  Huang et al., QMIX by  Hu et al., FOP by  Zhang et al., and HARL by  Zhong et al..
-Our method, Soft-QMIX, shares the same code base as QMIX.
+We used FOP by  Zhang et al.
+and HASAC by Zhong et al..
+Our method, ME-IGM is built on Hu et al..
 The experiments were conducted on a computer equipped with 92 GB of RAM, a 40-core CPU, and a GeForce RTX 2080 Ti GPU.
 The following are the hyperparameters used in the experiments.
-Hyperparameter Value
-Learning Rate Actor 5e-4
-Learning Rate Critic 1e-3
-Data Chunk Length 8
-Env Num 8
-Episode Length 400
-PPO Epoch 5
-Actor Train Interval Step 1
-Use Recurrent Policy True
-Use Adv Normalize True
-Use Value Active Masks False
-Use Linear LR Decay True
-Table 6: MAPPO hyper-parameters used for SMACv2. We utilize the hyperparameters used in SMACv2 [3].
 Parameter Name Value
@@
 fixed_order False
-Table 7: HASAC hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper [16].
+Table 6: HASAC hyperparameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Liu et al. 2023).
 Parameter Name Value
-Action Selector epsilon greedy
-$\epsilon$-Start 1.0
-$\epsilon$-Finish 0.05
-$\epsilon$-Anneal Time 1e5 (5e5 for MPE)
 Runner parallel
@@
 Mixing Embed Dimension 32
-Hypernet Embed Dimension 64
+Hyper-net Embed Dimension 64
 Learning Rate 0.001
 $\lambda$ 0.4
-Table 8: QMIX hyper-parameters used for SMACv2 and MPE. We utilize the hyperparameters used in SMACv2 [3].
+$\bar{H}$ (zerg, protoss) 0.24 $\times$ num_ally
+$\bar{H}$ (terran) 0.32 $\times$ num_ally
+$\alpha$ Learning Rate 0.3
+Table 7: ME-QMIX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
 Parameter Name Value
-$\alpha$($\_5$ vs _ 5) 0.03
-$\alpha$($\_10$ vs _ 10) 0.005
-Table 9: Soft-QMIX hyper-parameters used for SMACv2. Note that besides the hyperparameters mentioned in the table, all other values for Soft-QMIX are the same as those for QMIX.
+Runner parallel
+Batch Size Run 8
+Buffer Size 5000
+Batch Size 128
+Optimizer Adam
+$t_{\max}$ 1005000
+Target Update Interval 200
+Mac basic_mac
+Agent rnn
+Agent Output Type q
+Learner dmaq_qatten_learner
+Mixer dmaq
+Mixing Embed Dimension 32
+Hyper-net Embed Dimension 64
+Learning Rate 0.001
+$\lambda$ 0.6
+$\bar{H}$ (zerg, protoss) 0.24 $\times$ num_ally
+$\bar{H}$ (terran) 0.32 $\times$ num_ally
+$\alpha$ Learning Rate 0.3
+Table 8: ME-QPLEX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
 Parameter Name Value
@@
 $\lambda$ 0.4
-Table 10: FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper [36].
-Appendix K Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Table 9: FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Zhang et al. 2021).
+Parameter Name Value
+Runner parallel
+Batch Size Run 4
+Buffer Size 5000
+Batch Size 512
+Optimizer Adam
+$t_{\max}$ 1005000
+Target Update Interval 200
+Mac n_mac
+Agent n_rnn
+Agent Output Type q
+Learner nq_learner
+Mixer qmix
+Mixing Embed Dimension 32
+Hyper-net Embed Dimension 64
+Learning Rate 0.001
+$\lambda$ 0.6
+$\alpha$ Learning Rate 0.3
+Table 10: ME-QMIX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).
+Parameter Name Value
+Runner parallel
+Batch Size Run 8
+Buffer Size 5000
+Batch Size 512
+Optimizer Adam
+$t_{\max}$ 1005000
+Target Update Interval 200
+Mac basic_mac
+Agent rnn
+Agent Output Type q
+Learner dmaq_qatten_learner
+Mixer dmaq
+Mixing Embed Dimension 32
+Hyper-net Embed Dimension 64
+Learning Rate 0.001
+$\lambda$ 0.6
+$\alpha$ Learning Rate 0.3
+Table 11: ME-QPLEX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).

### v4

@@
 ME-IGM: Individual-Global-Max in Maximum Entropy Multi-Agent Reinforcement Learning
-Abstract
+Abstract.
 Multi-agent credit assignment is a fundamental challenge for cooperative multi-agent reinforcement learning (MARL), where a team of agents learn from shared reward signals.
@@
 We empirically evaluate two variants of ME-IGM: ME-QMIX and ME-QPLEX, in non-monotonic matrix games, and demonstrate their state-of-the-art performance across 17 scenarios in SMAC-v2 and Overcooked.
-1 Introduction
-Figure 1: The figure illustrates the improvement of our approach compared to existing maximum entropy MARL methods.
+Key words and phrases:
+Multi Agent Reinforcement Learning; Maximum Entropy Reinforcement Learning; Centralized Training with Decentralized Execution
+doi:
+GYYC3346
+1. Introduction
+Figure 1.
+The figure illustrates the improvement of our approach compared to existing maximum entropy MARL methods.
 The left figure shows a straightforward approach to applying maximum entropy MARL in the CTDE context, where blue texts represent the desired objectives, and black texts indicate corresponding constraints.
@@
 This transformation is optimized using a loss function that minimizes the expected difference between $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$, which guarantees monotonic policy improvement in maximum entropy MARL.
+[…]
+This image demonstrates improvement of our approach compared to existing maximum entropy MARL methods
 Collaborative multi-agent tasks, where a team of agents works together to complete a task and receives joint rewards, are inherently challenging for reinforcement learning (RL) agents.
@@
 The challenge is further exacerbated by the exponential growth of the joint action space as the number of agents increases.
-In this work, we propose leveraging maximum entropy RL to promote exploration and adopting the individual-global-max (IGM) credit assignment mechanism (Rashid et al.
-2020) to effectively distribute joint rewards among the agents.
-Maximum entropy RL (Levine 2018), recognized for promoting exploration (Haarnoja et al.
-2018), smoothing the objective landscape (Ahmed et al.
-2019), and improving robustness (Eysenbach and Levine 2021), has been extensively applied in single-agent RL.
+In this work, we propose leveraging maximum entropy RL to promote exploration and adopting the individual-global-max (IGM) credit assignment mechanism (rashid2020monotonic) to effectively distribute joint rewards among the agents.
+Maximum entropy RL levine2018reinforcement, recognized for promoting exploration haarnoja2018soft, smoothing the objective landscape ahmed2019understanding, and improving robustness eysenbach2021maximum, has been extensively applied in single-agent RL.
 However, as illustrated in Figure 1, naively extending maximum entropy RL to multi-agent settings under the centralized training and decentralized execution (CTDE) paradigm leads to uniform credit assignment, which limits the ability of multi-agent reinforcement learning (MARL) agents to learn from joint rewards.
@@
 This limitation highlights a critical challenge in achieving effective credit assignment within the maximum entropy MARL framework.
-This challenge, as we will demonstrate, is not unique to IGM-based methods but also represents a broader obstacle for actor-critic-based maximum entropy MARL algorithms, such as FOP (Zhang et al.
-2021), hindering their ability to achieve optimal coordination.
+This challenge, as we will demonstrate, is not unique to IGM-based methods but also represents a broader obstacle for actor-critic-based maximum entropy MARL algorithms, such as FOP (zhang2021fop), hindering their ability to achieve optimal coordination.
+We propose ME-IGM, the first CTDE maximum entropy MARL method that is compatible with the IGM condition.
+In particular, to address the misalignment between stochastic local policies and the maximum global Q-values, we introduce an order-preserving transformation (OPT) that maps local Q-values to policy logits while preserving their relative order.
+This ensures that selecting actions with the highest logits collectively leads to a joint action that maximizes the global Q-value.
+Additionally, the OPT is trained with a theoretically grounded objective, guaranteeing monotonic policy improvement in CTDE MARL.
 Our main contributions are summarized as follows:
@@
 Additionally, we propose a theoretically grounded and straightforward objective for training these transformation operators.
-  • We confirm the effectiveness of our algorithm through extensive evaluation on matrix games, SMAC-v2 (Ellis et al.
-2022), and Overcooked (Carroll et al.
-2019), where our method attains state-of-the-art results.
-2 Background
-2.1 Multi-Agent Reinforcement Learning
-Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek and Amato 2016).
+  • We confirm the effectiveness of our algorithm through extensive evaluation on matrix games, SMAC-v2 (ellis2022smacv2), and Overcooked carroll2019utility, where our method attains state-of-the-art results.
+[^1] We have published our code at https://github.com/WentseChen/Soft-QMIX
+2. Background
+2.1. Multi-Agent Reinforcement Learning
+Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) oliehoek2016concise.
 Formally, a Dec-POMDP is represented as a tuple $(\mathcal{A},S,U,T,r,O,G,\gamma)$, where $\mathcal{A}\equiv\{1,...,n\}$ is the set of $n$ agents.
@@
 $\rho_{\pi}(s_{t},\mathbf{u_{t}})$ is used to represent the state-action marginals of the trajectory distribution induced by a policy $\pi$.
-2.2 Individual-Global-Max Condition
-Individual-Global-Max (IGM) (Rashid et al. 2020) is a commonly used credit assignment method in value-decomposition-based cooperative MARL and is defined as follows:
-Definition 2.1.
+2.2. Individual-Global-Max Condition
+Individual-Global-Max (IGM) rashid2020monotonic is a commonly used credit assignment method in value-decomposition-based cooperative MARL and is defined as follows:
+[…]
 Individual-Global-Max (IGM) For joint q-function $Q_{tot}$, if there exist individual q-functions $\left[Q_{i}\right]_{i=1}^{n}$ such that the following holds:
-$\underset{\mathbf{u_{t}}}{\mathrm{argmax}}\ Q_{tot}(\mathbf{u_{t}},s_{t})=\begin{pmatrix}\underset{u_{t}^{1}}{\mathrm{argmax}}\ Q_{1}(u^{1}_{t}|o^{1}_{t})\\ \vdots\\ \underset{u_{t}^{n}}{\mathrm{argmax}}\ Q_{n}(u^{n}_{t}|o^{n}_{t})\end{pmatrix}.$ (1)
-Then, we say that $[Q_{i}]_{i=1}^{n}$ satisfy IGM for $Q_{tot}$ under $s_{t}$.
-2.3 Related Works
-Multi-Agent Credit Assignment: CTDE MARL primarily consists of two algorithmic branches: actor-critic methods (Lowe et al.
-2017; Yu et al.
-2022; Lin et al.
-2023; Foerster et al.
-2018; Kuba et al.
-2021; Zhong et al.
-2023) and value decomposition methods (Sunehag et al.
-2017; Rashid et al.
-2020; Wang et al.
-2020).
-Actor-critic methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence between local policies and the softmax of the global Q-value.
-On the other hand, value decomposition methods address the credit assignment problem by decomposing the global value function into multiple local value functions while satisfying the IGM condition.
-In particular, VDN (Sunehag et al.
-2017) assumes that the global Q-function is the sum of local Q-functions.
-QMIX (Rashid et al.
-2020) allows the mixing function, which aggregates local Q-values into the global Q-value, to be any function with non-negative partial derivatives, i.e., $\frac{\partial Q_{tot}}{\partial Q_{i}}\geq 0,\forall i\in\mathcal{A}.$ QTRAN (Son et al.
-2019) introduces an additional set of inequality constraints to construct its loss function.
-QPLEX (Wang et al.
-2020) uses the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize.
-Our method, ME-IGM, enhances the exploration capability of these MARL algorithms by leveraging the maximum entropy RL framework and is compatible with all the aforementioned value decomposition methods.
-Maximum Entropy MARL: Maximum entropy RL is demonstrated to encourage exploration in single-agent RL (Haarnoja et al.
-2018).
-Most maximum entropy MARL works utilize an actor-critic architecture, minimizing the KL divergence between the local actors and the softmax of the centralized critic (He et al.
-2022; Guo and Wu 2022; Zhang et al.
-2021).
-However, as demonstrated later, these methods cannot guarantee that the joint action (at each state) determined by local policies maximizes the global Q-value, thus violating the IGM condition.
-In this case, we propose a novel order preserving transformation to ensure that the IGM condition is met while employing maximum entropy MARL.
-3 Misalignment Between Local Policies and the Maximum Global Q-Value
+$\underset{\mathbf{u_{t}}}{\mathrm{argmax}}\ Q_{tot}(\mathbf{u_{t}},s_{t})=\begin{pmatrix}\underset{u_{t}^{1}}{\mathrm{argmax}}\ Q_{1}(u^{1}_{t}|o^{1}_{t})\\ \vdots\\ \underset{u_{t}^{n}}{\mathrm{argmax}}\ Q_{n}(u^{n}_{t}|o^{n}_{t})\end{pmatrix}$.
+(1)
+Then, we say that $Q_{i}]_{i=1}^{n}$ satisfy IGM for $Q_{tot}$ under $s_{t}$.
+2.3. Related Works
+Multi-Agent RL.
+In the CTDE framework, MARL primarily involves two common algorithmic approaches: policy gradient methods lowe2017multi; yu2022surprising; lin2023tizero; foerster2018counterfactual; kuba2021trust; zhong2023heterogeneousagent and value function decomposition methods sunehag2017value; rashid2020monotonic; wang2020qplex.
+Policy gradient methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence.
+On the other hand, Value function decomposition addresses the credit assignment problem by decomposing the global value function into multiple local value functions.
+VDN sunehag2017value assumes that the global Q-function is the sum of local Q-functions.
+QMIX rashid2020monotonic; peng2021facmac permits the mixer function to be any function with non-negative weights.
+QTRAN son2019qtran optimizes an additional pair of inequality constraints to construct a loss function.
+QPLEX wang2020qplex use the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize.
+The expressive power of the critic networks in these algorithms increases sequentially, but experimental results show that QMIX has better performance hu2021rethinking.
+Therefore, in this work, QMIX is used as the mechanism for credit assignment.
+Maximum Entropy MARL.
+Maximum entropy RL levine2018reinforcement is demonstrated to have advantages such as encouraging exploration haarnoja2018soft and increasing robustness eysenbach2021maximum in single-agent RL scenarios.
+In maximum entropy MARL, most works utilize an actor-critic (AC) architecture and typically aim to minimize the KL divergence between the local actor and the centralized critic he2022multiagent; guo2022learning; pu2021decomposed; zhang2021fop.
+Comparison with Actor-Critic Approaches.
+Most existing maximum entropy MARL methods (e.g., mSAC pu2021decomposed, FOP zhang2021fop, and HASAC liu2023maximum) rely on the Actor-Critic framework.
+Our value-based ME-IGM differs from them in three aspects.
+First, Order Preservation: AC-based methods may suffer from a misalignment where the order of actions preferred by the local actor does not match the local Q-function due to approximation errors.
+In contrast, our method uses Order-Preserving Transformations (OPT) to strictly enforce this alignment, ensuring the IGM condition is met.
+Second, Objective Function: AC methods typically minimize KL divergence to distill policies.
+Our method minimizes the MSE loss between logits, which has been shown to be superior in distillation contexts kim2021comparing.
+Third, Specific Baselines: FOP zhang2021fop employs a credit assignment mechanism similar to QPLEX but relies on the Individual-Global-Optimal (IGO) assumption.
+While FOP argues that locally optimizing policies reduces the joint policy gap, it does not explicitly guarantee alignment.
+HASAC liu2023maximum avoids the IGO assumption through sequential optimization but ignores the misalignment problem between stochastic policies and Q-values.
+Our work identifies this misalignment as a critical bottleneck and proposes OPT to resolve it, thereby achieving state-of-the-art performance.
+3. Misalignment Between Local Policies and the Maximum Global Q-Value
 The Misalignment Problem of Maximum Entropy MARL: While the IGM condition defines the relationship between local Q-values and the global Q-value, what is often more important in practice is the alignment between local policies and the global Q-value.
 In other words, joint actions that maximize the global Q-value can be obtained by querying local policies.
-Previous work in maximum entropy MARL (Zhang et al.
-2021; Guo and Wu 2022) often decompose the global Q-value into individual local Q-values to satisfy the IGM condition.
+Previous work in maximum entropy MARL zhang2021fop; guo2022learning often decompose the global Q-value into individual local Q-values to satisfy the IGM condition.
 These local Q-values are then used to guide the training of local stochastic policies.
@@
 Consequently, the learned policies may struggle to achieve optimal global coordination during execution.
-Figure 2: Illustration of misalignment between local policies and the maximum global Q-value.
+Figure 2.
+Illustration of misalignment between local policies and the maximum global Q-value.
 When naively combining the IGM condition with maximum entropy MARL, local policies often select suboptimal joint actions, leading to a lower Q-value, depicted by the blue curve.
 In contrast, the optimal joint action should achieve the global-max Q-value, shown as the orange curve.
-Empirical Demonstration: To analyze the performance gap caused by the misalignment between local policies and the maximum global Q-value, we evaluate the difference between the maximum global Q-value and the global Q-value selected by local policies in the SMAC-v2 zerg 5vs5 scenario, using a state-of-the-art Maximum Entropy MARL algorithm (Zhang et al.
-2021).
+[…]
+Illustration of misalignment
+Empirical Demonstration: To analyze the performance gap caused by the misalignment between local policies and the maximum global Q-value, we evaluate the difference between the maximum global Q-value and the global Q-value selected by local policies in the SMAC-v2 zerg 5vs5 scenario, using a state-of-the-art Maximum Entropy MARL algorithm (zhang2021fop).
 Figure 2 shows the results, averaged over five random seeds with standard deviations included.
 We can observe that there is a significant Q-value gap, emphasizing the potential for improvement by developing algorithms that enable the selection of joint actions with the highest global Q-value through local stochastic policies.
-4 Order-Preserving Transformation
+4. Order-Preserving Transformation
 Here, we propose a novel order-preserving transformation (OPT) to solve the aforementioned misalignment problem between local policies and the maximum global Q-value.
-4.1 Formulation of OPT
+4.1. Formulation of OPT
 OPT is used to map local Q-values to the logits of local policies while preserving their relative order. Formally, the OPT is defined as:
@@
 Notably, when all entries in $W^{1}$ are non-negative, the relative order of elements in $x$ is preserved in each of the $d_{1}$ hidden variables.
-Next, a nonlinear activation function $\sigma$, such as ELU (Clevert 2015), is applied, followed by another order-preserving linear transformation using $W^{2}$ and $b^{2}$.
+Next, a nonlinear activation function $\sigma$, such as ELU clevert2015fast, is applied, followed by another order-preserving linear transformation using $W^{2}$ and $b^{2}$.
 Consequently, the final output, $\text{OPT}(x,s)$, maintains both the dimension and the relative order of the entries in the input $x$.
@@
 Thus, the proposed OPT addresses the aforementioned misalignment issue commonly found in maximum entropy MARL.
-4.2 Training Scheme of OPT
-Figure 3: The overall pipeline of ME-IGM.
+4.2. Training Scheme of OPT
+Figure 3. The overall pipeline of ME-IGM.
+[…]
+The overall pipeline of ME-IGM.
 Here, we first formulate the ideal objective of Maximum Entropy MARL and its practical counterpart that admits an analytical-form solution.
@@
 Also, we find that the policy improvement using Equation (4) no longer guarantees a monotonic increase in the global Q-value, unlike the approach based on Equation (3).
-Specifically, we have the following theorem:
-Theorem 4.1.
-Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$.
+Specifically, we have the following theorem: {theorem} Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$.
 Updating the policy with Equation (3) ensures monotonic improvement in the global Q-value.
@@
 10:  for each gradient step do
-11:   calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq. 12
+11:   calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq.
 12:   $\theta\leftarrow\theta-\lambda_{\theta}\nabla_{\theta}L(\theta)$
@@
 $\epsilon\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim\mathcal{P}}\big[C\sqrt{2D_{KL}\big(\pi_{\text{old}}(\cdot|s_{t+1})|\pi_{\text{new}}(\cdot|s_{t+1})\big)}\big]$ (6)
-where $\gamma$ is the discount factor, $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function and $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})\triangleq|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}),s_{t+1})|$.
-Proof is provided in the Appendix.
-This theorem implies that we must enforce $\epsilon\leq 0$ to guarantee monotonic policy improvement when using Equation (4) as the objective.
+where $\gamma$ is the discount factor, $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function, $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})$, and $\Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})\triangleq|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})$- $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}),s_{t+1})|$.
+The proof is provided in Appendix B, where we also analyze how this error affects the general convergence properties of policy improvement.
 This motivates us to minimize the factor $C$ and adopt the following objective to train the OPTs.
 $\min_{{\text{OPT}_{1:n}}}\mathbb{E}_{s,\mathbf{u}\sim\rho_{\pi_{jt}^{new}}}[\Delta_{Q}(s,\mathbf{u})^{2}],$ (7)
-$\pi_{jt}^{new}=(\pi_{1}^{new},\cdots,\pi_{n}^{new})$ and $\pi_{i}^{new}$ is defined in Eq. (5).
+where $\pi_{jt}^{new}=(\pi_{1}^{new},\cdots,\pi_{n}^{new})$ and $\pi_{i}^{new}$ is defined in Eq. (5).
+Table 1.
+(a) presents the payoff matrix for a one-step matrix game involving two agents: the row player and the column player.
+Each agent has three possible actions: $\{\text{A, B, C}\}$.
+The rewards for each joint action are given in a $3\times 3$ matrix.
+This payoff structure is non-monotonic, as the optimal action for one player depends on the action chosen by its teammate.
+(b)–(d) illustrate the results of ME-QMIX, FOP, and QMIX, respectively.
+The first row and first column display the policies adopted by the row and column players after $10k$ training steps.
+The $3\times 3$ matrix in the bottom right corner represents the global Q-function $Q_{\text{tot}}$ learned by each algorithm.
+We can see that ME-QMIX is the only method that assigns the highest probability to selecting the optimal joint action, highlighting its effectiveness in overcoming the misalignment issue and achieving optimal coordination.
+$\epsilon$ in (d) denotes the exploration rate used in $\epsilon$-greedy exploration.
 (a) Payoff Matrix
@@
 (d) Result of QMIX
-[…]
-$\epsilon$8.5 -8.5 -8.5
-[…]
-$\epsilon$8.5 0.2 0.2
-[…]
-5 The Overall Framework: ME-IGM
+$\epsilon$ 1-$2\epsilon$ $\epsilon$
+$\epsilon$ -8.5 -8.5 -8.5
+1-$2\epsilon$ -8.5 0.2 0.2
+$\epsilon$ -8.5 0.2 0.2
+5. The Overall Framework: ME-IGM
 Figure 3 illustrates the framework of ME-IGM, which consists of three types of networks.
@@
 For instance, during evaluation, we utilize the Mixer networks of QMIX and QPLEX, leading to two variants of our algorithm: ME-QMIX and ME-QPLEX..
-$[\text{OPT}_{i,\phi}]_{i=1}^{n}$ are $n$ order-preserving transformations, and $\text{OPT}_{i,\phi}(f_{\theta}(o_{t}^{i}),s_{t})$ defines a local stochastic policy for agent $i$.
+$\text{OPT}_{i,\phi}]_{i=1}^{n}$ are $n$ order-preserving transformations, and $\text{OPT}_{i,\phi}(f_{\theta}(o_{t}^{i}),s_{t})$ defines a local stochastic policy for agent $i$.
 The update rules for these networks are described as follows.
@@
 $\mathbb{E}_{\tau_{t}}\left[\frac{1}{2}\left(Q_{tot_{\theta}}(s_{t},\mathbf{u}_{t})-\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2}\right]$ (8)
-where $\tau_{t}=\{o_{t:t+1}^{1:n},u_{t:t+1}^{1:n},s_{t:t+1},r_{t}\}$, $Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big{(}f_{\theta}(o^{1},u^{1}),\ldots,f_{\theta}(o^{n},u^{n}),s\big{)}$, and the estimated Q-target $\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})=r_{t}+\gamma Q_{tot,\theta^{-}}(s_{t+1},\mathbf{u}_{t+1})$.
+where the variables are defined as:
+$\tau_{t}=\{o_{t:t+1}^{1:n},\;u_{t:t+1}^{1:n},\;s_{t:t+1},\;r_{t}\}$
+$Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big(f_{\theta}(o^{1},u^{1}),\ldots,f_{\theta}(o^{n},u^{n}),s\big)$
+and the estimated Q-target is:
+$\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})=r_{t}+\gamma Q_{tot,\theta^{-}}(s_{t+1},\mathbf{u}_{t+1})$
 As a common practice, we utilize a target Q network $Q_{tot,\theta^{-}}$ to stabilize training, with $\theta^{-}$ being the exponentially weighted moving average of $\theta$.
@@
 This offers two advantages: (1) the softmax policy function enhances exploration, and (2) incorporating state $s$ in the sampling process allows for better utilization of global information during centralized training.
-Further, to train $[\text{OPT}_{i,\phi}]_{i=1}^{n}$, we adopt a sample-based version of Equation (7), denoted as $L(\phi)$:
+Further, to train $\text{OPT}_{i,\phi}]_{i=1}^{n}$, we adopt a sample-based version of Equation (7), denoted as $L(\phi)$:
 $\mathbb{E}_{\tau_{t}}\left[\left(\sum_{i=1}^{n}\text{OPT}_{i,\phi}(f_{\theta}(o^{i}_{t},u^{i}_{t}),s_{t})-Q_{tot,\theta}(s_{t},\mathbf{u}_{t})\right)^{2}\right]$ (9)
@@
 $\pi_{i}(u^{i}|o^{i},s)=\text{softmax}\left(\text{OPT}_{i,\phi}\left(f_{\theta}(o^{i}),s\right)/\alpha_{\omega}\right)$ (10)
-$\alpha_{\omega}$ is the temperature and, similar to Haarnoja et al. (2018), can be updated by minimizing the following loss function:
+$\alpha_{\omega}$ is the temperature and, similar to haarnoja2018soft, can be updated by minimizing the following loss function:
 $L(\omega)=\mathbb{E}_{\tau_{t}}\left[-\alpha_{\omega}\log\pi_{jt}(\mathbf{u}_{t}|s_{t})-\alpha_{\omega}\bar{\mathcal{H}}\right],$ (11)
@@
 The pseudo code of ME-IGM algorithm is presented in Algorithm 1.
-6 Experiments
+6. Experiments
 In this section, we first evaluate our method on a classic matrix game, demonstrating its capability to learn a globally optimal policy despite the non-monotonic nature of the overall payoff matrix.
@@
 Moreover, we conduct ablation studies to evaluate the benefits of entropy-driven exploration, assess the necessity of incorporating the order-preserving transformation, and analyze the impact of key hyperparameters.
-Figure 4: The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked.
+Figure 4.
+The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked.
 It shows that ME-QMIX and ME-QPLEX achieve higher returns and exhibit faster convergence, compared to QMIX and QPLEX which do not adopt maximum entropy.
+[…]
+The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked.
+Table 2.
+Average win rates of different algorithms on SMAC-v2.
+The results we recorded demonstrate that our method, ME-IGM, outperforms all previous baseline algorithms and achieves performance comparable to IL+MARL methods.
+Note that the results for MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, and InQ are cited directly from bui2024mimicking as mean values; therefore, no standard deviations are reported for these baselines.
 MARL ME-IGM (Ours) IL+MARL
 Task Scenario MAPPO IPPO QMIX QPLEX HASAC FOP ME-QMIX ME-QPLEX IMAX-PPO InQ
-Protoss 5_vs_5 58.0 54.6 70.2 53.3 22.9 37.8 75.9$\pm$3.0 69.8$\pm$1.8 68.1 78.7
-10_vs_10 58.3 58.0 69.0 53.7 12.1 8.6 78.5$\pm$2.7 78.7$\pm$3.4 59.6 79.8
-10_vs_11 18.2 20.3 42.5 22.8 5.7 0.4 50.4$\pm$2.9 44.9$\pm$3.2 21.3 48.7
-20_vs_20 38.1 44.5 69.7 27.2 - 2.0 73.9$\pm$3.0 38.3$\pm$4.0 76.3 80.6
-20_vs_23 5.1 4.1 16.5 4.8 - 0.3 23.4$\pm$1.0 18.7$\pm$1.3 11.8 24.2
-Terran 5_vs_5 52.0 56.2 58.4 70.0 37.9 57.1 70.2$\pm$2.6 70.0$\pm$4.8 53.3 69.9
-10_vs_10 58.1 57.3 65.8 66.1 20.4 15.8 72.6$\pm$1.8 69.4$\pm$0.2 58.4 72.2
-10_vs_11 28.6 31.0 39.4 41.4 9.1 7.5 49.6$\pm$1.6 51.0$\pm$0.4 28.4 53.9
-20_vs_20 52.8 49.6 57.6 23.9 - 0.2 59.3$\pm$1.3 45.1$\pm$7.2 35.9 65.4
-20_vs_23 11.2 10.0 10.0 7.0 - 0.0 18.7$\pm$0.7 17.4$\pm$2.9 4.7 17.7
-Zerg 5_vs_5 41.0 37.2 37.2 47.8 29.1 35.2 50.8$\pm$4.6 52$\pm$3.9 48.6 55.0
-10_vs_10 39.1 49.4 40.8 41.6 17.9 8.0 49.1$\pm$0.6 42.9$\pm$1.8 50.6 57.6
-10_vs_11 31.2 26.0 28.0 31.1 14.0 1.5 32.6$\pm$1.7 45.3$\pm$0.6 34.8 41.5
-20_vs_20 31.9 31.2 30.4 15.8 - 0.2 34.3$\pm$5.2 42.1$\pm$5.1 26.7 43.3
-20_vs_23 15.8 8.3 10.1 6.7 - 0.2 14.8$\pm$6.4 17.5$\pm$3.2 8.2 21.3
-Table 2: We record the average win rates of different algorithms on SMAC-v2.
-The results demonstrate that our method, ME-IGM, outperforms all previous baseline algorithms and achieves performance comparable to IL+MARL methods, despite not having access to expert opponent demonstrations or opponent modeling.
-6.1 Evaluation on Matrix Games
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+6.1. Evaluation on Matrix Games
 We first demonstrate the effectiveness of our algorithm in a classic one-step matrix game, shown as Table 1.
-It is a non-monotonic payoff matrix commonly used in previous studies (Son et al.
-2019; Wang et al.
-2020).
-We evaluate ME-QMIX, a variant of ME-IGM built upon QMIX, by comparing its performance against two state-of-the-art MARL methods: QMIX (Rashid et al.
-2020), a value decomposition MARL method, and FOP (Zhang et al.
-2021), a Maximum Entropy MARL method.
+It is a non-monotonic payoff matrix commonly used in previous studies son2019qtran; wang2020qplex.
+We evaluate ME-QMIX, a variant of ME-IGM built upon QMIX, by comparing its performance against two state-of-the-art MARL methods: QMIX (rashid2020monotonic), a value decomposition MARL method, and FOP (zhang2021fop), a Maximum Entropy MARL method.
 We are concerned with whether the joint action with the highest selection probability corresponds to the globally maximal Q value, ensuring optimal returns when executing a deterministic policy.
@@
 However, its policy tends to select suboptimal joint actions more frequently due to the misalignment between local policies and the maximum global Q-values, as discussed in Section 3.
+This observation highlights that the misalignment between learned policies and optimal joint actions is not confined to IGM-based approaches.
 ME-QMIX exhibits smaller estimation errors for the payoff of optimal joint actions (which is our primary concern) while having larger errors for suboptimal ones.
 Crucially, the joint action with the maximum Q-value in ME-QMIX aligns with the true optimal joint action, and the local policy assigns the highest probability to selecting this optimal joint action, avoiding the misalignment issue through the use of OPTs.
-6.2 Evaluation on Overcooked
-To demonstrate the advantage of employing the principle of maximum entropy, we compare ME-IGM with two leading value decomposition MARL methods (which do not involve maximum entropy): QMIX and QPLEX, on the Overcooked benchmark (Carroll et al.
-2019) – a widely used coordination-focused multi-agent environment.
+6.2. Evaluation on Overcooked
+To demonstrate the advantage of employing the principle of maximum entropy, we compare ME-IGM with two leading value decomposition MARL methods (which do not involve maximum entropy): QMIX and QPLEX, on the Overcooked benchmark carroll2019utility – a widely used coordination-focused multi-agent environment.
 In particular, we evaluate them in two representative layouts: cramped room and asymmetric advantages.
@@
 As shown in Figure 4, ME-QMIX and ME-QPLEX exhibit significantly faster convergence and achieve higher performance levels in the early stages of training compared to their counterparts, QMIX and QPLEX.
-6.3 Evaluation on SMAC-v2
-We further conduct a comprehensive evaluation of ME-IGM on the SMACv2 (Ellis et al.
-2022) benchmark across 15 scenarios, comparing its performance against 8 baselines.
-The results, as shown in Table 2, are obtained by training each algorithm on each scenario for 101010 millions of environment steps.
+6.3. Evaluation on SMAC-v2
+We further conduct a comprehensive evaluation of ME-IGM on the SMACv2 ellis2022smacv2 benchmark across 15 scenarios, comparing its performance against 8 baselines.
+The results, as shown in Table 2, are obtained by training each algorithm on each scenario for $10$ millions of environment steps.
 Each value in the table represents the average performance over three runs with different random seeds.
 ME-IGM outperforms the baseline algorithms across all scenarios by a significant margin.
-MAPPO and IPPO, as on-policy algorithms, have lower sampling efficiency, resulting in lower win rates within  [^10] M environment steps.
+MAPPO and IPPO, as on-policy algorithms, have lower sampling efficiency, resulting in lower win rates within 1010M environment steps.
 QMIX and QPLEX, lacking exploration, tend to converge to local optimum.
-HASAC (Liu et al.
-2023) and FOP perform poorly across all scenarios.
+HASAC (liu2023maximum) and FOP perform poorly across all scenarios.
 Surprisingly, ME-IGM achieves performance on par with (or even surpassing) IL+MARL methods, despite not relying on expert opponent demonstrations or opponent modeling.
-The results of MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, InQ are taken from (Bui, Mai, and Nguyen 2024).
+The results of MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, InQ are taken from bui2024mimicking.
 A detailed comparison with actor-critic baselines is presented in the Appendix.
-6.4 Ablation Study
-Figure 5: Comparison of exploration strategies in CTDE MARL: ME-QMIX with maximum entropy vs. QMIX with epsilon-greedy and extended epsilon annealing period.
+6.4. Ablation Study
+Figure 5.
+Ablation study on exploration strategies in QMIX.
+Comparison between ME-QMIX and a modified QMIX with extended epsilon annealing on the Protoss 5v5 map.
+The results show that simply increasing epsilon-greedy exploration is insufficient for improving performance.
+In contrast, ME-QMIX enables QMIX agents to perform more structured exploration through the maximum entropy framework, leading to higher returns.
+[…]
+Comparison of exploration strategies in CTDE MARL
 Benefits of entropy-driven exploration: To further demonstrate the benefits of using the maximum entropy framework for exploration, we compare ME-QMIX with a modified version of QMIX, which incorporates significantly higher exploration by applying an extended epsilon annealing period in the epsilon-greedy method, on Protoss 5v5.
@@
 ME-QMIX achieves superior performance, with the systematic exploration enabled by the maximum entropy framework.
+Table 3.
+Our algorithm (+OPT) vs.
+its two ablated versions: +Entropy, obtained by removing OPTs, and QMIX, obtained by further removing softmax local policies.
+Results on three SMAC-v2 scenarios show that removing OPTs (+Entropy) leads to lower win rates due to misalignment between local and global objectives, while incorporating OPTs ensures consistent and monotonic policy improvement.
 QMIX +Entropy +OPT
@@
 Zerg 5vs5 0.41 $\pm$ 0.05 0.40 $\pm$ 0.04 0.48 $\pm$ 0.05
-Table 3: Our algorithm (+OPT) vs. its two ablated versions: +Entropy, obtained by removing OPTs, and QMIX, obtained by further removing softmax local policies.
 The necessity of using OPTs: In Table 3, we compare QMIX with its two variants: +Entropy and +OPT.
@@
 (b) Impact of the target entropy $\bar{\mathcal{H}}$
-Figure 6: The impact of different hyperparameters for training $\alpha_{\omega}$ on the performance of ME-QMIX.
+Figure 6.
+The impact of different hyperparameters for training $\alpha_{\omega}$ on the performance of ME-QMIX.
 (a) A lower learning rate for $\alpha_{\omega}$ corresponds to increased exploration.
 (b) ME-QMIX’s performance is not sensitive to the choice of the target entropy $\bar{\mathcal{H}}$ in Equation (10).
+[…]
+The impact of different hyperparameters on the performance of ME-QMIX
+Table 4.
+Ablation study on the effect of the Order-Preserving Transformation (OPT).
+We compare ME-QMIX (OPT) with a variant that replaces OPT layers with a standard 3-layer MLP, which does not preserve the order of input values.
+Results on three SMAC-v2 scenarios show that ME-QMIX (OPT) consistently achieves higher win rates, highlighting that enforcing order preservation is essential for mitigating local–global misalignment and achieving stable policy improvement.
 ME-QMIX (MLP) ME-QMIX (OPT)
-Protoss 5v5 0.58±0.030.58$\pm$ 0.030.58 ± 0.03 0.74±0.040.74$\pm$ 0.040.74 ± 0.04
-Terran 5v5 0.61±0.040.61$\pm$ 0.040.61 ± 0.04 0.70±0.050.70$\pm$ 0.050.70 ± 0.05
-Zerg 5v5 0.41±0.040.41$\pm$ 0.040.41 ± 0.04 0.48±0.050.48$\pm$ 0.050.48 ± 0.05
-Table 4: ME-QMIX (OPT) vs. its ablated version: ME-QMIX (MLP), where OPT is replaced with a 3-layer MLP.
+Protoss 5v5 $0.58\pm 0.03$ $0.74\pm 0.04$
+Terran 5v5 $0.61\pm 0.04$ $0.70\pm 0.05$
+Zerg 5v5 $0.41\pm 0.04$ $0.48\pm 0.05$
 To further validate the design of the order-preserving transformation, we replace the OPT layers in ME-QMIX with a 3-layer MLP, which does not enforce non-negative weights, referring to this variant as ME-QMIX (MLP).
@@
 Please check the Appendix for the tables of hyperparameters.
-7 Conclusion
+7. Conclusion
 In this paper, we introduce ME-IGM, a novel maximum entropy MARL approach that applies the CTDE framework and satisfies the IGM condition.
@@
 This work was supported in part by the U.S. Army Futures Command under Contract No. W519TC-23-C-0030.
-[…]
-[…]
+Appendix A TD($\lambda$)
+In this work we use TD($\lambda$) to update the Q-value.
+In this paragraph, we will derive the expression of TD($\lambda$) in the context of Maximum Entropy Reinforcement Learning.
+Specifically, we aim to demonstrate the derivation of the following equation:
 $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u_{t}})-Q_{tot}(s_{t},\mathbf{u_{t}})=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{V}_{t+l}$ (12)
@@
 It is important to note that when computing the TD lambda error in the context of a reverse view and maximum entropy reinforcement learning, it is necessary to subtract the $\log(\pi)$ term from the $t+1$ step’s return as well as subtract $\log(\pi)$ from $Q_{t+1}$ in order to obtain results consistent with Equation 15.
-Appendix B B. Proof of Theorem 4.1
-Lemma B.1 (Joint Soft Policy Improvement).
+Appendix B Proof of Theorem 3
+[…]
 Suppose $\pi_{jt}^{\text{old}}\in\Pi$ and consider $\pi_{jt}^{\text{new}}$ to be the solution to Equation 3.
@@
 where we can repeatedly expand $Q^{\pi_{jt}^{\text{old}}}$ on the RHS by applying the soft Bellman equation, progressing from the second line to the final one.
-Theorem B.2.
+{theorem}
 Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$.
@@
 $\leq\mathbb{E}_{s_{t+1}\sim p}\big[C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big],$
-where $C=\underset{\mathbf{u}_{t+1}}{\max}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality (Pinsker 1964), we have:
+where $C=\underset{\mathbf{u}_{t+1}}{\max}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality pinsker1964information, we have:
 $\epsilon\leq\frac{\gamma}{1-\gamma}|\epsilon_{1}+\epsilon_{2}|$ $\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim p}\big[C\cdot\sum_{\mathbf{u}_{t+1}}|\pi_{jt}^{\text{old}}(\mathbf{u}_{t+1}|s_{t+1})-\pi_{jt}^{\text{new}}(\mathbf{u}_{t+1}|s_{t+1})|\big]$ (25)
 $\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim p}\big[C\cdot\sqrt{2D_{KL}\big(\pi_{jt}^{\text{old}}(\cdot|s_{t+1})|\pi_{jt}^{\text{new}}(\cdot|s_{t+1})\big)}\big]$
-Appendix C C. Hyper-network Structure
-Figure 7: (a) The overall network architecture.
+B.1. Convergence Analysis
+In this section, we provide a theoretical analysis of the convergence properties of the proposed ME-IGM algorithm.
+Our analysis leverages the theoretical framework of Regularized Modified Policy Iteration (RMPI) established in previous literature
+B.1.1. Theoretical Foundation: Regularized MPI
+As proved by Geist et al. (smirnova2019convergenceapproximateregularizedpolicy), the convergence of policy iteration with time-varying regularization can be established through the following theorem:
+{theorem}
+[RMPI Convergence] Consider the Regularized Modified Policy Iteration algorithm with time-varying regularization functions $\Omega_{t}$.
+Let the sequence $(\lambda_{t})_{t}$ be a uniform bound for $\Omega_{t}$, such that:
+$\sup_{\pi}\|\Omega_{t}(\pi)\|_{\infty}:=\sup_{\pi,s}|\Omega_{t}(\pi(\cdot|s))|\leq\lambda_{t}.$ (26)
+The gap between the utility of the learned policy after $N$ iterations, $V_{N,\Omega}$, and the optimal utility $V^{*}$ is bounded by:
+$\|V_{N,\Omega}-V^{*}\|_{\infty}\leq\frac{2}{1-\gamma}\left(\Lambda_{N}+\gamma^{N}\|V_{0,\Omega}-V^{*}\|_{\infty}\right),$ (27)
+where $\Lambda_{N}:=\left(1+\frac{1-\gamma^{m}}{1-\gamma}\right)\sum_{t=1}^{N-1}\gamma^{N-t}\lambda_{t}$.
+Notably, if $\lambda_{t}\to 0$ as $t\to\infty$, the algorithm converges to the optimal value function $V^{*}$.
+B.1.2. Mapping ME-IGM to RMPI
+The discrepancy induced by the Order-Preserving Transformation (OPT) and the decentralized policy update can be analyzed as a time-varying regularization term $\Omega_{t}(\pi)$ within the Regularized MPI framework (which we denote as $\epsilon$ in our derivation).
+As derived in Theorem 3, the error term $\Omega_{t}(\pi)$ is bounded by:
+$\lambda_{t}=\frac{\gamma}{1-\gamma}\mathbb{E}\left[\Delta_{Q}\cdot\sqrt{D_{KL}(\pi_{t-1}||\pi_{t})}\right].$ (28)
+Here, $\Delta_{Q}$ represents the alignment gap between the global Q-value and the transformed individual Q-values: $\Delta_{Q}(s,u)\triangleq|Q_{tot}(s,u)-\sum_{i}OPT_{i}(Q_{i},s)|$.
+By substituting Eq.
+28 into the expression for $\Lambda_{N}$ in Eq.
+27, we can quantify the total error induced by the misalignment $\Delta_{Q}$.
+B.1.3. Empirical Convergence Analysis
+Our current empirical evidence strongly supports the convergence of ME-IGM:
+  1. (1)
+    Vanishing $\Delta_{Q}$: Empirical results show that $\Delta_{Q}^{2}$ decreases by three orders of magnitude during the training process, indicating that the OPT effectively aligns the decentralized policies with the global Q-function.
+  2. (2)
+    Stable Policy Updates: The term $D_{KL}(\pi_{t-1}||\pi_{t})$ remains non-increasing during training, suggesting that the policy updates become increasingly stable.
+Given these observations, we can approximately view $\lambda_{t}\to 0$ as training progresses.
+Under this approximation, Theorem B.1.1 guarantees that the ME-IGM algorithm converges to the optimal value function $V^{*}$, fulfilling the IGM condition while enjoying the exploration benefits of maximum entropy.
+Appendix C Hyper-network Structure
+Figure 7.
+(a) The overall network architecture.
 Single Q-Net takes local observations as input and outputs a distribution over local actions $u^{i}$, also denoted as local $Q_{i}$.
@@
 OPT 64 - 1
-Table 5: Hyperparameters used for hyper-networks.
-Appendix D D. Comparison with Baselines
+Table 5. Hyperparameters used for hyper-networks.
+Appendix D Comparison with Baselines
 The following paragraphs present the differences between our method and those based on the Actor-Critic framework:
@@
 Actor-critic methods typically minimize the KL divergence between the actor and the softmax critic, whereas our method minimizes the MSE loss between their logits.
-(Kim et al.
-2021) pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
+kim2021comparing pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
 3.
@@
 In fact, it’s challenging to theoretically prove that value-based methods are inherently superior to AC methods.
-However, through experiments, we have demonstrated that our method surpasses both FOP (Zhang et al.
-2021) and HASAC (Liu et al.
-2023).
+However, through experiments, we have demonstrated that our method surpasses both FOP zhang2021fop and HASAC liu2023maximum.
 Below are comparisons with individual papers.
-mSAC (Pu et al.
-2021) introduces a method similar to MASAC.
+mSAC pu2021decomposed introduces a method similar to MASAC.
 It employs the AC architecture, thus encountering issues 1 and 2 mentioned above, but does not fully utilize global information to train the critic, hence lacking the advantage noted in point 3.
-FOP (Zhang et al.
-2021) shares a similar issue with (Pu et al.
-2021), in that the AC framework it utilizes can actually train the critic with global state information.
-Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX (Wang et al.
-2020).
-(Hu et al.
-2021) experimentally shown that QMIX outperforms QPLEX.
+FOP zhang2021fop shares a similar issue with pu2021decomposed, in that the AC framework it utilizes can actually train the critic with global state information.
+Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX wang2020qplex.
+hu2021rethinking experimentally shown that QMIX outperforms QPLEX.
 Additionally, FOP’s proof concept is that under the IGO conditions, if we locally move local policies towards local optimal policies, the distance between the joint policy and the optimal joint policy also decreases.
 In contrast, our proof approach first establishes that our method is equivalent to the single-agent SAC algorithm and then completes the proof using existing conclusions.
-HASAC (Liu et al.
-2023), in contrast to FOP (Zhang et al.
-2021), avoids the restrictive Individual-Global-Optimal (IGO) assumption used for joint policy factorization by employing a theoretically grounded sequential optimization of individual policies via joint soft policy decomposition.
+HASAC liu2023maximum, in contrast to FOP zhang2021fop, avoids the restrictive Individual-Global-Optimal (IGO) assumption used for joint policy factorization by employing a theoretically grounded sequential optimization of individual policies via joint soft policy decomposition.
 However, our primary contribution addresses the misalignment problem, which is orthogonal to HASAC’s focus.
@@
 Nevertheless, we empirically demonstrate that our method consistently outperforms HASAC across benchmark tasks presented in the paper.
-PAC (Zhou, Lan, and Aggarwal 2022) utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
-Appendix E E. Hyper-parameters
+PAC zhou2022pac utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
+Appendix E Hyper-parameters
 For all baseline algorithms, we implemented them using the corresponding open-source frameworks and chose the hyperparameters provided by these frameworks for replication.
-We used FOP by  Zhang et al.
-and HASAC by Zhong et al..
-Our method, ME-IGM is built on Hu et al..
+We used FOP by zhang2021fop and HASAC by zhong2023heterogeneousagent.
+Our method, ME-IGM is built on hu2021rethinking.
 The experiments were conducted on a computer equipped with 92 GB of RAM, a 40-core CPU, and a GeForce RTX 2080 Ti GPU.
@@
 critic_lr 0.0005
-[…]
-[…]
+$\text{auto\_}\alpha$ True
+$\alpha\text{\_lr}$ 0.0003
 $\gamma$ 0.99
@@
 fixed_order False
-Table 6: HASAC hyperparameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Liu et al. 2023).
+Table 6. HASAC hyperparameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper liu2023maximum.
 Parameter Name Value
@@
 $\alpha$ Learning Rate 0.3
-Table 7: ME-QMIX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
+Table 7. ME-QMIX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 ellis2022smacv2.
 Parameter Name Value
@@
 $\alpha$ Learning Rate 0.3
-Table 8: ME-QPLEX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
+Table 8. ME-QPLEX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 ellis2022smacv2.
 Parameter Name Value
@@
 $\lambda$ 0.4
-Table 9: FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Zhang et al. 2021).
+Table 9. FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper zhang2021fop.
 Parameter Name Value
@@
 $\alpha$ Learning Rate 0.3
-Table 10: ME-QMIX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).
+Table 10. ME-QMIX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked carroll2019utility.
 Parameter Name Value
@@
 $\alpha$ Learning Rate 0.3
-Table 11: ME-QPLEX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).
+Table 11. ME-QPLEX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked carroll2019utility.